A Complete Bibliography of Publications in Perspectives in Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

15 May 2024 Version 1.37

Title word cross-reference

//doi.org/10.1162/posc_{a}_{0}0337[RB22].

1572 [Met97]. 17th [De 06, Gar95a]. 1860s [Fri21]. 1920s [Lef22]. 1930s [Fri21]. 1940s [Lef22]. 1950s [Bul16]. 1960s [Bar20, SD17]. 1970s [Bar20]. 1976-1999 [Bur12]. 19th [De 06].

2001h [Sal00]. 2003b [Sab00]. 28 [RB22].

6 [Sal00]. 6th [Réd14].

8 [Sab00].

Abduction [Mor21]. Abstraction [CM23, Chi20]. Academia [Dei99]. Academic [Sch03]. Acceleration [Lan19]. Acceptability [Tei16].

Acceptance [Sha96]. Accessibility [Jac08]. Accommodation [McI01]. According [Gig12, Urr10]. Account [Ded20, Fis03b, dC22b]. Accountability [Ueb20]. Accounting [Sch21]. Accounts [De 06, Dow09]. Acids [Jac21b]. Acoustics [Wit16]. Across [FH07b]. Act [Che14]. acting [Pic05c]. Action [Ner19]. Actor [Cha23]. Acts [Cor21]. Adaptation [Bur08, Hyu22]. Adaptedness [Fle13]. Additional [Swe13]. Addressing [dMMI14]. Adequality [KSS13]. Adjustment [dC22b]. Ado [DG95, Sny97]. Adventured [Sin14]. Aeolipile [Mar16]. Aeronautical [Ham01]. Aesthetics [Dow09]. aestu [Ako22]. Affair [Pet13]. Affairs [Lef22, Omo22]. Affective [JV23]. Afraid [dMMI14]. After [Ako22, Bae24, Gar21, Des03, Dup12, Wit16]. Afterlives [AAM24]. Against [SP15, Mag11, Sha02]. Age [Leo14, Mét15]. Agency [FH07a]. Agent [Kuh21, Šeš21, Yli14]. Agent-Based [Kuh21, Šeš21, Yli14]. Agents [Mét15]. Ages [Gra04]. Agonistes [Sch95]. Agreement [Lun18]. Agricultural [Har05b]. Aid [IJ16]. Aided [All16]. Air [Bul16]. Alan [Har09]. Alchemical [New09a]. Alchemy [New09b]. Alexis [Hac24]. Alfonsine [Cha02]. Algorithms [JV23, Ste11]. Alignment [Geo17]. Almost [Hol12, KSS13]. Alternative [Sol10]. Always [Rou18]. Amazonia [DS17]. Ambivalent [Car17]. America [Lóp17, Rob17, SDGDV17]. American [RB22, BR20a, Har96, Hyu22, Kev15, Mol17, Ric20]. Americanization [Die96]. Amerindians [Mol17]. among [Joy95]. Anachronism [Sta12]. Analogies [GG11]. Analogy [Nan11, SC19, Tay95]. Analysis [Boh12, Kle05a, Lam03, Nor94, Rud98, Zup17]. Analyzing [Duc03]. Anarchism [Tso03]. Anatomy [Gol16a, Sga22]. Ancient [Kos04]. Ancients [Bab22]. Andrew [Har05a]. Anglo [Har96]. Animal [Sch06, Smi09]. Animals [Fis03b, SJ18]. Animism [Koc21, Koc22]. Annual [Ano97d]. Answer [Zup17]. Answers [Duc10]. Anthropic [Del93]. Anthropological [Cle22, CHB22, JM17, dC22a]. Anthropology [Bar22a, Hyu22, JZ98]. Anti [Car07, Fee07]. Anti-Augustinianism [Car07]. Anti-Metaphysical [Fee07]. Anticipative [FH07a, FH07b]. Antoni [Fra19]. Anything [Pan15]. Ape [FH07a, FH07b]. Ape-Language [FH07a]. Aporetic [Sab98, Sab00, Sal00]. Apparent [Nor94]. Appendix [Swe12, Swe13]. Apples [Gar95b]. Application [Bar15, Chi20]. Applying [FH07a, FH07b]. Approach [BERW16, Rol12, Sha16]. Approaches [Man19, Yea16]. Appropriate [Ore08]. Arabic [Sab00, Sal00, Sab98, Sal00]. Archaeological [Kok19, Wyl96]. Archaeologists [Wyl99]. Archaeology [Ear13, SG01]. Architecture [De 04, Shi13]. archives [Don06]. Argentina [Nov20]. Arguing [Di 11]. Argument [Cra15, Eig20, Ste16a]. Argumentation [RS08]. Ariew [Col18, Péc18]. Aristotelian [Gre00]. Aristotelianism [Hat16, Koc21]. Aristotelico [Gre93a]. Aron [Sim22]. Árpád [Mát06]. Art [Elk99, Jal16, Kin10, dO17]. Articles [Ano12a, Ano13a, Ano13b, Ano13c]. Artifice [Wol09]. Artificial [Gre21]. Artisanal [New09a]. Artists [Gag08]. Aspects [Art98, Eig20, KM15]. Assessment [Sch15, TV22]. Assistant [Jac10]. Association [Kyl22]. Assumptions [Che14]. Astrology [Gra00, Hei15, Mar22, Swe12, Swe13]. astronomical [Sha02]. Astronomy [BG98, Bar02b, Bar22b, Bok14, Bow07, Hut96, Sab98, Ano03b, Gol97, Sab00, Sal00, HZ09]. Asymmetric [PT19]. Asymmetry [Nor94]. Athens [BM04]. Atlas [Len15]. Atom [Cha21, Ach07]. Atomism [Gar95b, Hat16]. Atomist [Fis03b]. Atomistic [SE23]. Atoms [New09a, Urr10]. Attempt [Fre05]. Attending [Nel23]. Attraction [Geo17]. Aufbau [Fee07]. Augustinianism [Car07]. Authority [AO22, Bar22b, BH17, Kyl22, Lug22, Omo22, Tau22]. Autonomous [Sch22]. avant [Kle05c]. Axiomatic [Maj14, Réd14]. Axiomatics [Abr14, Lup14, Sie14].

B [Jac10]. Bachelard [Rhe05a]. Back [Cur21, HO96, Jac08]. Background [Nej19, Nor02]. Backyard [SD17]. Bacon [Cas21b, CV12, Gar21, Gig12, Har12, Jal12, Jal16, KH17, Rus17, You17]. Baconian [Gau00, Lan12]. Bang [Bru93]. Barnes [Pic05c]. Barry [Pic05c]. Based [Kuh21, OA22, Sch15, Šeš21, Yli14, Mar13]. Bayesian [Rud98]. Be [Gal02, Han22, Mag03]. Bear [Fel17]. Bearing [Gun20]. Before [Ako22, Wit16]. Beginning [De 06, Zik01]. begins [Hei10]. behind [Hol12]. Being [dBV13]. Belief [Cam93, Fel17, Nor94]. Believe [Lan14]. Bellarmine [Omo22]. ben [Gol97]. Bending [Bar15]. Benefits [BR20b]. Benjamin [Cha06]. Bernard [Bar02b]. Bernoulli [Syl98]. Bertrand [Hac24]. Between [CD19, Gis09, GY13, Lyn19, SL23, Alo24, Bes15, Bla18, Bur02, Cat07, Che05, Fri21, Gra10, Kel94, Kin10, Mas10, Mát06, PCC^{+}19, Rea13, Stö18, dR24]. betwixt [Slo09]. Beyond [KSS13, Sto24, Kos23]. Bias [Res00]. Biases [Fer23, Juk17]. Biblical [Luo13]. Bibliographic [Lev96]. bid [Jha06a]. Bifurcation [Whi99]. Big [Bru93]. Bio [Tam21a]. Bio-Inspired [Tam21a]. Biochemists [Dei99]. Bioengineering [Ner19]. Biographies [Gre93b]. Biography [Bur12]. Biological [Fis03a, Gis09, KL21, Mag11, Pal15]. Biology [ASOVS15, Blu93, Car11, Cra15, Fag11, Gre00, Gut21, Hun19, Mét15, MMC93, Mol17, Poj11, Rai93, Riv19, Rud98, TPDB21, Toe21, dC20, Gre97, Cla93b]. Biomedical [DS17, Nuk19, SJ18]. Biomimetics [Kro21]. BioParody [Nor21]. Bioscience [Whi99]. BioTechnology [Nor21]. Biran [Alo24, AAM24, Bae24, San24, Vin24, dR24]. Biranian [Hac24]. Biranism [AAM24]. Bits [Urr10]. Blocks [New09a]. Blood [De 04, Lef22, SD17]. Bloor [Rhe05b]. Bodies [Cam16]. Bodin [Bla94]. Body [De 06, DeH99, Hac24, Hei07, Kel94, Sha03]. Body-Machine [Sha03]. Bohm [Hug06]. Bohm-Pines [Hug06]. Bohr [Bok08]. Book [Gou22, Jes04]. Books [Ano93a, Ano93b, Ano94a, Ano95a, Ano95b, Ano96a, Ano96b, Ano97a, Ano97b, Ano98a, Ano99, Ano01a, Ano02, Ano03a]. borderlands [MMC93]. Borrowed [ASOVS15]. Boscovich [Ako22]. Botanical [Bal20, Seg22]. Both [Lan19, Rai93]. Boundaries [Kos23, RC15]. Brain [Chi20]. Brazil [SM20]. Brazilian [dC20]. Brian [New09b]. Bridging [NKvdP22]. Bringing [vD23]. Brittleness [Spe15]. Broadening [KM15]. Bromberg [Ano16a]. Brownian [Nis16]. Bruno [Gra10, SL23, Sim19]. Builders [LL95]. Building [Lef22, New09a, Tay95]. Bureaucracy [Spe15].

c [KH17, Omo22]. C. [Das16, Tub20]. Cabinet [Wol09]. cadente [Pie93]. Caesar [Bla08]. Calculate [Sch21]. Calculus [Jes98]. Calibration [Fra97, Ski04]. Call [Ano97c]. Calvinist [Gar12]. Came [Pal10]. Can [Ant23, CD19, Coo04, Goe11, Han22, Hla18, Rol99]. Capturing [SGD17]. Cardano [Mar22]. Carnap [Dam22]. Cartesian [Ari21, Col18, Gau00, Gor05, Gre93a, Pet18, Sch18, Sch95]. Cartesianism [Lew07, Sto24]. Cartesians [Ari94, Ari18, Col18, Péc18, Sch18]. Cartesiomania [Des95]. Cartography [Bal96, Len15]. Case [AM19, ASOVS15, Bur01, Cam09b, Chv20, DS07, Ell00, Ger13, Gis09, Har05b, Hol12, Jes07, Kle05a, LL95, Mag11, Miz20, Ore08, Pas12, Pit01a, PCC^{+}19, PH21, Riv19, Seg22, SP15, Van15, Way96, Zol97, Cha02]. Cases [Gre93a]. Cassirer [Hei10, Moy03]. Categorical [MR14, Mol17]. Categories [Che94, PCC^{+}19, Sch03]. Cauchy [KK11]. Causal [Des19, Kos19b, Lan19, Toe21]. Causation [Cha23, Hat00]. Causes [Pal10]. Cautionary [JZ98]. Cavaillès [Hyd03]. Cavell [Kin10]. Cavendish [Lew01]. CCR5 [Jac15a]. CDF [RS08]. Cecile [Bla94]. Cell [Fag11, Liu16]. Censor [Rei97]. Censorship [Ari94, Gor96]. Centennial [Pie06]. Center [Cla93b]. Century [AMM19, Jac21a, Wat98, All97, BG98, Bla18, De 06, Dij08, Dör01, Gar95a, Gar95b, GG11, Gis02, Gre93a, Jal12, Kle05a, La 04, Lin19, Lüt00, Man19, PB23, Ric02b, Sha03, Sto24, Tam21b, YB16, Zik01]. CERN [Kar17]. Certainty [Dom10]. Certified [Nor94]. Chagas [CP99]. Challenges [MMMN19]. Challenging [Cop18, Jac15a]. Chance [Len10, Nor23]. Change [CD19, Che05, Che14, Hof18, Mir22, NKvdP22, Sch06, Spe04, Ste16a, TV22]. Changes [Kai96]. Chaperoned [Bur08]. Chaplin [Hol12]. Chapter [Nis16]. Charles [Hor05, Kin93]. Chemical [Bes15, Ell00, Jac21b, JR23, Nee96]. Chemistry [Bar15, Kor94, Llo21, Ram98, WH20, Zol97]. Chemists [Dei99]. Chene [Hat00, Men00]. cherts [Le 93]. Chicago [Blu93, Cit93, Cla93b, Kin93, MMC93, Rai93]. Children [Hyu22]. Chimeras [Mir22]. Choice [Mar96, Sch14]. Choke [Le 93]. Choke-holds [Le 93]. Chomskyian [Har94b, Har94a]. Chrestienne [Gar12]. Christian [Cam16]. Christoph [Gor96]. Chromosome [RB22, San20]. circa [All16, Cla93b, Stu97]. Circle [Ueb08, Ueb20]. Circularity [Cha95]. Circulation [Wil21]. Cities [MC22]. Citizen [BR20b]. Cladistic [Ric02a]. claims [Sup93]. Clarification [Ell00]. Clarifying [Sha16]. Classical [Cam07, GMHP21]. Classification [KH17, Pie06, VD17]. Clerk [Cha21]. Climate [BH17, Che14, Gun20, Par14, PH21, Ste16a, TV22, Wil21]. Clinic [MM15]. Clock [DS07]. Closing [vdS12]. Coding [Ste11]. Coenesthesia [Alo24]. Cognition [Che14, Smi12, SD17, SDGDV17]. Cognitive [Bar01, Fer23, Ger13, Ner19, PCC^{+}19, Stu14, Twe06b]. Cognitive-Historical [Twe06b]. Cognitivism [Dam22]. Cohen [Moy03]. Coherence [Bro16a, Ric08]. Cohesion [Art98]. COI [Jan02]. Collaboration [Lyn24, RS08]. Collaborations [Kos23]. Collaborative [Bar20, Kar17]. Collecting [Pan15]. Collective [Lyn24]. Collèges [Bro95]. Collider} [Kar17]. Colloid [Twe06a]. Colonial [Roq22]. Coloniality [Bar22a]. Colonies [Kev15]. Color [Jac08, Mar09, SvJB02, Sha96]. Colors [Dij08]. Combinatorics [Nor23]. Combined [Hat16]. Comet [Met97]. Commensurability [Nor99]. Commentary [Gou22, dC22a]. Comments [Bur02, Dup12, Har05a, Lon08, Sab00, Kai96, Sal00]. Commercialization [Pin15]. common [Vic95]. Communication [Bul16, HE22]. Communicatory [Kar17]. Communities [Eck16, Wra97]. Community [Joh09, MSF94, Mik97, Gar95a]. Companion [SJ18]. Comparative [Tan19, Chv20]. Comparing [GMHP21]. Comparison [Man13, Ste11]. Compass [Bia22]. Compiling [Len15]. Complexity [Gre93b, Li 21, Li 23, SE23]. Comprehending [BERW16]. Computation [MR14]. Computational [Car11, Das08, Joh09, Vig22, WH20]. Computationalism [Chi20]. Computer [All16, Ear13, IJ16, SE23]. Computer-Aided [All16]. Computers [Kel94]. computing [MH11]. Comte [Pin20]. Conant [Jac10]. Concept [Bor21, Bro23, MSF94, Nuk19, Sha03, Smi12, GH05]. Conception [Dam22, Gra10]. Conceptions [Goo18, HS17]. Concepts [Ant23, Bok14, Cam07, Che05, Smi18, Wit16]. Conceptual [CD19, Che05, Ell00, Gut21]. Conceptually [Kle05a]. Concerns [dMMI14]. Conditions [CF15, Ink17]. Condorcet [Pin20]. Conference [Ano97d, Ano97c]. Configuring [Sab98, Sab00, Sal00]. Confined [Cha21]. Confirmation [McI01]. Conflation [Sha16]. Conflict [Omo22]. Conflicting [Goo18]. Conflicts [Ell16, Res98]. Confusion [Pal15]. Connection [Kin10, Jha06a]. Connections [CHB22, PCC^{+}19]. Connexion [Psi04]. Conscience [Ric08].

Consciousness [Sch06]. Consensus [Ore08, Wyl08, dMMI14]. Consent [Zah17]. Consequences [Ham01]. Conservation [Fel17]. Considerations [Gut21]. Considered [Tha14]. Consolidation [Bro23]. constitution [Gar95a]. Constitutional [vD23]. Constitutive [Maj14, Ric02b]. Constructing [All97, Bar02a, Bor21, Cam09a, Tau22, Tuu01]. Construction [Das16, Goo18, Tay95]. Constructionism [Wra10, Tay95]. Constructive [Sta12]. Constructivism [Gal02]. Constructivist [Gin16]. Contamination [Gra01, HS01]. Contemporary [Dup04, Jes10, Zol97]. Content [Ant23, Zup17]. Contested [RC15]. Context [BR20a, Cam09b, Gre93b, Hed07, RB22, Tuc07]. Contexts [Das16, RH03a]. Contextual [Duc10, Sch02, Rup06]. Contextualist [Col18]. Continental [Ore08]. Contingency [McD10]. Continuing [Cha06]. Continuum [Art98, KK11]. Contributes [Bor19]. Contributions [CF15]. Contributors [Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano04, Ano05b, Ano05c, Ano05d, Ano05e, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano08a, Ano08b, Ano08c, Ano09a, Ano09b, Ano09d, Ano09e, Ano09f, Ano09g, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12b, Ano12c, Ano12d, Ano12e, Ano13d, Ano13e, Ano13f, Ano13g, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano15d, Ano16b]. Control [KL21, Sol96]. Controversies [ASOVS15, Bue19]. Controversy [Bla18, DS07, La 04, Mil18, Whi19, Joy95]. Coordination [Eck16]. Copenhagen [Cam09a]. Copernican [Omo17, Wes13]. Copernicus [Bar02a, Swe12, Swe13]. Cordelia [Sha95]. Core [Che14]. Corpore [Pie93]. Corpses [De 06]. Corpuscles [Bla08]. Correlation [Cha23]. Correspondence [Len10, Syl98]. Coselection [Cam93]. Cosmic [Bru93]. Cosmogenesis [DG95]. Cosmogony [Beg20]. Cosmological [Kra97, Omo22]. Cosmologies [Gau00]. Cosmology [AO22, Bab22, Kra13, SP23, Tes22, VFN22]. Cosmopolitan [Cit93]. Cosmopolitics [vD23]. Cottage [Lev10]. Coulston [Hor05]. Counteract [Fer23]. Counterfactuals [Psi04]. Courts [Kev15]. Cousin [AM19, Whi19]. Cowles [Cit93]. Craft [You17]. Crafting [And06]. Creating [Ink17]. Creativity [Das08]. Credentialling [Sup93]. Credibility [ASOVS15, Car17, Sha95]. Cremonini [Omo22]. Crises [BERW16]. Crisis [Car17, SGD17]. Critical [Hei07, Hei10]. Criticism [Sol10]. Critics [Sol08]. Critique [Chi20, DG95, Gin16, Gra10, Li 23]. Crops [Hic17]. Crossing [MMC93]. Crossroad [Mas10]. Crossroads [dCGLP21]. Crucis [Mar09]. Cultivation [Har12]. Cultural [Lyn19, Rea13]. Culture [Gal99, Gra01, Kus19a, LR94, Wat14]. Cumulative [Spe04]. Curious [Smi12]. Current [ASOVS15, Fer23, Gre97, Pas12, Sch14]. Cybernetic [Chi20, Hof18]. Cyborg [Pic95, Sen00, Ano95c]. Cycloid [Jes07]. Cytogenetic [Ric20]. Cytogenetics [RB22, San20].

Damned [Ari94]. Daniel [Beg20, Bla11]. Darkness [Roj20]. Darstellungen [Ham02]. Darwin [Fle13, Gre93b, Len10, Mag11]. Darwin/Gray [Len10]. Darwinian [Nan11]. Darwinism [Cra15]. Data [Bul16, Cur21, Hof20, Leo14, Rea13, Vig22]. David [MR14, Rhe05b]. Davy [Le 93]. Dayton [Kon14]. Dead [Lau00]. Deal [Sch22]. Dealing [SE23]. Debate [La 04, Pas12, Bok08, Sab00, Sal00, Sny97]. Debates [AMM19, Kos19b, RC15]. Debiasing [Tei16]. Decentering [Pic05a]. Dechales [Mal22]. Decision [MC22, OA22, Wyl08]. Decision-Making [MC22]. Decolonial [Hla18]. Defects [Nor94]. Defending [Ste16a]. Defense [Bul16, Nas18, Sch15, Sha02, Rup06]. Deference [Bou16]. Define [AB23]. Defining [Ink17, Mik97, Tes22]. Degérando [Man19]. Della [Rus17]. Delusion [Gig12]. Demarcating [CBRG17, Ink17]. Demarcation [Bow07]. Democracy [Hic17, Liu16]. Democratic [Van15]. Demonstration [McD03]. Dennis [Men00]. Dependence [GM21]. Depicted [Jac10]. Deprecation [You17]. Depth [Kos19a]. derived [Che94]. Desaparecidos [SGD17]. Descartes [Ale07, Ari07, Ari18, Bal20, Bro95, Col18, Des95, Des03, Gar01, Glo95, Hat00, Hut16, Jes07, Joy95, Len95, Pet13, Rou18, Sch18, Sch95, Sto24, Str23]. Design [All16, Hof01, Höf21, Len10]. Designs [Ell16]. Details [Ore08]. Determination [Cok20]. Development [Coz10, Jes10, Nuk19]. Developmental [Nur13]. Developments [dMMI14]. Deviance [Pan15]. Devil [Ore08]. Devo [Cra15]. Deweyan [HE22]. d’Histoire [Wol09]. Diagrammatic [Kar17]. Diagrams [Bro18a, DR18, Kar17, Mey18, Stö18]. Did [Chv20, Lan14]. Diderot [Wol09]. Different [Wüt18]. Differentiating [All97]. Diffusion [Cha02]. Digital [Ear13, Leo14]. Dilemma [Bur01, Pit01a]. Dilemmas [Wyl96]. Dimension [Gut21]. Dimensional [Bue19]. Dimensions [And06, Cok20]. Diophantus [KSS13]. Dirac [Bok08, Pas12, Sta12]. Directed [FH07a, FH07b]. Disagreement [Lun18]. Disagreements [Che94]. Disappearance [Gut21]. Disappearing [DS17]. Disciplinary [Hat97, Wyl96]. Discipline [LL95]. Disciplines [PT19]. Disciplining [Jal16]. Disconnect [BH17]. Discontinuity [BS01]. Discordant [Fra95]. Discourse [Pea17]. Discoveries [Zol97]. Discovering [Twe06a]. Discovery [Bor19, Cam09b, Cop18, Fra93a, Jos03, Kis06, Twe06a].

Discursive [Gal09]. discuss [Pie93]. Discussing [Ako22]. Discussion [Len10]. Disease [Bid16, CP99, DS17]. Diseases [Mat05, SD17]. Disembowelled [De 06]. Disharmony [Gag08]. Disponent [Geo17]. Disputes [Ach07]. Dissensions [JV23]. Dissent [Nas18, dMMI14]. Distinction [Clo08, Sta99]. Distrust [Bou16]. Disturbance [Hof01]. Disturbances [HS01]. Disunities [Wyl99]. Divergence [Fle13]. Division [Art98]. dixit [AO22]. Do [AB23, Ari94, Bro18a, Kon14, Lüt00, Che94]. Dobzhansky [SM20, dC20]. Doctrine [Kev15]. documenting [GH05]. Does [Del93, Hut16, Hut96]. Doesn’t [Sch21]. doi.org [RB22]. Doing [Nis18]. Domains [All97, Tan19, Sis95]. Double [dC22b]. Double-Adjustment [dC22b]. Doubt [Pal10]. Douglas [Ded20]. down [Rup06]. Drang [Har94b]. Draughtsmen [Fra19]. Drawing [Lop09]. Dreams [Kon18]. Dreyfus [SM20]. Drift [Ore08]. Driven [CW18, Kle05a]. Driving [JR23]. Drop [Ano94b, Stu94]. Drops [Jac21a]. Drosophila [Mar20]. DSM [Bue19]. DSM-Controversies [Bue19]. Ducheyne [Gui13]. Duhem [Nee96]. Dunes [Cit93]. Dupré [Kou12]. During [Jac01]. Dutch [Bal20]. Dyes [Pic05b, Pic05a]. Dynamic [Cit93]. Dynamica [Duc98].

Early [AO22, Ari06, Bla94, Bow02, Des95, FH07a, FH07b, Fis03a, Har96, Hei15, Hen19, Kle05a, Kus19a, Kus19b, LR94, Mar16, Rei19, Ric96, RB22, San20, Tes22, Ueb08, You17, Bar20, Bal20, Cas21a, Glo95, Man19, Seg22]. Early-Modern [Hei15]. Earth [Sin14]. Ecliptic [HZ09]. Ecology [Bro23, Cit93, Ink17, MSF94, Mik97, vD23]. Economic [Han06]. Economics [Car99, GY13, Mäk99]. Economicus [Sch21]. Economist [Rei97]. Eddington [Ryc03]. Editor [Jha06b, Pit00, Pit01b, Twe06b]. Editors [Ano01b, Ano09c, BN99]. Edmond [Sin14]. Education [Bur99, Bur08, Har05b, PCC^{+}19, Van11]. Effort [Lyn24]. Effortlessness [McA16]. Egalitarianism [Ger13]. Egger [dR24]. Eighteenth [Dij08, GG11, Lin19, Sto24, Wat98]. Eighteenth-Century [Dij08, Sto24]. Einstein [Bok08, How93, Lan14]. Elective [Eig20]. Electrical [Kra97]. Electrodynamics [Wüt18]. Elements [Sti14, Gou22]. Elusive [Bar05]. Embryology [Vig22]. emergence [Syl98]. Emergentism [Bla18]. Emerging [Liu16]. Empires [Seg22]. Empirical [Ach07, Clo08, Gur06, Miz20, OA22]. Empirical/Non [Clo08]. Empirical/Non-Empirical [Clo08]. Empiricism [AMM19, Dam22, Gar21, Leh03, Lon08, Man19, Mil18, Rup06, Ueb13, Whi19, Wyl08]. Empiricist [Mas10]. Encyclopaedia [Van11]. End [De 06]. Ends [Des00]. Energy [Shi13]. Engagement [Sim22]. Engineering [Kou16, Nor02, YB16]. England [Ric96]. Enlightenment [Ari06, Hor05]. Enquiry [dC22b]. Enriching [SP23]. Ensemble [Riv19]. Entrenched [Sha16]. Enzyme [Mar96].

Epidemiology [Vig22]. Epigrams [vD23]. Epilepsy [Bla11, De 04]. Epistemic [ASOVS15, BR20b, Bid16, Blo05, CF15, Cur21, Hof18, KM15, ON17, Rhe05b, Ric08, WH20, Wyl08, Kou16]. Epistemological [Hun19, IJ16, Juk17]. Epistemologist [Rei97]. Epistemology [Cam16, Gin16, Hyd03, Jes07, Kro21, Mer23, Mey18, Mir22, PT19, Sga22, Sim22, Wyl08]. Equal [KSS13]. Era [Bro16a, Mag11]. Ernst [De 04, Moy03, Poj11]. Errata [Ano94b, Ano95c]. Erratum [Ano20, RB22]. Error [All97, All01, FH07a, Nor94, Rud01, Sch02]. Especially [Hol12]. Essay [Cla93a, Des95, Fra93b, Gie94, Gre93b, Gre97, Har96, Har97, Lev96, Ram98, Ric96, RB22, Row97, San20, Sch95, She94, Vic95]. established [Wat98]. Establishing [Car17, Nor99]. Ethical [LWR20, Wyl96]. Ethics [Kai96, KC96, Sol96]. Ethnography [Ner19, Ris00]. Eucharist [Ale07]. Euclid [Gou22]. Eugene [Bro23]. Eugenics [Kyl22]. Euler [Ped08]. Europe [VFN22]. European [Koc22]. Europeans [Mol17]. Evaluate [Ach07]. evaluation [Bru93]. Eve [Ach07]. Even [Dre23]. Event [Che05]. Events [Sta99]. Ever [Sch01]. Ever-Present [Sch01]. Everything [Rou18]. Evidence [Bes15, BS01, Ded20, Jan02, Li 21, OA22, Vic00]. Evidence-Based [OA22]. Evo [Cra15]. Evo-Devo [Cra15]. Evolution [BR20a, JR23, Lyn19, RB22, Roj20]. Evolutionary [ASOVS15, Cra15, FG12, Hun19, Nor94, Rud98, TPDB21, dC20]. Evolutionay [Vic00]. Exact [Car99]. Examined [Dre23]. Examples [Vig22]. Except [Dre23]. Exchange [MGY13]. Exercice [Bro95]. Exigencies [Wüt18]. Existence [Alo24]. Experience [AMM19, Fee07, Gol16a, HE22, Jal16, Jos05, Man19, Vin19]. Experience/Science [Fee07]. Experiencing [Mor13]. Experiment [Bai99, Daw16, Elg14, Gar95a, Lug22, McD03]. Experimental [AM19, Fra93b, Jal16, Kar17, Kle05a, Mar16, Tei16, Tuu01, Vin19, Zup17]. Experimentation [Bal20, KL21, Mar96, Vin19]. Experimenter [Zup17]. Experimenting [Bur08, Cav06, Jal16, Mat20]. Experiments [Bok01, Bro14, Che94, Fag11, FS14b, Hop14, Kle05a, Kra97, Nej19, Pal10, Sch22, Ste02, Stu14, Tha14]. Experimentum [Mar09]. Expertise [Bou16, Kyl22]. Experts [Eig20, Lun18]. Explain [Del93]. Explaining [Bro14, Jeb18]. Explanation [Des19, Ger13, Jan02, Mar13, Ris00, SE23]. Explanations [Bat19, CW18, Kos19a, Kos19b, Lan19, Riv19]. Explanatorily [Lan14]. Explanatory [Kos19a, NKvdP22, OS18]. Explication [Nee96]. Exploit [Wyl99]. Exploration [Bar21, FH07a, KL21, Tay95]. Explorative [Sch22]. Exploratory [Bor21, FGS21, Fri21, Kuh21, Wil21]. Exploring [And06, Dou16, Gre21, Mir22, Sch01, Šeš21, Van15]. Exposing [Sch02]. Expulsion [Dei99]. Extension [Bla13]. Extravagant [Sin14].

F [Cob12]. Facing [Car17]. Facts [Leh03]. Fail [Chv20]. Fairness [Wyl08]. faith [Gor96]. Fall [CP99, Pal10]. Family [Bar21]. Faraday [And06, Cav06, Goo06, Twe06a, Twe06b]. Farbenlehre [Jac08]. Fast [DS17]. Fate [Tan19]. Features [Sch02]. Feeling [Alo24]. Feminine [Roj20]. Feminism [Dup12]. Feminist [Rol12, Sis95]. Fermat [KSS13]. Fertility [Lóp17]. Feyerabend [Tso03]. Feynman [Bro18a, DR18, Mey18, Stö18, Wüt18]. Fiction [Elg14]. Fictionalism [Pur18, Sha02]. Field [Goo06, PT19, Réd14]. Fieldwork [Ric20]. Fieldwork/Cytogenetic [Ric20]. Fights [Han06]. Financial [BERW16, Ell16, Res00]. ‘Fine [Sta12]. ‘Fine-Tuning [Sta12]. First [Ari18, Bae24, Bie04, Col18, De 06, Gou22, Péc18, Sch18, Twe06a, dO17, Ano97d, Ric02b, Tam21b]. First-hand [De 06]. Fish [Gar12]. Fishing [Bar21]. Fission [Stu94, Ano94b]. Flayed [De 06]. Fleck [Gin16, Hed07]. Fleming [Cop18]. Flexibility [Die93]. Florentine [De 06]. Fluctuations [Yea16]. Folk [Bok14]. Forays [BN99]. Forces [Toe21]. Foreign [Roq22]. Forerun [FS14b]. Forgetting [Wri16]. Forgot [Nor23]. Form [Ari06, Gut21, Lup14, Tam21b, Tam21a]. Formal [MR14, Sie14]. Formalizing [Boh12]. Formation [Bor21, Wyl96]. Forms [Bla11, Coo04, Toe21]. Forthcoming [Ano05a, Ano12a, Ano13a, Ano13b, Ano13c]. Fortunio [Bla13]. Forwards [Cur21]. Foucault [Hyd03]. Found [Bar21, Twe06a]. Foundation [Gal09]. Foundational [For10]. Foundations [Bar02b, Gal09, Jes98, Ano03b, Sab00, Sal00]. Founder [Gar21]. Four [Bar02b, Tan19]. Fractals [Glo95]. Fragrance [Bar15]. Framework [Ded20]. Framing [Cam09b]. France [AMM19, Hor05, Kle05a, Sch03]. Francis [CV12, Gig12, Har12, Jal12, Jal16, Rus17]. Frank [Tub20]. Franklin [Cha06]. Franz [Ric20]. Free [Gun20, Hud16, Mäk99, Pal10, SGMB18, Wal01]. Frege [Hei10]. French [Sto24, Bro95, Gis02, Sch18, Tuc07]. Freud [Hol12, SG01]. Frey [Bla94]. Friedman [Ric02b]. friends [Joy95]. Fringe [CBRG17]. Frontiers [Dor00]. Frontlines [Gar04]. Function [Ari06, Kuh21]. Fundamental [Cam09b, Cha21, Duc10]. Fusion [Rei19]. Future [Loo02].

G [Sab00]. Gabriel [Mat20]. Galilean [Pal04]. Galilei [Kon18]. Galileo [Ari04, Ber04, Bia22, Bie04, Daw16, Gar04, Gor96, Hei12, Jes04, Mas10, Mil18, Omo17, Pal10, Sga22, She94]. Galison [Sta99]. Games [Zam06]. Gap [NKvdP22]. Gaps [Slo09, vdS12]. Gassendi [Bro95, Cor21, Fis03b, Gar21, Gar95b, Glo95, Joy95, Len95, Osl11, Pal04]. Gaston [Rhe05a]. Geddes [Van11]. Geminus [Bow07]. Gender [Nov20, Rol99, SGMB18]. Gendered [RB22, San20]. Gene [Bar21, Jac15a, MM15, SP15]. Genealogies [Mar20]. Genealogy [MG23]. General [Hof18, Wil21]. Generalized [Ste16a]. Generation [Des03, Fis03b, Smi09]. Generative [Smi09]. Genes [Bar21, Die96]. Genesis [Har05b, Wüt18]. Genetic [CHB22]. Genetically [Hic17]. Geneticists [Bar20, Die96]. Genetics [Kyl22, Mar20, SM20, Ski04, dC20]. Genome [Bal96, Bar21, Hun19]. Genomic [Mag11]. Genomics [Dup04]. Geocentric [VFN22]. Geology [Rai93, Di 11]. Geometrical [Dom10]. Geometries [Ryc03]. Geometry [Glo95, Goo18, Gou22, Sti14]. Georg [De 04, VFN22]. George [Sha02]. German [Bar22a, Har97, Har05b]. Germany [Dei99, Hof01, Moy03, Rem04, Wat98]. Gerson [Gol97]. Gestalt [Cat07, Fee07]. Gestalts [Cat07]. Getting [Koc11, dBV13]. Giambattista [Rus17]. Giants [Glo95, Kou16]. Giere [Kou12]. Gilbert [Geo17, Koc21, Lug22]. Gilles [Bab22]. Gillispie [Hor05]. Giordano [Gra10]. Girolamo [Mar22]. Glasses [Wat14]. Gleiter [Nor09]. Glimpse [Psi04]. Global [Nor99]. Gloom [Roj20]. Glouberman [Len95]. Goal [Che94]. God [Slo09, Gra10]. Gods [Glo95, Gou22, Pie94]. Goethe [Jac08, Mar09]. Golden [Sta99]. Goldschmidt [Die96]. Goldstein [Bar02b]. Goldsteinian [Sha02]. Good [HS17]. Gorlaeus [Hat16]. Governance [Cha23]. Gradual [Sha96]. Grand [BA11, Kra97]. Grant [Gol13]. Granularity [Des19]. Graphic [Ear13, Kus19b, Van11]. Gravitational [Mag03]. Gravity [Fra10]. Gray [Len10]. Great [Wit16]. Greece [Lef22]. Greek [Sab00, Bow02, DeH99, Sal00]. Grey [Des19]. Gross [Har09]. Grossman [Fre05]. ground [Vic95]. Grounding [Bou16]. Group [Lef22, SM20]. Guidelines [Ano93c, Ano93d, Ano93e, Ano93f, Ano94c, Ano94d, Ano94e, Ano94f, Ano95d, Ano95e, Ano95f, Ano95g, Ano96c, Ano96d, Ano96e, Ano96f, Ano97e, Ano97f, Ano97g, Ano97h, Ano98b, Ano98c, Ano98d, Ano05b, Ano05c, Ano06a, Ano09a, Ano09b]. GWS [Sch14].

Habit [dR24]. Hadron [Kar17]. Half [Pia24]. Half-sleep [Pia24]. Halley [Sin14]. Hallucinations [Smi18]. hand [De 06]. Harmful [Tha14]. Harmonising [Psi04]. Harmony [Art98, Mir22, Wat98]. Harms [Ste16b]. Harper [Gui13]. Harvey [Gol16a]. Harwood [Pic05b]. Hattab [Des00]. Health [Sha03]. Heartbeat [Pet13]. Heather [Ded20]. Heavens [Ham16]. Hegel [FG12]. Hegemony [HO96]. Heidegger [Koc11, dBV13]. Heilbron [Wes13]. Heisenberg [Cam07]. Helen [Rup06]. Helmholtz [McD03]. Henricus [Str23]. Henry [Cit93]. Heraclitian [Pit01a]. Herbert [Nor09, Aug00, Sen00]. Heredity [BR20a, Fis03b, Kyl22, RB22]. Hermann [Moy03]. Hermeneutic [Gin16]. Herrick [Kin93]. Herschel [Cob12]. Hertz [Ham02]. Hesitancy [Gol16b]. Hessen [Fre05]. Heterogeneity [Mol17]. heterogeneous [Tay95]. Heuristic [Ban06, Kis06, Sch15]. Heuristics [Mir22]. Hevelius [Tau22]. Hidden [VFN22]. Hierarchy [Hut16, KH17]. Hilbert [Abr14, Lup14, MR14, Réd14, Row97, Sie14, Sti14]. Hippocratic [DeH99]. Hippocratism [Mar22]. Historian [Mal22]. historians [Pie93]. Historical [Ash19, Cam09b, Hyd03, KM15, Ore08, Osl11, Sim22, Twe06b]. Historicist [Mir22]. Histories [Jos03, Lan12, MH11]. Historiographical [Man19]. Historiography [Pal93, Pin20]. History [Ano97c, Ano97d, Ano09c, AS12, Bai99, BM04, Bar02b, Bow02, Bur01, Bur02, CD19, Col18, Dor00, Elk99, Gal09, Gar12, Gra00, Jac15a, Jal12, Jan02, Kin05, Kle05a, La 99, Mla07, Nis16, Nor23, Pal93, Pih08, Ram98, Rem04, Ric97, Ric02b, RB22, San20, Sim22, Smi12, Ste02, SB02, Ste11, TPDB21, You17, Ano95c, Ano03b, Mát06, Pic95, Vic95, RH03b]. Hobbes [Jes93, Jes04]. holds [Le 93]. Hollow [Sin14, Ste16b]. Homage [MR14]. Homo [Sch21]. Honor [Bar02b]. Hooke [Hen19, Lew06]. Hormesis [Ell00, SF00]. Horrocks [Bar22b]. Household [Shi13]. HPS [Sch11]. https [RB22]. Hughes [Ric20]. Hughes-Schrader [Ric20]. Human [CHB22, Das08, Gig12, Hyu22, Mol17, MGY13, RB22, Roj20, San20, Smi12, SDGDV17, dC17, Bal96, Hun19]. Humanistic [Kin93]. Hungarian [Jha06b]. Hunt [Ste16b]. Husserl [Hyd03]. Huygens [Mag03]. Hybrid [Hyu22, Nuk19]. Hydraulic [Kou16]. Hypernormal [CSB^{+}23]. Hypotheses [Ham16]. Hypothesis [Sin14, Wyl99]. Hypothetical [Cob12].

Iberian [CE04, Seg22]. Ibero [RB22, BR20a]. Ibero-American [RB22]. IBM [Wri16]. Idea [Rea13]. Ideal [Hud16, Kit99, Kou12, SGMB18, Sti14]. Idealism [Ryc03]. Identities [ON17]. Identity [And06, Wyl96]. Ideologies [RC15, Rol99]. Idiosyncrasy [Bar21]. Idiosyncratic [Ros22]. Idolatry [Jal12]. Idols [CV12]. if [Ari94]. Ignored [CE04]. II [Ano95c, FH07b, Har94b, Pic95]. Illustration [Lop09]. Image [DeH99, IJ16, Sta99, dO17, BN99, Gal99]. Image/Logic [Sta99]. Imagery [Smi18]. Images [Dup08, Elk99, Fra19, De 06]. Imagination [CV12, Dom10, Gra04, PH21]. Imaging [RB22, San20]. Immaterial [Bla13]. Immediate [Alo24]. Immediately [Che14]. Immersion [Kou16]. Impact [Kos23]. Impediments [Sch01]. Implications [Bau03, Kai96, OS18]. Important [Kra13]. Imposing [KBG22]. Imre [Don06, Mát06]. Incentive [Sch15]. Incentive-Based [Sch15]. Including [Jac10]. Inclusion [Hic17]. Incommensurability [Bar01, Bes15, BS01, Sha16, Che94]. Incompleteness [For10]. Independent [Smi18]. Indeterminacy [Cam07, Sha16]. Index [Rei97]. India [Bar22a]. indígenas [SD17]. Indigenous [DS17, VD17]. induction [Sny97]. Inductive [Bid16, Ded20, Eig20, Ste16a]. Inductivist [Cob12]. Inequality [Lyn19]. Infectious [DS17, Mat05].

Inference [Cam16]. Inferential [Kok19]. Infinitesimal [Jes98]. Infinitesimalmethode [Moy03]. Influence [Poj11, Rol99]. Influx [Wat98]. Information [Ant23, Sol96, Wri16]. Informed [Zah17]. Infuence [Gur06]. Ingold [Koc22]. Ingrained [SC19]. Initiative [Gal09]. Innovation [Bla94]. Inquiries [JV23]. Inquiry [CM23, Cob12, Gig12, Juk17, Kle05a, Šeš21, SDGDV17]. Insider [JZ98]. Inspired [Tam21a]. Institutes [Hof01]. Institutional [Bou16]. Instruction [Gar12]. Instructions [Bia22]. Instrument [Dup08, HZ09, LL95, LR94, Sch21, Ste11, Zik01]. Instrumentalism [BG98]. Instrumentation [Bor19]. Instruments [Sch01, Zik01]. Integrated [TV22]. Integrating [AS12, Ded20, LWR20]. Intellectual [Jac15a, Omo22]. Intelligence [Gre21]. Intelligent [Len10]. Interactions [Gis09]. Interactive [All16, FH07a]. Intercalation [Nor99]. Interdependence [Hut16]. Interdependent [Smi18]. Interdisciplinarities [Ner19]. Interdisciplinarity [Ash19, Mat05, Pie06]. Interdisciplinary [BERW16, MMMN19, Man13, MGY13, ON17, PCC^{+}19]. Interest [Ell16, Res98]. Interests [Res00]. Internal [Bai99, Pal93, Zup17]. International [Ano97c, Kyl22, dC22b]. Internationalist [Van11]. Interplay [Wol09]. Interpretation [Ano94b, Cam09a, Kok19, Leo14, Met97, Stu94]. Interpretative [Die93]. Interpreting [TPDB21]. Intersection [Kle05a]. Intersubjective [Ueb20]. Introducing [HZ09]. Introduction [AO22, AAM24, ACT22, AMM19, AS12, Ari04, BN99, Bro18b, Cas21a, CHB22, CGJ12, Don06, FS14a, FGS21, Gar95b, How97, Jac15b, Jha06b, Kle05b, Kus19b, Lup14, MMMN19, Mod09, Pit01b, Ros22, SV23, SM20, SB02, Twe06b, Van16, dCGLP21, dRP14]. Inventions [Bia22, Kev15]. Investigating [MMMN19]. Investigation [IJ16]. Invisible [Nor09]. Involved [Sha16]. Iraq [Pic05b]. Islandworld [CW18]. Isn’t [Dre23]. Isolates [Roq22]. Isolation [Gra01]. Issue [Ano05a, BR20a, RH03b, RB22]. Issues [Fer23, Gre97, LWR20]. Italian [Omo22]. Italy [Omo17]. Itself [Ric02b].

J. [Jac10]. Jacob [Syl98]. James [Cha21]. Janet [Dup12, Vin19, Vin24]. Japan [Hyu22]. Japanese [Hyu22]. Jean [Ach94, Bla94]. Jean-Cecile [Bla94]. Jeremiah [Bar22b]. Jesuit [Gor96]. Jewish [Dei99]. Joad [Tub20]. Joan [Ano16a]. Johannes [Beg20, Tau22]. John [Cob12, Wes13]. Joint [Ano09c]. Jollos [Die96]. Jonathan [Pic05b]. Jones [Le 93]. Joseph [Ako22, Fra10]. Journal [Juk17]. Journals [Ano09c]. Judgments [Ell16, SF00]. Judson [Kin93]. Julius [Bla08]. Just [Seg22]. Justice [Nuk19]. Justification [Pin20, Whi99, Fra93a].

Kac [Nis18]. Kalmár [Gur06]. Kant [Dom10, Goo18, Jan04, Sch03, Wat97]. Kantian [Mas10]. Kantianism [Hei07, Obe14]. Karl [Kra13]. Kepler [Gin02, GH05, HZ09]. Kettlewell [Rud01]. Key [Dre23]. Kin [Lyn19]. Kind [Dre23]. Kinds [Bok14]. Kinematic [Sab98, Sab00, Sal00]. Kinks [Smi12]. Kinship [Mar13]. Kintsugi [Mir22]. Klein [New09a]. Know [Rou18]. knowing [Pic05c]. Knowledge [Boh12, Bor19, Cha21, Che05, DS17, Hat16, Joh05, Joh09, PCC^{+}19, RS08, Ric20, Seg22, Spe04, Tes22, Tub20]. Kochan [dBV13]. Kourany [Dup12]. Kuhn [Bar01, Ach01, Buc10, Jac10, Kin05, Kin10, Lev10, MG23, Mla07, Pin20, Pit01b, Ric02b, Sol10, Wra10, dO17]. Kuhnian [Gar01, Ric02a, Sch14, TPDB21].

L. [Tub20]. Labor [Lef05, You17]. Laboratories [Ner19]. Laboratory [Goe11, MM15, Ski04, Lyn24]. Lakatos [Don06, Gur06, Jha06a, Jha06b, Kis06, Mát06, Pal93]. Lakatos-Polanyi [Jha06a]. Lamarck [Gis09]. Lamarckism [Gis02]. Language [CD19, FH07a, FH07b]. Languages [MR14]. Large [Kar17]. Largest [Bar21]. László [Gur06]. Late [Gis02, Jal12]. Later [For10, Fra10, Sch11, Kin05]. Latin [Lóp17, Mol17, Rob17, SDGDV17]. Latour [Mer23, Nel23, SL23, Sim19, Tur23]. Latourian [vD23]. Law [Gor05, Jac15a, vD23, Pal10, Pie93, Pie94]. Laws [Des00, KH17, Pea17, Wat97]. Leapt [Cop18]. Learn [FH07b, Goe11]. Learned [Gar04]. Learning [CF15, FH07a, FH07b]. leaves [Hei10]. Leeuwenhoek [Fra19]. Left [Dam22, Ueb20]. Left-Wing [Dam22]. Legacy [Cur21, Hor05, Hun19, Koc21, Lew01, Pan15, Pit01b]. Legal [LWR20]. Legitimacy [Cla93b, Omo22]. Leibniz [Art98, Bae24, Duc98, Duc03, GA98, Jes98, McD10, Syl98, Wat98]. Leibnizian [Loo02]. Leonard [Jac10]. Leonhard [Ped08]. Lessons [Dow09, Llo21]. lettre [Kle05c]. Leuven [Pet13]. Levels [Bro23, Wüt18]. Levi [Gol97]. l’homme [Str23]. Liceti [Bla13]. Life [Cam16, Coo04, Des03, Hut16, LWR20, Mor13, Sch06]. Light [Bla13, Das08, Dij08, Mar09, Nej19, Ped08, Sha96]. Like [Seg22, Cop18]. Lillie [Cla93b]. Limit [Nis16]. Limited [Fle13]. Limits [Cam07, Car99, Car11, Cha21, Die93, Gig12, Mar22, Pin15]. Line [Ger13]. Link [Gur06]. Liquid [Ano94b, Beg20, Stu94]. Liquid-Drop [Ano94b, Stu94]. Lisa [Ano16a]. Literary [LR94]. Literature [Dör01, Vic00]. Lithic [Lop09]. Litigation [Gol13]. Live [Lau00]. Lizard [Sol03]. location [Pic05c, Sha16]. Locke [Dow97]. Lockean [Dow97]. locker [Le 93]. Logic [Abr14, Ari21, Cas21a, Cas21b, Cor21, Elk99, Gar21, Kis06, Pet18, Sta99, Zol97, BN99, Gal99]. Logical [Dam22, Ueb13]. Logico [Wag08]. Logico-Mathematical [Wag08]. Logics [Hla18]. logistic [Hei10]. Long [Lau00]. Longer [CE04]. Longino [Rup06]. Looking [Cur21, Kou16, Wat14]. Looting [Wyl96]. Lore [vdS12]. Lost [Beg20, Gut21]. Love [Gar04, Sha95, Vis01]. Lunar [Mil18]. Lyell [Di 11].

M [Tub20]. Mach [Poj11]. Machine [Chi20, Kel94, Sha03]. Machines [Jes10]. made [Duc10]. Magia [Rus17]. Magnete [Geo17, Koc21]. Magnetic [LL95]. Magnetism [Jos05]. Magni [Nej19]. Magnitudes [Jes98]. Maharashtra [Bar22a]. Mahoney [CK13]. Maine [Alo24, AAM24, Bae24, San24, Vin24, dR24]. Make [Sin14]. Makers [LL95]. Making [Dur10, JM17, Kus19b, Lóp17, Mat05, MC22, OA22, Sha08, SP15, SGD17, dC17]. Makiritare [Mol17]. Malebranche [Det03]. Mammalian [Bar21]. Man [De 06]. Manifesto [Ueb08]. Manipulating [DS07]. Manual [You17]. Manuscript [dO17]. Mapping [BR20b]. Marburg [Gio17, Obe14]. Margaret [Lew01]. Maria [Kon14, RB22]. Marin [Kon18, Urr10]. maris [Ako22]. Mark [Nis18]. Market [Mäk99, Sch21, Wal01]. Marks [Loo02]. Marvelous [Cit93]. Marx [Dor00]. Maser [Bro16a]. materia [Gar12]. Material [Fri21, Gal99, Spe15, Tam21a]. Materialized [KL21]. Materials [JM17, Nor09, Pal15]. Mathematical [Bia22, Cam09b, Fri21, Kos19b, Wag08, Yea16, Jes93, Syl98]. Mathematician [Nis18]. Mathematics [Gur06, Jes07, Jha06b, MC10, Mal22, Obe14, Pal10, Ros22, Stö18, Stu97, Mát06, Pie94]. Mathematization [Mas10]. Matter [Bie04, Gau00, Jeb18, Kor94, Mat20, Sch14, Gor96]. Matters [Bar21, Hei12, HS01]. Maupertuis [Mag03]. Maxwell [Cha21]. may [Mag03]. Mean [You17]. Meaning [And06, Bes15, Nor99, Wat14]. Means [JV23, Nel23, SL23, SV23]. Measurement [GM21, Llo21, Mor21, Tal21, dCGLP21, Cha95]. Measurements [Che00, Cle22]. Measuring [Nis16, PCC^{+}19, Sch21]. Mechanical [Das16, Gol13, Slo09, You17, Bro95]. Mechanicks [You17]. Mechanics [Jes10, Ham02]. Mechanism [Jes10, Sha03]. Mechanisms [JR23, Kos04, Psi04]. Mechanistic [Wüt18]. Media [KL21]. medica} [Gar12]. Medical [Boh12, Gol13]. Medicine [ACT22, Ano09c, Blu93, Cam16, CV12, DeH99, Har96, Jal12, Vin19]. Meditations [Hla18]. memes [Bur12]. Memetics [Chv20]. Memoriam [Ano16a]. Memories [Bar05]. Mendel [Ore10]. Menn [Des00]. Mental [Smi18]. Merely [Slo09]. Mersenne [Ber04, Coz10, Gar04, Gra10, Kon18, MC10, Pal10, Urr10]. Mess [Sch20]. Messengers [Gra00]. Mestizos [Mol17]. Meta [Bes15, Boh12]. Meta-Analysis [Boh12]. Meta-Incommensurability [Bes15]. Metabolism [VD17]. Metallic [Twe06a]. Metaphor [SC19].

Metaphors [CM23, Mik97]. Metaphysical [Fee07, Gor05]. Metaphysics [Slo09]. Metascience [JM17, SF00]. Metempsychosis [Pia24]. Meteorological [Mar22]. Method [AM19, CW18, Jac21a, KSS13, Maj14, Miz20, PB23, Jes93]. Methodological [Ach07, Fer23, Gut21, MMMN19, OS18, SF00]. Methodology [Cas21a, For10, Hla18, Kis06]. Methods [Mik97, Sch01, Tei16]. Metrology [Dör01]. Mexican [SD17]. Mexico [Bar20, Roj20, VD17]. Michael [And06, CK13, JM17, Twe06b]. Micro [Gra01]. Micro-Organisms [Gra01]. Microbiome [OS18]. Microphysics [Gal99]. Microscope [La 04]. Microscopy [La 99, Sch01]. microwaves [Bru93]. Mid [SD17]. Mid-1960s [SD17]. Middle [Gra04, Kok19]. Middle-Range [Kok19]. Migrantes [SGD17]. Migrate [Seg22]. Militaristic [Nel23]. Mill [Sny97]. Milliet [Mal22]. Mind [Bla13, CV12, Cor21, Har12, Hei07, Smi12]. Mind-Body [Hei07]. Minds [Gre21]. Minimal [Kos19a, dBV13]. Miriam [Cle22, Lon08]. Mirror [Wat14]. Miscitation [Vic00]. Misconceptions [MJLG23]. Misled [SC19]. Misogyny [Cla93a]. Misrecognition [Gue17]. Missemination [Tub20]. Missing [And19, Gur06, SGD17]. Misunderstanding [Gol16b, Bur12]. Mixed [MC10]. Model [Ano94b, Duc03, GY13, HE22, Hof18, Höf21, Mar13, Mar16, Sch14, Stu94, Vig22, Way96]. Model-based [Mar13]. Modeling [Bor21, Ear13, FGS21, Gre21, Ham16, Joh09, KL21, Kou16, Kuh21, Luo13, Man13, Mar13, Mat05, Mor13, MGY13, Nis18, Nur13, Rea13, Rio13, Sab98, Stö18, TV22, Sab00, Sal00, Tay95]. Modelling [KM15, Šeš21, Wüt18]. Models [Aug00, BJ02, Bar15, CF15, De 06, Dow09, FGS21, Fri21, KBG22, Kro21, Rea13, Rio13, Sch15, Shi13, Ski04, Van15, WH20, Wil21]. Modems [Bul16]. Modern [AO22, Bal20, Bla94, Cas21a, Fis03a, GMHP21, Har96, Hei15, LR94, Mar16, Seg22, Smi09, Tes22, You17, dC20]. Modernism [Eig20]. Modernity [Glo95]. Modes [Gre21]. Modified [Hic17]. Molecular [Ach94, DS07, SE23]. Molecules [Bar15, Joh09]. Money [Cla93b]. Monochord [Urr10]. Monomorphism [Gra01]. Moon [De 04]. Moral [Bur99, NKvdP22, San24]. Morphogenesis [Höf21, Tam21b]. Morphogenetic [Kro21]. Morphology [Tam21a, Toe21]. Most [Kra13]. Motion [CHB22, Gor05, Nis16, Wat97]. move [GH05]. Movement [Cor21, Kyl22]. Movements [CHB22]. Movius [Ger13]. MR1696321 [Sal00]. MR1873205 [Sab00]. Much [CE04, DG95, Sny97]. Muckraking [Mla07]. Multi [Bue19]. Multi-Dimensional [Bue19]. Multifaceted [Hun19]. Multilevel [Riv19]. Multiple [Cok20, Hof20, KL21]. Multiverse [Pea17]. Mundane [Kra97]. Music [Jac01]. Mutability [Die96]. Mysteries [Kon18]. Myth [Cam09a, Fuc93]. Myths [Tal21].

Name [Hyu22, Ueb13]. Nanotechnology [Nor09, Gal09, Joh09]. Narrating [Ric02b]. Nash [Jac10]. Nation [Bar22a, Lef22]. National [Gal09, Har97]. Nationalism [dC22a]. Natural [Bla94, Bok14, Gar12, Gre93a, Har12, Jal12, Lan12, Lev96, Lüt00, Mar16, SM20, Smi12, Sch95]. naturalis [Rus17]. Naturalism [Lon08, Wol09]. Naturalistic [Sol95]. Naturalizing [Kuk08]. Nature [Ari21, DR18, Ink17, KH17, Kev15, Mar09, Mas10, McD10, Ros22, Wol09, Gar95a, Jes04]. Naturelle [Wol09]. Nazi [Dei99, Rem04]. Neglected [KC96]. Neither [Rai93]. Neo [Cra15, Hei07, Jal12, Obe14]. Neo-Darwinism [Cra15]. Neo-Kantianism [Hei07, Obe14]. Neo-Stoic [Jal12]. Network [Cha23]. Networks [Bar20, Kyl22, Mar13]. Neumann [For10]. Neurath [Rei97, Van11]. Neurons [Wat14]. Neuropsychiatry [Blu93]. Neutral [Sch14]. neutrality [Rup06]. Neutrinos [Fra10]. Never [Rou18]. Newton [Ako22, CW18, Dij08, Dup08, Gag08, Gui13, Mag03, Mar09, Pie93, Pie94, Sha96, Sha08, Slo09, Wat97]. Newtonian [Dow97]. Newtonianism [Dow97, Ric08]. Nexus [Dou16]. Nietzsche [Kon14, SP23]. Nine [Sha08]. Nineteenth [AMM19, Dör01, Gis02, Kle05a, La 04, Man19, PB23]. Nineteenth-Century [Dör01, Gis02, Kle05a, La 04, Man19]. Nineties [She94]. No [Kev15, Lug22, Lun18, Sab00, Sal00]. Nobel [Kor94]. Noel [Wes13, Sto24]. Noise [Bro16a, Bul16, Mon17, Nis16, Wit16, YB16]. Non [Dam22, Des19, Eig20, Ger13, Jac21a, Kos19b, Lan19, OA22, Sch21]. Non-Causal [Des19, Kos19b, Lan19]. Non-Cognitive [Ger13]. Non-Cognitivism [Dam22]. Non-Empirical [Clo08, OA22]. Non-Experts [Eig20]. Non-Market [Sch21]. Non-Uniform [Jac21a]. Nonsense [TPDB21]. Normal [Mag03]. Normality [Pan15]. Normative [Gun20, Ste16a]. Normatively [Ore08]. Norms [Lon08]. Norwegian [Kyl22]. Note [Ano01b, Pit00]. Notes [Ano93g, Ano94g, Ano94h, Ano94i, Ano94j, Ano95h, Ano95i, Ano95j, Ano95k, Ano96g, Ano96h, Ano96i, Ano96j, Ano97i, Ano97j, Ano97k, Ano97l, Ano98e, Ano98f, Ano98g, Ano04, Ano05d, Ano05e, Ano06b, Ano06c, Ano06d, Ano06e, Ano08a, Ano08b, Ano08c, Ano09d, Ano09e, Ano09f, Ano09g, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12b, Ano12c, Ano12d, Ano12e, Ano13d, Ano13e, Ano13f, Ano13g, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano15d, Ano16b, PT19]. Nothing [DG95]. Notion [Fle13, Rhe05a]. Nova [Met97]. Novelty [Kin10]. Novum [Cas21b]. Nuclear [Ano94b, Goe11, LL95, Stu94]. Numbers [KBG22]. Numerical [Ber04]. Nutrition [VD17].

object [Bur12]. Objective [RS08]. Objectivity [BR20b, Cam93, Dör01, Ger13, Gie94, Gol13, Kuk08, Lin19].

Objects [Ham02, HS01, Tuu01]. Observation [You16, Zah17]. Observations [Mil18]. Observing [Ric96]. Occasionalism [Det03]. Occult [New09b]. Odum [Bro23]. off [Hei10]. Offspring [Hon98]. Old [Pin20]. Once [Goe11]. One [Hac24, Lug22]. Ontic [CW18]. Ontic-Driven [CW18]. Ontological [Che14, HO96, Mer23]. Ontology [Pur18, SL23]. Opacity [Kou16, WH20]. Operational [Gal09]. Operations [Höf21, Vig22]. Opportunities [Cha06]. Opportunity [DS17]. Optical [HZ09]. Opticks [Dij08, Gag08, Sha08]. Optics [Coz10, McD10]. Oranges [Gar95b]. orbits [GH05]. orbs [GH05, Sha02]. Order [KBG22, Ste16a]. Ordered [Pin15]. Ordinary [CD19]. Organ [Kon18]. Organic [Duc03]. Organism [Kel94, Lau00]. Organisms [Gra01]. Organization [Kyl22]. Organizational [Fuc93]. Organized [Kyl22]. Organum [Cas21b]. Oriented [ASOVS15]. Origin [Ano94b, Lyn19, Stu94]. Origins [FS14b, Nor09, Ris00]. Other [Coo04, Jac10, JV23, Kou16, Nel23, SL23, SV23, Ear13]. Otlet [Van11]. Otomi [VD17]. Otto [Rei97, Van11]. Outcomes [Tei16]. Outsider [JZ98]. Overdiagnosis [Bid16]. Own [Hac24].

P [Bro23]. Palætiological [Tan19]. Paleontological [Mir22]. Paleontology [Rai93]. Paper [Fri21, Har05a]. Papers [Ano97c, Bar02b]. Paradigm [Hoo18, Nee96, Ore10, Sch22, Sch06]. Paradigms [Cop18]. Paradox [Han06]. Parameterization [WH20]. Paratext [Tau22]. Paris [Ari04, La 04]. Parsimony [Ric02a]. Part [HZ09, Ano05a]. Participant [Zah17]. Participation [Dur10]. Particles [Cha21]. Partnership [Mar20]. Pascal [Ale07, Ari07, Car07, Jes07]. Past [Dre23, Loo02, Sch02]. Pastoralism [Gar12]. Pasts [Ear13]. Patent [Jac15a]. Patenting [Pal15]. Patents [Bro16b, Jac15a, MM15, SP15]. Path [Sol10]. Pathogenic [Gra01]. Pathological [De 04]. Patrick [Van11]. Patterns [Kis06]. Paul [Bok08, Stu14, Van11, Vin19]. Paulo [dC20]. pective [PB23]. Pedagogy [Jos05]. Peer [Juk17]. Peirce [Pie06, PB23]. Peirce-pective [PB23]. People [CHB22, Seg22]. Perception [McD03]. Perfect [Fle13]. Performative [And06]. Period [Bal20, Cas21a]. Peripheral [Das16]. Perrin [Ach94]. Persistence [Kha24]. Personne [Bab22]. Perspect [Sab00, Sal00]. Perspective [Ash19, Hun19, Jac21b, KM15, MM15, Row97, Syl98]. Perspectives [Gin16, GMHP21, Har96, JZ98, RB22]. Pharmacology [Bla18, Lin19]. Phase [Nis18]. Phenomena [Duc03]. Phenomenology [Goo06]. Phenomenotechnique [Rhe05a]. Philipp [Tub20]. Philosopher [Jan04]. Philosophers [Gal02]. Philosophical [Duc10, Ham01, Llo21, McD03, Nej19, dC22b, Kin05]. Philosophically [Kra13]. Philosophies [Jha06b]. Philosophy [Alo24, Ano97c, Ano97d, AS12, Bae24, Bai99, BM04, Bla94, Bur01, Bur02, Cat07, Col18, Dow09, FS14b, Fis03a, Gie12, Gig12, Goo18, Gre93a, Gun20, Gur06, Hac24, Har12, Hoo18, Kin05, Kos04, Lev96, LWR20, Lun18, Lüt00, Mar16, Miz20, Mla07, Pih08, Pit01a, Ram98, Ric97, RH03a, San24, Sch03, Sol95, Ste02, SB02, Tub20, Ueb20, Wyl99, Bro95, Gre97, Hei10, Mát06, RH03b, Sch95, Vic95, Dup12]. Phlogiston [All97]. Phoranomus [Duc98]. Photometers [Che00]. Phylogeography [ASOVS15]. Physical [Ant23, Art98, Bow07, Kra13, Nis16, Pal10, Rol99, Wat98, Gol97, Sha02]. Physicist [Goe11, Nis16]. Physicists [Tub20]. Physics [Ban06, Car99, Gor05, Hof01, Maj14, MR14, Nis18, Poj11, Stö18, Wri16, YB16]. Physiologia [Men00]. Physiology [Alo24, And19, Hut16]. Piaget [Bur08, Cav06]. Pickering [Har05a]. Pictures [Dow09]. Picturing [Mey18]. Pierre [Cor21, Gar12, Vin24]. Pines [Hug06]. Pipe [Kon18]. Pittsburgh [BM04]. Place [Clo08]. Plagiarism [Str23]. Planet [Bok14]. Planetary [Bow02, Ham16]. Planning [MC22]. Plants [Fis03b]. Plaster [Fri21]. Plato [Hut96]. Plattes [Mat20]. Plausible [Cam93]. Player [Das16]. Playing [Bab22]. Plein [Bro95]. Pluralism [Car17, Cra15, Van15]. Pluralist [Bue19, Fel17]. Pluto [Bok14]. Pneumatic [Kon18]. Poems [Tau22]. Poetry [Bar22b]. Point [Nis16, Rud01, Hei10]. Poisons [Bla11]. Polanyi [JM17, Jha06a]. Polar [Fel17]. Police [Hei15]. Policy [Bau03, BH17, CBRG17, Dou16, Dur10, Ell00, Gun20, Liu16, Mét15, Pal15, dMMI14]. Policy-Related [Ell00]. Policy-Relevant [Gun20]. Political [Bal96, Mer23, Tuc07, Tur23]. Politicizing [JV23]. Politics [Cle22, Dam22, Goe11, Mer23, Moy03, Nel23, Ris00, SL23, SV23, Ueb20]. Polymorphism [Gra01]. Popper [Jha06a, Kra13, Nan11]. Population [Lóp17, Mar20, Rob17, SM20, Ski04, SD17, dC20]. Population-Making [Lóp17]. Populations [CHB22, DS17, Gue17, SDGDV17, VD17, dC17]. Porcupine [HO96]. Porta [Rus17]. Portuguese [Roq22]. posc_{a}_{0}0337[RB22].Positivism[Fuc93, Ueb13].Positron[Pas12].Possibility[FH07a].Possible[Cam93, SP23].Post[Mag11, Omo17, Wag08].Post − Copernican[Omo17].Post − Genomic[Mag11].Post − Structural[Wag08].Postwar[Hyu22].Potentia[Duc98].Power[Che00, Geo17, Gol13, Smi09].Powerless[Lan14].Practice[ASOVS15, Das16, HS17, La 04, MMMN19, Rei19, Rol99, SP23, Wyl96, You16, Hug06].Practice − Oriented[ASOVS15].Practices[Kus19b, MGY13, SDGDV17, vD23].Practicing[Rob17].Praeceptor[VFN22].Pragmatic[Sol95].Pragmatics[Des19].Pragmatist[Pih08].Pre[Wat98].Pre − established[Wat98].Precarious[Bur02].Prediction[Bru93, McI01, Pas12].Predictions[Mar22].Preformation[Det03].Prehistory[Gio17].Presence[Roj20].Present[Dre23, Loo02, Sch01, Sch02].Preserve[Cit93].Prestige[Nor09].Primary[Swe12, Swe13].Primitive[Ink17].Principal[Hla18].Principle[Del93, Fle13, Hut16, Lan14].Principles[Ham02, Kro21, Obe14, Di 11].Print[Dup08, LR94].Prinzip[Moy03].Priori}[Ric02b].Priority[Han06].Prism[TPDB21].Privacy[Zah17].Prize[Kor94].Probability[Cam09b, Syl98].Problem[Alo24, Art98, Bok14, BH17, Bul16, Gol16b, Gra01, Hat00, Hei07, Hoo18, Lüt00, Mik97, Sab98, Smi09, SC19, Wüt18, Sab00, Sal00, Sta12, Réd14].Problem − Solving[Hoo18].Problematic[VD17].Problems[Sch20].Procedures[Kar17].Process[BA11, Bur08, Höf21, RS08, Vic00].Processes[Nis16, Nur13].Processing[IJ16].Proclus[Gou22].Production[Spe04, Ueb08].Productivity[Ano05a, Kle05b].Products[Kev15].Products − of − Nature[Kev15].Professor[Swe13].Program[Gie12, Ric20, Hun19].Progress[Bau03, Bor19, Bro16b, Duc10, FG12].Prohibitorum[Rei97].Project[Fee07, Bal96].Projections[PH21].Promises[Li 23].Proper[Clo08].Property[Jac15a].Propositional[Abr14, CF15].Protestantism[Jal12].Proust[Pia24].Psychiatry[Vin19].Psychic[Pia24].Psychobiology[Kin93].Psychological[Vin24].Psychology[Cat07, Fee07, GY13, Hat97, Nur13, Poj11, Sch03, Vin19, Zol97, Tay95].Psychophysical[Lam03].Psychophysics[Gio17].Ptolemy[Bow07, Ham16, Sha02].Public[Dur10, Gol16b, MB24, MC22, dMMI14].Publicization[Bro23].Puerto[Mol17].Pure[Gra01].Purity[Dör01].Purposes[Sch02].pursuit[Fra93a].Putting[Jac08].Pythagorean[Ban06, Kon18].

quadripartita [Rou18]. Quantified [Han22]. Quantum [DG95, MR14, Réd14, Wüt18]. quartet [Hug06]. Quechua [Mol17]. Question [Jes98, Wes13]. Questions [Duc10, Fra93b, KC96, Nor09]. Quid [Slo09]. Quis [AO22].

R [Bar02b]. Race [Mag11, Roq22, Smi12]. Racial [Cle22, Hyu22, Lef22]. Racializing [Bar22a]. Radar [Bul16]. Radiobiogical [SF00]. radiolarian [Le 93]. Raman [Das16]. Random [Yea16]. Range [Kok19]. Rational [Bou16, Slo09, Wra97]. Rationale [Ink17]. Rationality [Joy95]. Raymond [Sim22]. Rays [Gol13]. Re [Dre23]. Re-Examined [Dre23]. Reading [Dij08, Pic01, Vin24]. Readings [Wag08]. Real [Koc11, Lun18, dBV13]. Realism [BG98, Fel17, Gal02, Hei07, Mag11, Pas12, dBV13]. Realist [How93]. Reality [Ach94, Jes98, Kha24]. Really [How93, Lan14]. Reappraisal [BG98, Gra10]. Reason [Ric02b]. Reasoning [Cor21, Daw16, Goo06]. Reassessment [Pie06, Str23]. Received [Ano93a, Ano93b, Ano94a, Ano95a, Ano95b, Ano96a, Ano96b, Ano97a, Ano97b, Ano98a, Ano99, Ano01a, Ano02, Ano03a]. Reception [Dup08, Mag03, Sch03, Ueb08, Wat98, Bro95]. Receptions [Des95]. Reclaiming [KC96]. Reconsidered [Lug22, Sol03]. Reconsidering [Pit01b, Tso03]. Reconstructing [Stu97]. Reconstruction [Mir22]. Rectilinear [Gor05]. Reductionism [Bla18]. Refashioning [Dij08]. Referees [Ano96k]. Reference [Jac10]. Referent [Cam93]. Reflections [Duc10, For10, Gal99, Gar01]. Reflective [Che00]. Reflex [Chi20]. Reflexivity [Mäk99]. Reform [Bae24, dC22b]. Reframing [Gol16b]. Refuse [SG01]. Refutation [Stu14]. Regime [Toe21, Pic95, Ano95c]. Regimes [Li 21]. Regiomontanus [Sha02]. Régis [Des03]. Regius [Str23]. Regnault [Sto24]. Regress [Zup17]. Regressus [Mil18]. Regulating [BERW16, Mét15]. Regulative [Kit99]. Rehabilitation [Fre05]. reinforcing [JR23]. Reintrepretation [Pal04]. Reinvention [BA11]. Reinvigorating [PB23]. Reject [Hud16]. Rejection [Lug22]. Related [Ell00, Hei12, Luo13]. Relation [Bur08, Cat07, Cav06, Gra10, Ham02, Kin05, Mát06]. Relationism [Slo09]. Relations [GY13]. Relationship [Bur02, Rea13]. Relativism [Lyn19]. Relevance [Car17]. Relevant [Gun20]. Reliability [Car17, Cha95]. Religion [Har96, Lan12]. Reluctant [Che14]. Remark [Gal02]. Remarks [Sch18]. Remembering [CK13]. Reminiscence [Buc10]. Renaissance [CE04, Ros22]. Repatriation [Wyl96]. Replicating [Bia22]. Replication [Sha16]. Reply [Hei13, Kou12, Len95, Pic05b, Rhe05b, Sab00, Swe13, Wes13]. Representation [DR18, Dow09, Eig20, Ham02, Len15]. Representational [Tal21, Zup17]. Representations [Jeb18, Kar17]. Representative [Eig20]. Representing [All16, KBG22, Kar17]. Reproductive [Cla93b]. Research [Dor00, Ell16, FH07a, FH07b, Fer23, Gra01, Ham01, HS17, Hof01, HS01, Kar17, Kos23, Lef22, Liu16, LWR20, Ner19, Nuk19, OS18, PCC^{+}19, Res00, Ric20, SJ18, Spe04, Twe06b, Whi99, dC22a, Kyl22]. Resistance [Kha24, Sol10]. resolutae [Cha02]. Resolution [Fra95, Han06]. Resolved [Bur01]. Resonance [LL95]. Resources [Omo22]. Response [Ano09c, Bue19, Des00, Hat00, Hei10, New09a, New09b, Pea17, Pic05c, Rup06]. Responses [Sol08]. Responsible [Kou12, Sol12]. Results [Fra95, Zup17]. Rethinking [Bok01, Di 11, Rus17, Wyl99]. Review [Cla93a, Des95, Fra93b, Gie94, Gre93b, Gre97, Har96, Har97, Juk17, Lev96, Ram98, Ric96, Row97, Sch95, She94, Vic95]. Reviews [Boh12]. Revisited [Ari18, Lew07, Poj11, SV23, Wes13]. Revisiting [Joh05]. revival [Bru93]. Revolution [FH07b, Gar01, Gar04, Har94b, Har94a, Jac01, Jac21b, Kin10, Obe14, Ueb08]. Revolutionary [Mag03, GH05, Hor05]. Revolutions [Kuh70, Kuh62, Kuh96, KH12, Kin05]. RG [Bat19]. Rheticus [VFN22]. Rhetoric [Har09, McA16, Nel23, Whi99]. Rican [Mol17]. Richard [Die96, Rei19]. Riddles [Tam21b]. Riehl [Hei07]. Rise [Boh12, CP99, Glo95]. Risk [Bid16, Ded20, Eig20, Ste16a]. Risks [BR20b]. Road [HO96]. Robert [Hei12, Hen19, Lew06]. Roberval [Bab22]. Robotics [Sch22, Tam21a]. Robotics-Inspired [Tam21a]. Robots [Tam21a, vD23]. Rocket [Lan19]. Roger [Ako22, Col18, KH17, Péc18]. Role [Ber04, Car17, Dup08, Gar95b, Gau00, Kro21, Maj14, Mla07, RS08, You16]. Roles [Bou16]. Rolin [Kou12]. Roman [Hei15]. Romantic [Mar13]. Romanticism [Lev96]. Roots [Gor05]. Rotten [De 06]. Rouse [Koc11, dBV13]. Royal [Hen19, Kus19a, Kus19b, Rei19]. Ruins [Vis01]. Rule [Nuk19]. Rulers [Hut96]. Rules [Bar15]. Russell [Hei10, Obe14]. Rust [Jac21b].

S [CK13]. Sabra [Sal00]. Saliba [Sab00]. Salience [Nor21]. Sally [Ric20]. Santesmases [RB22]. São [dC20]. Scaffolding [Kok19]. Scaliger [Bla08]. Scarcity [Cur21]. Scenarios [PH21]. Scene [Pea17]. Scenographic [Van11]. Scheiner [Gor96]. Schelling [FG12, Whi19]. Schlichting [Beg20]. Schlick [Hei07]. Schliemann [SG01]. Schmitt [Tur23]. Scholarly [BA11]. Scholars [Cla93a]. Scholarship [Duc10, Loo02, Ric96]. Scholastic [Pet18]. Scholasticism [Sch18]. School [Gio17]. Schrader [Ric20]. Sci [Sab00, Sal00]. Science [ACT22, Ano97c, Ano09c, AS12, BM04, Bar01, Bar22b, Bau03, BR20b, Bie04, Bor19, Bur01, Bur02, Bur99, CE04, Car17, Car99, Cat07, CD19, Cha06, Clo08, CBRG17, CSB^{+}23, Cop18, Das16, Dor00, Dou16, Dow09, Dup12, Dur10, Ell00, FH07a, FH07b, Fee07, Fer23, FGS21, For10, Fuc93, Gal09, Gar95b, Gie12, Gin16, GG11, Glo95, Gol16b, HO96, HE22, Han06, Har96, Har94a, Har09, Har97, Hat97, Hof18, Hol12, Hoo18, Hor05, Hud16, Hul98, Jac01, Jac10, JZ98, Jan04, Jan02, Jha06b, Kai96, Kin05, Kin10, Kin93, KC96, Koc22, Kos23, Kos04, Kou12, La 99, Lef05, Liu16, LWR20, Lop09, LR94, Lun18, Lyn24, Mag03, Mäk99, MJLG23, McA16, Mét15, Miz20, Mla07, MB24, Moy03, Nas18, Nor99, Nor09, Nov20, Omo17, Ore10]. Science [ON17, Pih08, Pin15, Pin20, Pit01a, PCC^{+}19, Rem04, Res98, Ric96, RH03a, Ric08, RB22, RC15, Rol12, Roq22, Sch20, Sch03, SL23, SV23, Sha95, SP23, Sim19, Sol95, Sol96, Sol10, Sol12, Spe15, Ste02, SB02, Stu14, Tam21b, Tuc07, Tur23, Van15, Vic00, Wal01, Wil21, Wyl99, You17, Zam06, Zik01, dO17, PB23, RH03b, Rup06, Sab00, Sal00, Vic95, Ano97d, Rem04, Lan12].

Sciences [Alv19, Cam16, Hyd03, Kra13, Lev96, Li 21, Li 23, LWR20, Man13, Mar13, MGY13, Ner19, Pie06, Rio13, Rol99, Sch06, Seg22, Vig22, GA98]. Scientific [And06, BJ02, BR20b, BA11, Bou16, BH17, Bur99, CM23, Cha21, CW18, Dam22, Das16, Daw16, Ded20, DS07, Duc10, Eck16, Gar01, Gar04, Ger13, Gol13, Gra04, IJ16, Jac01, Kin05, Kos19a, Kuh70, KM15, La 04, Lin19, Luo13, Nor02, PB23, PH21, Rei19, Ric97, Sch15, SE23, Šeš21, Spe04, Ste11, Van15, Vic95, Vic00, You16, Zik01, Zol97, dC22b, dMMI14, dRP14, GH05, Kuh62, Kuh96, KH12, Sis95, Sup93]. Scientific/Scholarly [BA11]. Scientifically [Hul98]. Scientist [BJ02, HS17, Sen00, SGMB18]. Scientists [Che94, Gun20]. Scopes [Kon14]. Screenshots [All16]. Search [Osl11]. Second [Ste16a, Ano97c, Pie93, Pie94]. Second-Order [Ste16a]. Secondary [Hat00, HS01]. Secrecy [Kai96]. Secret [Psi04]. Secularization [Sim19]. See [Che00]. Seeing [Bro14]. Select [Mét15]. Selection [Lyn19, Nuk19]. Selenographia [Tau22]. Self [FH07a, FH07b, Gig12, JR23]. Self-Delusion [Gig12]. Self-Directed [FH07a, FH07b]. Self-reinforcing [JR23]. Semantics [Har94a, MG23, Pur18, Zup17]. Sennert [Bla11]. Sensation [Gio17]. Sense [TPDB21]. Sensibility [Nov20]. Sensing [Cha23]. Sentiments [San24]. Sequence [Ste11]. Sequences [Ste11]. Series [Jal16]. Set [Cam09b]. Seventeenth [Gar95b, GG11, Gre93a, Lüt00, Sha03, Zik01]. Seventeenth-Century [Gar95b, Gre93a, Lüt00]. Severe [ACT22, Sch22]. Sex [Cla93b]. Shades [Des19]. Shared [Car11, Fel17]. Shedding [Das08]. Shift [Duc98, Sch22]. Should [Hud16]. Shoulders [Kou16]. SI [dC22b]. Significance [CSB^{+}23, KH17, McD03, Wit16, dC22b]. Signs [Gag08, Mar22]. Silks [GMHP21]. Sils [Kon14]. Sils-Maria [Kon14]. Simon [Aug00, Sen00]. Simple [Hol12]. Simplicius [Bow02]. Simply [Kor94]. Simulation [Gre21, Len15, McD03, Par14, Yli14, dRP14]. Simulations [Jeb18, SE23]. Situating [Kel94]. Sixteenth [BG98, Bla18, Jal12]. Sixteenth-Century [Bla18]. Skepticism [Stu14]. sketch [Tay95]. Skill [You16]. Skills [You17]. Sky [vdS12]. sleep [Pia24]. Slips [dR24]. Smith [Gol13]. Social [Eck16, Fag11, Gis09, Juk17, KM15, LWR20, Lon08, Lop09, Luo13, Man13, Mar13, MGY13, Pic05a, Rio13, Sch03, Sha95, Sol10, Wyl08]. Social-Scientific [Luo13]. Socialism [Har97]. Socially [Kou12, Sol12]. Societal [Kos23]. Society [Har96, Hen19, Kus19a, Kus19b, Rei19]. socioeconomic [Tay95]. Sociological [Yli14]. Sociologists [Coo04]. Sociology [Blo05, Gis02, Pic05a, Rhe05b, Vic95]. Software [Spe15, WH20]. sole [Bar22b]. Solomon [Clo08, Kou12, Lon08, Ric08, Wyl08]. Solving [Hoo18, Sab98, Sab00, Sal00]. Some [Duc10, Gar01, Gre93a, Ham01, Kai96, KC96, Llo21, Nas18, Urr10, Vic95, Gut21]. Soul [AM19]. Sources [Swe12, Swe13]. South [Ric20]. Space [FH07a, Hei07, JR23, Slo09, CD19]. Special [Ano05a, BR20a, RH03b, RB22]. Speculation [Daw16]. Sphairopoiia [Ham16]. Spider [GMHP21]. Spinoza [And19]. Spiritual [Gar12, Jal12]. Spiritualism [Hac24]. Spiritualist [AM19]. Splitting [Pia24]. Stability [Mik97]. Stahl [De 04]. Stakes [Whi19]. Standard [Way96]. Standardization [Jac21a]. Standardized [Ell16]. Standardizing [Cle22]. Standing [Kou16]. Standpoint [Ryc03]. standpoints [Sis95]. Star [Met97]. Starry [Gra00]. State [Hor05, Pal15]. States [Pal15]. Statisticians [Rud01]. Statistics [Sta99]. Stebbing [Tub20]. Stem [Fag11, Liu16]. Stepping [Cur21]. Stewardship [Wyl96]. Still [Dor00]. Stimulus [Gio17]. Stink [Wüt18]. Stoic [Jal12]. Stories [Jan02]. Strang [Pan15]. Strangeness [Bor21]. Strategies [Nor21, OS18, Rud98]. String [RC15]. Strings [Fri21]. Structural [Kai96, Sie14, Ste16a, Wag08]. Structure [Kin05, Kos19a, Kuh70, Ryc03, dO17, Kuh62, Kuh96, KH12, Pic01]. Structuring [Spe04]. Struggle [Kin93]. STS [Lun18]. Students [Bia22]. Studies [Bau03, Bur01, HO96, Jha06b, Luo13, Ore10, Pit01a, Sha95, SM20, Tuc07, Zam06]. Study [ASOVS15, Cam09b, Chv20, CHB22, Ell16, Har05b, Hul98, Hut96, Ink17, Jos05, Kle05a, Miz20, Par14, Pas12, PCC^{+}19, Tam21a, Tan19, Zol97]. Studying [Hul98]. Sturm [Har94b]. styles [Vic95]. Subconscious [Pia24]. Subject [Jac08]. Subjective [Ach01]. Subjectivity [Lin19]. Subjects [Nuk19]. Subordinate [Bla11]. Substantivalism [Slo09]. Substitution [Nee96]. Subvert [Hla18]. Summa [Rou18]. Supernaturalism [Det03]. Surplus [Bor19, Ryc03]. Surprises [Urr10]. Surveyors [Lóp17]. Survived [Jac21a]. Susan [Tub20]. Sustainability [Goe11]. Swerdlow [Wes13]. Switching [Cat07]. Sylva [Rus17]. sylvarum [Rus17]. Symbolic [Lup14]. Sympathy [San24]. Symposium [BM04]. Syntax [Har94a]. Syntheses [Boh12]. Synthetic [Mét15, Pic05a]. System [KH17, Tes22, dC22b]. Systematic [Boh12]. Systems [AB23, Alv19, Bul16, Rea13, Tuu01]. Szabó [Mát06].

Tables [Ber04, Cha02]. Tabulae [Cha02]. Tacitus [Gig12]. Taiwan [Liu16]. Tale [JZ98]. tales [Sch95]. Targeting [Nas18]. Taught [VFN22]. Taxonomic [Lüt00]. Taxonomy [Smi09]. Teaching [Bur99, Roj20]. Technique [You16]. Technological [AB23, Kle05a]. Technologies [Lóp17, MC22, SJ18, Spe04, Hon98]. Technology [ACT22, Ano09c, BM04, Har97, Joh05, LR94, You16]. Technomoral [NKvdP22]. Technoscience [Ano05a, Bar05, Har05b, Kle05c, Pic05c]. Technoscientific [Kle05b]. Telegraphs [Kel94]. Telescope [Dup08, Zik01]. Tell [CD19]. Temporal [Che05]. Tennemann [Man19]. Terminologies [Mar13]. Terrestrial [Jos05]. Tertium [Slo09]. Test [Mäk99]. Text [Wag08].

Textbooks [Bro23, Dij08]. Texts [And06]. Thagard [Stu14]. Their [Gun20, Hon98, Kai96, Zup17, OS18, Smi18]. Them [Ach07, Fer23]. Themes [Bro23, Gre93a, Sab98, Sab00, Sal00, Sha02]. Theodosius [dC20]. Theology [Met97, Tur23]. Theon [Sha02]. Theorems [For10]. Theoretical [Bar15, Cra15, Duc98, Hug06, Riv19, Way96, Wüt18]. Theories [Bes15, BA11, Dij08, Kok19, Lan14, Sch03]. Theory [Bow02, Bow07, Cha23, Gau00, GM21, Goo06, GY13, Hof18, Kra97, Mag03, Mar96, McD03, Mir22, Nor23, Pal04, Ped08, Pic05a, Rea13, Réd14, RS08, RC15, Sch14, Sha96, Tuc07, Vin24, Zam06, Bru93]. Theory-Change [Mir22]. Theory-Choice [Sch14]. There [Dor00, Lan19, Mon17, Sch21]. therein [Duc10]. Thermodynamics [Wri16]. These [Gie94, Pea17]. Thesis [Fre05]. Theurgy [Gou22]. Things [Blo05, Rhe05b]. Thinking [Gis09, Kis06]. Thomas [Buc10, Jac10, Lev10, MG23, Pit01b, dO17]. Thought [Bok01, Bro14, Cam07, Elg14, FS14b, Fis03a, Hei15, Hop14, Stu14, Tha14]. Thoughts [BJ02, Sch11]. Threat [Ano09c]. Three [Cor21]. Thrombin [Mar96]. Tics [dR24]. Tides [Ako22, Pal04]. Tildesley [Cle22]. Time [Goe11]. Times [Gie94]. Tongue [dR24]. Tool [dR14]. Tools [PH21]. Topology [MR14]. Traces [Loo02]. Tractarian [Lam03]. Tractatus [Ham02, Nor02]. Trade [BC99]. Tradescant [Pan15]. Trading [BN99]. Tradition [Bla94, Mas10]. Traditional [vdS12]. Training [PT19]. Traité [Str23]. Trajectories [Hon98]. Trajectory [SvJB02]. Trans [Sch02, Wyl96]. Trans-Contextual [Sch02]. transcend [Jha06a]. Transcendental [Ryc03]. Transfer [PCC^{+}19]. Transformation [Cam07, Van11]. Transformations [Hat97]. Transformative [Gol13]. Transforming [Che05, Tuu01]. Transitions [Nis18]. Translations [Swe12, Swe13]. Transnational [CHB22, Kyl22, Roq22]. Transnationalism [dC22a]. Transparent [Hol12]. Transwar [CHB22]. Trebizond [Sha02]. Trends [Rem04]. trial [Gor96, Kon14]. Tropes [Leh03]. True [Whi19]. Trust [Bou16, Fel17, Lug22]. Truth [GM21, Joy95]. Tuning [Sta12]. Turn [Mer23, Sol95, Tam21a]. Twentieth [All97, YB16]. Twentieth-Century [All97, YB16]. Twenty [Sha08, Ric02b, Tam21b]. Twenty-First [Ric02b, Tam21b]. Twenty-Nine [Sha08]. Two [Pie94, PT19, Tal21, Wil21, Yea16]. Types [All01]. Typographic [LR94].

U.S. [Gol13, Kev15]. Uncertain [Jac21a]. Uncertainties [OA22]. Uncertainty [ACT22, Han22, MC22, NKvdP22, Sch22, Ste16a, TV22]. Unconscious [Alo24, Vin24]. Underdetermination [Cur21, Die93, DS07]. Undermining [dMMI14]. Understand [Coo04]. Understanding [Bro14, Dup04, Kos19a, Mey18, MB24, Par14, Twe06b, Wüt18, YB16, Yli14, dRP14, dR14].

Understandings [Jes10]. Unexpected [Lun18]. Unfaithful [Hon98]. Unification [Kit99]. Unified [Dow09]. Uniform [Jac21a]. Uniformitarianism [Dre23]. Uniformity [Di 11]. Unifying [Hof18, Kos19b]. Unit [Jac21a]. United [Pal15]. Units [dC22b]. Unity [Nor99, Smi09, Sol12, Stu97, Way96, Wyl99]. Universal [Van11]. Universality [Bat19]. Universals [Hat16]. Universe [Kra97, Sab98, Sab00, Sal00, Gra10]. University [Blu93, Kin93, Rai93, dC20]. Unknowability [SGD17]. unscientific [Bur12]. Unusual [Cop18, DS17]. upon [Goe11, Poj11]. Urge [Hut96]. Use [Che94, Lug22, Rus17, Sta12, Str23]. Uses [Che00, GG11, Smi18, Wil21]. Using [Bur01, Sch02, Sha16, Tay95]. Utopias [Van11].

V [Das16, Gol13]. Vaccine [Gol16b]. Vacuum [Bla08, Nej19]. Valerian [Nej19]. Validity [Pal10]. Value [Ell16, Gun20, Hud16, Kar17, SF00, SGMB18, Whi99, Rup06]. Value-Bifurcation [Whi99]. Value-Free [Gun20, Hud16, SGMB18]. value-neutrality [Rup06]. Values [Bur99, Clo08, Dou16, Ger13, Ric02a, Rol12, Sol12]. Varignon [Mag03]. Variously [KL21]. Vectors [Wyl08]. Venus [Bar22b]. Verborum [Rei97]. Vermittler [Mar09]. Vernacular [Shi13]. Versuch [Mar09]. versus [Kra97, Mar09, Sab00, Sal00]. Vertebrate [Rai93]. very [Hei10]. via [Šeš21]. Vicissitudes [AO22]. Vickers [New09b]. Victims [AB23]. Victor [AM19, Die96, dR24]. Victorian [Ric96]. Vienna [Ueb20, Ueb08]. View [Gur06, MR14, Nis16, Rud01]. Views [Ach01, Gun20, SGMB18]. Vincenzo [Kon18]. Violin [Das16]. Violin-Player [Das16]. Virchow [SG01]. Viret [Gar12]. Virtual [Len15]. Virtues [Bur01]. visa [Bar22b]. Visible [Kus19b]. Vision [Car11]. Visitors [De 06]. Visual [Goo06, Hen19, Jeb18, Kus19a, Kus19b, Rei19, Van11]. Visualization [dRP14, dR14]. Void [Nej19]. Vole [Cop18].

W [Cob12]. Waller [Rei19]. Wanted [Rou18]. War [AB23, Ano95c, Pic95, Wüt18, Wit16]. Warrants [Kok19]. Was [Goe11, How93, Bar21, Gar21]. Waste [HS01]. Wave [Ped08]. Waves [Fra10]. Wax [De 06]. Way [Kou16, SP23, Rup06]. Ways [AS12, Fer23, Sie14]. Weasel [Cop18]. Weather [Mar22, Par14]. Weber [Fra10]. Webs [GMHP21]. Well [Pin15]. Well-Ordered [Pin15]. Western [Gra00]. Westman [Hei12, Hei13, Swe13]. Weyl [Ryc03]. where [Hei10]. Whewell [Duc10, Ric96, Sny97, Tan19]. Whiptail [Sol03]. Who [AB23, Gar21, Sch21, dMMI14]. Wicked [Sch20]. Wilhelmine [Moy03]. William [Gol16a, Koc21, Lug22, Ric96, Tan19]. Winds [Mar22]. Wing [Dam22]. Witch [De 04]. without [Fel17, Ric08]. Wittgenstein [Ham01, Ham02, Nor02]. Wolff [Cam16]. Woman [De 06]. Women [Bar20, Mar20, RB22, San20]. Words [CD19]. Work [Bro18a, Fra10, Gin16, Gra00, Gre00, Nis18, Ram98, Van11, Kin05]. Working [Wyl99]. Workplace [Bar20]. Works [Mat20]. World [Ano95c, Cav06, Dam22, Hen19, KBG22, Roq22, Ryc03, Pic95]. World-Conception [Dam22]. Worm [De 04]. Write [Pih08]. Writing [Ueb08]. Wrong [Sta99]. Wrote [Rou18]. Wundt [Hat97].

X [Gol13]. X-Rays [Gol13].

Years [Sch11, Sha08]. Yes [Kev15]. Yourself [Lug22].

Zone [BN99]. Zoophysiological [Lyn24]. Zur [Jac08].

References

Aloisi:2024:IMB

[AAM24] Alessandra Aloisi and Delphine Antoine-Mahut. Introduction: Maine de Biran and the afterlives of Biranism. Perspectives on Science, 32(1):1–14, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/32/1/1/2210162/posc_e_
00595.pdf.

Albornoz:2023:HDT

[AB23] María Belén Albornoz and Javier Andrés Jiménez Becerra. How do technological systems define who war victims are? Perspectives on Science, 31(1):186–205, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
1/186/2073149/posc_a_00586.pdf.

Abrusci:2014:HAP

[Abr14] V. Michele Abrusci. On Hilbert’s axiomatics of propositional logic. Perspectives on Science, 22(1):115–132, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.1.abrusci.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v022/22.1.abrusci.pdf.

Achinstein:1994:JPM

[Ach94] Peter Achinstein. Jean Perrin and molecular reality. Perspectives on Science, 2(4):396–427, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Achinstein:2001:SVK

[Ach01] Peter Achinstein. Subjective views of Kuhn. Perspectives on Science, 9(4):423–432, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v009/9.
4achinstein.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.4achinstein.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361401760375811.

Achinstein:2007:AEE

[Ach07] Peter Achinstein. Atom’s empirical eve: Methodological disputes and how to evaluate them. Perspectives on Science, 15(3): 359–390, Fall 2007. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v015/15.3achinstein.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v015/15.
3achinstein.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2007.15.3.359.

Andreoletti:2022:ISU

[ACT22] Mattia Andreoletti, Daniele Chiffi, and Behnam Taebi. Introduction: Severe uncertainty in science, medicine, and technology. Perspectives on Science, 30(2):201–209, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
2/201/2005010/posc_e_00411.pdf.

Akopyan:2022:DTB

[Ako22] Ovanes Akopyan. Discussing tides before and after Newton: Roger Joseph Boscovich’s De aestu maris. Perspectives on Science, 30(6):1042–1064, November/December 2022. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/30/6/1042/
2059520/posc_a_00567.pdf.

Alexandrescu:2007:DPE

[Ale07] Vlad Alexandrescu. Descartes and Pascal on the Eucharist. Perspectives on Science, 15(4):434–449, Winter 2007. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.4alexandrescu.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v015/15.4alexandrescu.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2007.15.4.434.

Allchin:1997:TCP

[All97] Douglas Allchin. A Twentieth-Century phlogiston: Constructing error and differentiating domains. Perspectives on Science, 5(1):81– 127, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://search.ebscohost.com/login.
aspx?direct=true&db=aph&AN=6664445&site=ehost-live.

Allchin:2001:ET

[All01] Douglas Allchin. Error types. Perspectives on Science, 9 (1):38–58, Spring 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v009/9.1allchin.html;
http://muse.jhu.edu/journals/perspectives_on_science/v009/
9.1allchin.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/10636140152947786.

Allen:2016:RCA

[All16] Matthew Allen. Representing computer-aided design: Screenshots and the interactive computer circa 1960. Perspectives on Science, 24(6):637–668, November/December 2016. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/635021;   http://muse.
jhu.edu/article/635021/pdf.

Aloisi:2024:CIF

[Alo24] Alessandra Aloisi. Coenesthesia or the immediate feeling of existence: Maine de Biran and the problem of the unconscious between physiology and philosophy. Perspectives on Science, 32(1): 47–69, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/32/1/47/2210164/posc_a_00598.pdf.

Alvargonzalez:2019:SS

[Alv19] David Alvargonzález. Sciences as systems. Perspectives on Science, 27(6):839–860, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00327.

Antoine-Mahut:2019:EMS

[AM19] Delphine Antoine-Mahut. Experimental method and the spiritualist soul: The case of Victor Cousin. Perspectives on Science, 27(5):680–703, September/October 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00321.

Antoine-Mahut:2019:IDE

[AMM19] Delphine Antoine-Mahut and Silvia Manzo. Introduction: Debates on experience and empiricism in Nineteenth Century France. Perspectives on Science, 27(5):643–654, September/October 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_e_00319.

Anderson:2006:CSM

[And06] Ronald Anderson. The crafting of scientific meaning and identity: Exploring the performative dimensions of Michael Faraday’s texts. Perspectives on Science, 14(1):7–39, Spring 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.1anderson.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v014/14.1anderson.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.

2006.14.1.7. Special Issue: Cognitive Psychology and Faraday’s Experiments.

Andrault:2019:SMP

[And19] Raphaële Andrault. Spinoza’s missing physiology. Perspectives on Science, 27(2):214–243, March/April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00306.

Anonymous:1993:BRa

[Ano93a] Anonymous. Books received. Perspectives on Science, 1(1):147– 151, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1993:BRb

[Ano93b] Anonymous. Books received. Perspectives on Science, 1(3):560– 563, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1993:GCa

[Ano93c] Anonymous. Guidelines for contributors. Perspectives on Science, 1(1):??, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1993:GCb

[Ano93d] Anonymous. Guidelines for contributors. Perspectives on Science, 1(2):??, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1993:GCc

[Ano93e] Anonymous. Guidelines for contributors. Perspectives on Science, 1(3):??, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1993:GCd

[Ano93f] Anonymous. Guidelines for contributors. Perspectives on Science, 1(4):??, Winter 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1993:NC

[Ano93g] Anonymous. Notes on contributors. Perspectives on Science, 1 (1):??, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:BR

[Ano94a] Anonymous. Books received. Perspectives on Science, 2(3):350– 351, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:EOL

[Ano94b] Anonymous. Errata: The Origin of the Liquid-Drop Model and the Interpretation of Nuclear Fission. Perspectives on Science, 2 (2):254, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Correction of printers error in bottom three equations from page 90.

Anonymous:1994:GCa

[Ano94c] Anonymous. Guidelines for contributors. Perspectives on Science, 2(1):??, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:GCb

[Ano94d] Anonymous. Guidelines for contributors. Perspectives on Science, 2(2):??, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:GCc

[Ano94e] Anonymous. Guidelines for contributors. Perspectives on Science, 2 (3):353–354, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:GCd

[Ano94f] Anonymous. Guidelines for contributors. Perspectives on Science, 2(4):488–489, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:NCa

[Ano94g] Anonymous. Notes on contributors. Perspectives on Science, 2 (1):??, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:NCb

[Ano94h] Anonymous. Notes on contributors. Perspectives on Science, 2 (2):??, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:NCc

[Ano94i] Anonymous. Notes on contributors. Perspectives on Science, 2(3): 355–356, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1994:NCd

[Ano94j] Anonymous. Notes on contributors. Perspectives on Science, 2(4): 490–491, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:BRa

[Ano95a] Anonymous. Books received. Perspectives on Science, 3(1):146– 151, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:BRb

[Ano95b] Anonymous. Books received. Perspectives on Science, 3(3):421– 424, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:ECH

[Ano95c] Anonymous. Errata: Cyborg history and the World War II regime. Perspectives on Science, 3(2):??, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). See [Pic95].

Anonymous:1995:GCa

[Ano95d] Anonymous. Guidelines for contributors. Perspectives on Science, 3(1):??, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:GCb

[Ano95e] Anonymous. Guidelines for contributors. Perspectives on Science, 3(2):??, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:GCc

[Ano95f] Anonymous. Guidelines for contributors. Perspectives on Science, 3(3):??, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:GCd

[Ano95g] Anonymous. Guidelines for contributors. Perspectives on Science, 3(4):582–583, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi.

Anonymous:1995:NCa

[Ano95h] Anonymous. Notes on contributors. Perspectives on Science, 3 (1):??, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:NCb

[Ano95i] Anonymous. Notes on contributors. Perspectives on Science, 3 (2):??, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:NCc

[Ano95j] Anonymous. Notes on contributors. Perspectives on Science, 3(3): ??, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1995:NCd

[Ano95k] Anonymous. Notes on contributors. Perspectives on Science, 3(4): 584–585, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi.

Anonymous:1996:BRa

[Ano96a] Anonymous. Books received. Perspectives on Science, 4(1):141– 144, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:BRb

[Ano96b] Anonymous. Books received. Perspectives on Science, 4(3):387– 389, Fall 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:GCa

[Ano96c] Anonymous. Guidelines for contributors. Perspectives on Science, 4(1):??, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:GCb

[Ano96d] Anonymous. Guidelines for contributors. Perspectives on Science, 4(2):??, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:GCc

[Ano96e] Anonymous. Guidelines for contributors. Perspectives on Science, 4(3):??, Fall 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:GCd

[Ano96f] Anonymous. Guidelines for contributors. Perspectives on Science, 4(4):??, Winter 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:NCa

[Ano96g] Anonymous. Notes on contributors. Perspectives on Science, 4 (1):??, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:NCb

[Ano96h] Anonymous. Notes on contributors. Perspectives on Science, 4 (2):??, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:NCc

[Ano96i] Anonymous. Notes on contributors. Perspectives on Science, 4 (1):??, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:NCd

[Ano96j] Anonymous. Notes on contributors. Perspectives on Science, 4 (4):??, Winter 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1996:R

[Ano96k] Anonymous. Referees. Perspectives on Science, 4(1):??, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:BRa

[Ano97a] Anonymous. Books received. Perspectives on Science, 5(1):152– 154, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:BRb

[Ano97b] Anonymous. Books received. Perspectives on Science, 5(4):571– 577, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:CPS

[Ano97c] Anonymous. Call for papers: Second International History of Philosophy of Science Conference. Perspectives on Science, 5(2):282– 283, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:FAC

[Ano97d] Anonymous, editor. First Annual Conference for the History of Philosophy of Science. University of Chicago Press, Chicago, IL, USA and London, UK, 1997. ISSN 1063-6145 (print), 1530-9274 (electronic). i–iv and 285–480 pp. Selected papers from the conference (HOPOS) held in Roanoke, VA, April 19–21, 1996, Perspect. Sci. 5 (1997), no. 3.

Anonymous:1997:GCa

[Ano97e] Anonymous. Guidelines for contributors. Perspectives on Science, 5(1):155–157, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:GCb

[Ano97f] Anonymous. Guidelines for contributors. Perspectives on Science, 5(2):??, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:GCc

[Ano97g] Anonymous. Guidelines for contributors. Perspectives on Science, 5 (3):??, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special issue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Anonymous:1997:GCd

[Ano97h] Anonymous. Guidelines for contributors. Perspectives on Science, 5(4):??, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:NCa

[Ano97i] Anonymous. Notes on contributors. Perspectives on Science, 5 (1):158, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:NCb

[Ano97j] Anonymous. Notes on contributors. Perspectives on Science, 5 (2):??, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1997:NCc

[Ano97k] Anonymous. Notes on contributors. Perspectives on Science, 5(3): ??, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special issue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Anonymous:1997:NCd

[Ano97l] Anonymous. Notes on contributors. Perspectives on Science, 5 (4):??, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:BR

[Ano98a] Anonymous. Books received. Perspectives on Science, 6(3):331– 334, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:GCa

[Ano98b] Anonymous. Guidelines for contributors. Perspectives on Science, 6(1–2):204–206, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:GCb

[Ano98c] Anonymous. Guidelines for contributors. Perspectives on Science, 6 (3):335–337, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:GCc

[Ano98d] Anonymous. Guidelines for contributors. Perspectives on Science, 6(4):428–430, Winter 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:NCa

[Ano98e] Anonymous. Notes on contributors. Perspectives on Science, 6 (1–2):207–208, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:NCb

[Ano98f] Anonymous. Notes on contributors. Perspectives on Science, 6(3): 338–339, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1998:NCc

[Ano98g] Anonymous. Notes on contributors. Perspectives on Science, 6(4): 431–432, Winter 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Anonymous:1999:BR

[Ano99] Anonymous. Books received. Perspectives on Science, 7(2):285– 287, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.1999.7.2.285.

Anonymous:2001:BR

[Ano01a] Anonymous. Books received. Perspectives on Science, 9(1):116– 118, Spring 2001. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2001.9.1.116.

Anonymous:2001:EN

[Ano01b] Anonymous. Editors’ note. Perspectives on Science, 9(3):257–258, Fall 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://www.mitpressjournals.org/doi/abs/
10.1162/10636140160176134.

Anonymous:2002:BR

[Ano02] Anonymous. Books received. Perspectives on Science, 10(2):241– 246, Summer 2002. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/106361402321147559.

Anonymous:2003:BR

[Ano03a] Anonymous. Books received. Perspectives on Science, 11(3): 365–368, Fall 2003. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v011/11.3books_received.html;
http://muse.jhu.edu/journals/perspectives_on_science/v011/
11.3books_received.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/106361403773062696.

Anonymous:2003:NFH

[Ano03b] Anonymous, editor. New foundations in the history of astronomy. MIT Press, Cambridge, MA, USA, 2003. ISSN 1063-6145 (print), 1530-9274 (electronic). i–ii and 151–227 pp. Four papers in honor of Bernard R. Goldstein, Papers from the session of the History of Science Society Annual Meeting held in Pittsburgh, PA, 1999, Perspect. Sci. 10 (2002), no. 2.

Anonymous:2004:NC

[Ano04] Anonymous. Notes on contributors. Perspectives on Science, 12 (3):367–368, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/1063614042795408.

Anonymous:2005:FSI

[Ano05a] Anonymous. Forthcoming special issue: Technoscience and productivity, Part I. Perspectives on Science, 13(1):138–139, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/1063614053714081.

Anonymous:2005:GCa

[Ano05b] Anonymous. Guidelines for contributors. Perspectives on Science, 13(1):133–135, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/1063614053714072.

Anonymous:2005:GCb

[Ano05c] Anonymous. Guidelines for contributors. Perspectives on Science, 13(2):279–281, Summer 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/106361405774270502.

Anonymous:2005:NCa

[Ano05d] Anonymous. Notes on contributors. Perspectives on Science, 13(1): 136–137, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/1063614053714117.

Anonymous:2005:NCb

[Ano05e] Anonymous. Notes on contributors. Perspectives on Science, 13(2): 282–283, Summer 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/106361405774270511.

Anonymous:2006:GC

[Ano06a] Anonymous. Guidelines for contributors. Perspectives on Science, 14(4):574–576, Winter 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

www.mitpressjournals.org/doi/abs/10.1162/posc.2006.14.
4.574.

Anonymous:2006:NCa

[Ano06b] Anonymous. Notes on contributors. Perspectives on Science, 14(1): 125–126, Spring 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.1.125.

Anonymous:2006:NCb

[Ano06c] Anonymous. Notes on contributors. Perspectives on Science, 14(2): 255–256, Summer 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.2.255.

Anonymous:2006:NCc

[Ano06d] Anonymous. Notes on contributors. Perspectives on Science, 14 (3):385–386, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.3.385.

Anonymous:2006:NCd

[Ano06e] Anonymous. Notes on contributors. Perspectives on Science, 14(4): 577–578, Winter 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.4.577.

Anonymous:2008:NCa

[Ano08a] Anonymous. Notes on contributors. Perspectives on Science, 16(2): 234–235, Summer 2008. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v016/
16.2.contributors.pdf.

Anonymous:2008:NCb

[Ano08b] Anonymous. Notes on contributors. Perspectives on Science, 16 (3):306–307, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v016/
16.3.contributors.pdf.

Anonymous:2008:NCc

[Ano08c] Anonymous. Notes on contributors. Perspectives on Science, 16(4): 442–443, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v016/
16.4.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2008.16.4.442.

Anonymous:2009:GCa

[Ano09a] Anonymous. Guidelines for contributors. Perspectives on Science, 17(1):105–107, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc.2009.17.1.105.

Anonymous:2009:GCb

[Ano09b] Anonymous. Guidelines for contributors. Perspectives on Science, 17(2):232–234, Summer 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

www.mitpressjournals.org/doi/abs/10.1162/posc.2009.17.

2.232. Special Issue on History of Nanotechnology.

Anonymous:2009:JUT

[Ano09c] Anonymous. Journals under threat: a joint response from history of science, technology and medicine Editors. Perspectives on Science, 17(1):1–4, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.1.editor.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v017/17.
1.editor.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2009.17.1.1.

Anonymous:2009:NCa

[Ano09d] Anonymous. Notes on contributors. Perspectives on Science, 17(1): 108–109, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.1.contributors.html;

http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.1.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2009.17.1.108.

Anonymous:2009:NCb

[Ano09e] Anonymous. Notes on contributors. Perspectives on Science, 17 (2):235, Summer 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.2.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2009.17.2.235.

Anonymous:2009:NCc

[Ano09f] Anonymous. Notes on contributors. Perspectives on Science, 17 (3):??, Fall 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.3.contributors.pdf.

Anonymous:2009:NCd

[Ano09g] Anonymous. Notes on contributors. Perspectives on Science, 17(4): 510–511, Winter 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.4.contributors.pdf.

Anonymous:2010:NCa

[Ano10a] Anonymous. Notes on contributors. Perspectives on Science, 18(1): 116–117, Spring 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.1.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v018/
18.1.contributors.pdf.

Anonymous:2010:NCb

[Ano10b] Anonymous. Notes on contributors. Perspectives on Science, 18(2): 276–277, Summer 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.2.contributors.html;

http://muse.jhu.edu/journals/perspectives_on_science/v018/
18.2.contributors.pdf.

Anonymous:2010:NCc

[Ano10c] Anonymous. Notes on contributors. Perspectives on Science, 18 (3):381–382, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v018/
18.3.contributors.pdf.

Anonymous:2010:NCd

[Ano10d] Anonymous. Notes on contributors. Perspectives on Science, 18(4): 503–504, Winter 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v018/
18.4.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_x_00023.

Anonymous:2011:NCa

[Ano11a] Anonymous. Notes on contributors. Perspectives on Science, 19 (1):119, Spring 2011. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.1.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v019/
19.1.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_x_00027.

Anonymous:2011:NCb

[Ano11b] Anonymous. Notes on contributors. Perspectives on Science, 19(2): 233–234, Summer 2011. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v019/
19.2.contributors.pdf.

Anonymous:2011:NCc

[Ano11c] Anonymous. Notes on contributors. Perspectives on Science, 19 (3):358–359, Fall 2011. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.3.contributors.html;

http://muse.jhu.edu/journals/perspectives_on_science/v019/
19.3.contributors.pdf.

Anonymous:2011:NCd

[Ano11d] Anonymous. Notes on contributors. Perspectives on Science, 19(4): 485–486, Winter 2011. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v019/
19.4.contributors.pdf.

Anonymous:2012:FA

[Ano12a] Anonymous. Forthcoming articles. Perspectives on Science, 20 (1):??, Spring 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.1.article01.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v020/20.
1.article01.pdf.

Anonymous:2012:NCa

[Ano12b] Anonymous. Notes on contributors. Perspectives on Science, 20(1): 132–133, Spring 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.1.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v020/
20.1.contributors.pdf.

Anonymous:2012:NCb

[Ano12c] Anonymous. Notes on contributors. Perspectives on Science, 20(2): 271–272, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v020/
20.2.contributors.pdf.

Anonymous:2012:NCc

[Ano12d] Anonymous. Notes on contributors. Perspectives on Science, 20 (3):392–393, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v020/
20.3.contributors.pdf.

Anonymous:2012:NCd

[Ano12e] Anonymous. Notes on contributors. Perspectives on Science, 20(4): 533–534, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v020/
20.4.contributors.pdf.

Anonymous:2013:FAa

[Ano13a] Anonymous. Forthcoming articles. Perspectives on Science, 21 (1):??, Spring 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.1.article.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v021/21.
1.article.pdf.

Anonymous:2013:FAb

[Ano13b] Anonymous. Forthcoming articles. Perspectives on Science, 21 (2):i, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.2.article.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v021/21.

2.article.pdf. Special Issue on Modeling in the Social Sciences.

Anonymous:2013:FAc

[Ano13c] Anonymous. Forthcoming articles. Perspectives on Science, 21 (4):??, Winter 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.4.article.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v021/21.
4.article.pdf.

Anonymous:2013:NCa

[Ano13d] Anonymous. Notes on contributors. Perspectives on Science, 21(1): 140–141, Spring 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.1.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v021/
21.1.contributors.pdf.

Anonymous:2013:NCb

[Ano13e] Anonymous. Notes on contributors. Perspectives on Science, 21(2): 276–278, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v021/

21.2.contributors.pdf. Special Issue on Modeling in the Social Sciences.

Anonymous:2013:NCc

[Ano13f] Anonymous. Notes on contributors. Perspectives on Science, 21 (3):390–392, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v021/
21.3.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_x_00109.

Anonymous:2013:NCd

[Ano13g] Anonymous. Notes on contributors. Perspectives on Science, 21(4): 485–486, Winter 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v021/
21.4.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_x_00115.

Anonymous:2014:NCa

[Ano14a] Anonymous. Notes on contributors. Perspectives on Science, 22(1): 161–162, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.1.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v022/
22.1.contributors.pdf.

Anonymous:2014:NCb

[Ano14b] Anonymous. Notes on contributors. Perspectives on Science, 22(2): 309–310, Summer 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v022/
22.2.contributors.pdf.

Anonymous:2014:NCc

[Ano14c] Anonymous. Notes on contributors. Perspectives on Science, 22 (3):421–422, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v022/
22.3.contributors.pdf.

Anonymous:2014:NCd

[Ano14d] Anonymous. Notes on contributors. Perspectives on Science, 22(4): 650–651, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.4.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v022/
22.4.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_x_00154.

Anonymous:2015:NCa

[Ano15a] Anonymous. Notes on contributors. Perspectives on Science, 23(1): 146–147, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.1.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v023/
23.1.contributors.pdf.

Anonymous:2015:NCb

[Ano15b] Anonymous. Notes on contributors. Perspectives on Science, 23(2): 241–242, Summer 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.2.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v023/
23.2.contributors.pdf.

Anonymous:2015:NCc

[Ano15c] Anonymous. Notes on contributors. Perspectives on Science, 23 (3):379–380, Fall 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.3.contributors.html;
http://muse.jhu.edu/journals/perspectives_on_science/v023/
23.3.contributors.pdf.

Anonymous:2015:NCd

[Ano15d] Anonymous. Notes on contributors. Perspectives on Science, 23(4): 485–486, Winter 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.4.contributors.html;

http://muse.jhu.edu/journals/perspectives_on_science/v023/
23.4.contributors.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_x_00185.

Anonymous:2016:MJL

[Ano16a] Anonymous. In memoriam: Joan Lisa Bromberg, 1929–2015. Perspectives on Science, 24(1):144, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v024/24.1.article.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v024/24.1.article.pdf.

Anonymous:2016:NC

[Ano16b] Anonymous. Notes on contributors. Perspectives on Science, 24(1):142–143, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v024/24.
1.contributors.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v024/24.1.contributors.pdf.

Anonymous:2020:E

[Ano20] Anonymous. Erratum. Perspectives on Science, 28(5):655, September/October 2020. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_e_00354.

Anta:2023:CIC

[Ant23] Javier Anta. Can information concepts have physical content? Perspectives on Science, 31(2):207–232, March/April 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
2/207/2077437/posc_a_00424.pdf.

Akopyan:2022:IQD

[AO22] Ovanes Akopyan and Pietro Daniel Omodeo. Introduction: Quis dixit? The vicissitudes of authority in early modern cosmology. Perspectives on Science, 30(5):819–825, September/October 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
5/819/2053605/posc_e_00560.pdf.

Ariew:1994:DIY

[Ari94] Roger Ariew. Damned if you do: Cartesians and censorship, 1663– 1706. Perspectives on Science, 2(3):255–274, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Ariew:2004:IGP

[Ari04] Roger Ariew. Introduction: Galileo in Paris. Perspectives on Science, 12(2):131–134, Summer 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v012/12.
2ariew.html; http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.2ariew.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361404323119844.  Special Issue:

Galileo in Paris.

Arikha:2006:FFE

[Ari06] Noga Arikha. Form and function in the early enlightenment. Perspectives on Science, 14(2):153–188, Summer 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.2arikha.html; http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.2arikha.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/posc.2006.14.
2.153.

Ariew:2007:DP

[Ari07] Roger Ariew. Descartes and Pascal. Perspectives on Science, 15(4): 397–409, Winter 2007. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v015/15.4ariew.html; http://muse.
jhu.edu/journals/perspectives_on_science/v015/15.4ariew.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2007.15.4.397.

Ariew:2018:DFC

[Ari18] Roger Ariew. Descartes and the first Cartesians revisited. Perspectives on Science, 26(5):599–617, September/October 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00291.

Ariew:2021:NCL

[Ari21] Roger Ariew. The nature of Cartesian logic. Perspectives on Science, 29(3):275–291, May/June 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/3/275/1922997/posc_
a_00369.pdf; https://muse.jhu.edu/article/796016.

Arthur:1998:CDH

[Art98] Richard Arthur. Cohesion, division and harmony: Physical aspects of Leibniz’s continuum problem (1671–1686). Perspectives on Science, 6(1–2):110–135, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v006/6.

1arthur.html. Special Issue: Leibniz and the Sciences.

Arabatzis:2012:IWI

[AS12] Theodore Arabatzis and Jutta Schickore. Introduction: Ways of integrating history and philosophy of science. Perspectives on Science, 20(4):395–408, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v020/20.
4.arabatzis.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.4.arabatzis.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_x_00079.

Ash:2019:IHP

[Ash19] Mitchell G. Ash. Interdisciplinarity in historical perspective. Perspectives on Science, 27(4):619–642, July/August 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00318.

Arroyo-Santos:2015:POC

[ASOVS15] Alfonso Arroyo-Santos, Mark E. Olson, and Francisco Vergara-Silva. Practice-oriented controversies and borrowed epistemic credibility in current evolutionary biology: Phylogeography as a case study. Perspectives on Science, 23(3):310–334, Fall 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.3.arroyo-santos.html;   http://muse.
jhu.edu/journals/perspectives_on_science/v023/23.3.arroyo-
santos.pdf.

Augier:2000:MHS

[Aug00] Mie Augier. Models of Herbert A. Simon. Perspectives on Science, 8(4):407–443, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v008/8.
4augier.html; http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.4augier.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400753373768.

Blute:2011:RGT

[BA11] Marion Blute and Paul Armstrong. The reinvention of grand theories of the scientific/scholarly process. Perspectives on Science, 19(4):391–425, Winter 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v019/19.4.blute.html;
http://muse.jhu.edu/journals/perspectives_on_science/v019/
19.4.blute.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00046.

Babes:2022:PAC

[Bab22] Ovidiu Babes. Playing with the ancients: The cosmology of Gilles Personne de Roberval. Perspectives on Science, 30(6):950–981, November/December 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/6/950/2059522/posc_a_00565.pdf.

Baertschi:2024:RFP

[Bae24] Bernard Baertschi. On the reform of the first philosophy: After Leibniz, Maine de Biran. Perspectives on Science, 32(1):15– 27, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/32/1/15/2210168/posc_a_00596.pdf.

Baird:1999:IHP

[Bai99] Davis Baird. Internal history and the philosophy of experiment. Perspectives on Science, 7(3):383–407, Fall 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v007/7.3baird.html; http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.3baird.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.1999.7.3.383.

Balmer:1996:PCH

[Bal96] Brian Balmer. The political cartography of the Human Genome Project. Perspectives on Science, 4(3):249–282, Fall 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://search.ebscohost.com/login.aspx?direct=true&db=
aph&AN=6369646&site=ehost-live.

Baldassarri:2020:DDB

[Bal20] Fabrizio Baldassarri. Descartes and the Dutch: Botanical experimentation in the Early Modern Period. Perspectives on Science, 28 (6):657–683, November/December 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00355.

Bangu:2006:PHP

[Ban06] Sorin Bangu. Pythagorean heuristic in physics. Perspectives on Science, 14(4):387–416, Winter 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
4bangu.html; http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.4bangu.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.4.387.

Barker:2001:KIC

[Bar01] Peter Barker. Kuhn, incommensurability, and cognitive science. Perspectives on Science, 9(4):433–462, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.4barker.html; http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.4barker.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361401760375820.

Barker:2002:CC

[Bar02a] Peter Barker. Constructing Copernicus. Perspectives on Science, 10(2):208–227, Summer 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v010/10.
2barker02.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.2barker02.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361402321147531. New foundations in

the history of astronomy (Pittsburgh, PA, 1999).

Barker:2002:NFH

[Bar02b] Peter Barker. New foundations in the history of astronomy: Four papers in honor of Bernard R. Goldstein. Perspectives on Science, 10(2):151–154, Summer 2002. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/106361402321147496.

Barnes:2005:EMT

[Bar05] Barry Barnes. Elusive memories of technoscience. Perspectives on Science, 13(2):142–165, Summer 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v013/13.
2barnes.html; http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.2barnes.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361405774270520.  Special Issue:

Technoscientific Productivity, Part 1.

Barwich:2015:BMB

[Bar15] Ann-Sophie Barwich. Bending molecules or bending the rules?: The application of theoretical models in fragrance chemistry. Perspectives on Science, 23(4):443–465, Winter 2015. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.4.barwich.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v023/23.4.barwich.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00183.

Barahona:2020:WWC

[Bar20] Ana Barahona. Women and the workplace. collaborative networks of women geneticists in Mexico in the 1960s and early 1970s. Perspectives on Science, 28(2):119–126, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00338.

Barwich:2021:FGH

[Bar21] Ann-Sophie Barwich. Fishing for genes: How the largest gene family in the mammalian genome was found (and why idiosyncrasy in exploration matters). Perspectives on Science, 29(4):359–387, July/August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/359/1959242/posc_a_00375.pdf; https:/
/muse.jhu.edu/article/804403.

Barbosa:2022:RNN

[Bar22a] Thiago P. Barbosa. Racializing a new nation: German coloniality and anthropology in Maharashtra, India. Perspectives on Science, 30(1):137–166, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/1/137/1987176/posc_
a_00405.pdf.

Barton:2022:PJH

[Bar22b] William M. Barton. The poetry of Jeremiah Horrocks’s Venus in sole visa (1662): Astronomy, authority, and the ‘new science’. Perspectives on Science, 30(6):982–1004, November/December 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
6/982/2059524/posc_a_00566.pdf.

Batterman:2019:URE

[Bat19] Robert W. Batterman. Universality and RG explanations. Perspectives on Science, 27(1):26–47, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00298.

Bauer:2003:PSI

[Bau03] Henry H. Bauer. The progress of science and implications for science studies and for science policy. Perspectives on Science, 11(2):236–278, Summer 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v011/11.
2bauer.html; http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.2bauer.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361403322495902.

Baird:1999:WT

[BC99] Davis Baird and Mark S. Cohen. Why trade? Perspectives on Science, 7(2):231–254, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v007/7.2baird02.html;

http://www.mitpressjournals.org/doi/abs/10.1162/posc.1999.
7.2.231.

Begley:2020:LLC

[Beg20] Justin Begley. The lost liquid cosmogony of Johannes Daniel Schlichting (1705–1765). Perspectives on Science, 28(5):571– 609, September/October 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00351.

BertoloniMeli:2004:RNT

[Ber04] Domenico Bertoloni Meli. The role of numerical tables in Galileo and Mersenne. Perspectives on Science, 12(2):164–190, Summer 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.2meli.html; http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.2meli.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361404323119862.

Special Issue: Galileo in Paris.

Bandelj:2016:CRF

[BERW16] Nina Bandelj, Julia Elyachar, Gary Richardson, and James Owen Weatherall. Comprehending and regulating financial crises: an interdisciplinary approach. Perspectives on Science, 24 (4):443–473, July/August 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/625740;   http://muse.jhu.edu/article/
625740/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00216.

Best:2015:MIB

[Bes15] Nicholas W. Best. Meta-incommensurability between theories of meaning: Chemical evidence. Perspectives on Science, 23(3): 361–378, Fall 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.3.best.html;   http://
muse.jhu.edu/journals/perspectives_on_science/v023/23.
3.best.pdf.

Barker:1998:RIS

[BG98] Peter Barker and Bernard R. Goldstein. Realism and instrumentalism in Sixteenth Century astronomy: a reappraisal. Perspectives on Science, 6(3):232–258, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v006/6.3barker.html.

Brown:2017:DPS

[BH17] Matthew J. Brown and Joyce C. Havstad. The disconnect problem, scientific authority, and climate policy. Perspectives on Science, 25(1):67–94, January/February 2017. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/647989;   http://muse.
jhu.edu/article/647989/pdf.

Biagioli:2022:RMI

[Bia22] Mario Biagioli. Replicating mathematical inventions: Galileo’s compass, its instructions, its students. Perspectives on Science, 30 (3):437–462, May/June 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/3/437/2023839/posc_a_00422.pdf.

Biddle:2016:IRE

[Bid16] Justin B. Biddle. Inductive risk, epistemic risk, and overdiagnosis of disease. Perspectives on Science, 24(2):192–205, March/April 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00200.

Biener:2004:GFN

[Bie04] Zvi Biener. Galileo’s first new science: The science of matter. Perspectives on Science, 12(3):262–287, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v012/12.
3biener.html; http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.3biener.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/1063614042795462.

Bailer-Jones:2002:STS

[BJ02] Daniela M. Bailer-Jones. Scientists’ thoughts on scientific models. Perspectives on Science, 10(3):275–301, Fall 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.3bailer_jones.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v010/10.3bailer_jones.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/106361402321899069.

Blair:1994:TIE

[Bla94] Ann Blair. Tradition and innovation in early modern natural philosophy: Jean Bodin and Jean-Cecile Frey. Perspectives on Science, 2(4):428–454, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Blank:2008:JCS

[Bla08] Andreas Blank. Julius Caesar Scaliger on corpuscles and the vacuum. Perspectives on Science, 16(2):137–159, Summer 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.2.blank.html; http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.2.blank.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/posc.2008.16.
2.137.

Blank:2011:DSP

[Bla11] Andreas Blank. Daniel Sennert on poisons, epilepsy, and subordinate forms. Perspectives on Science, 19(2):192–211, Summer 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.2.blank.html; http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.2.blank.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00036.

Blank:2013:FLM

[Bla13] Andreas Blank. Fortunio Liceti on mind, light, and immaterial extension. Perspectives on Science, 21(3):358–378, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.3.blank.html; http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.3.blank.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00103.

Blank:2018:SCP

[Bla18] Andreas Blank. Sixteenth-Century pharmacology and the controversy between reductionism and emergentism. Perspectives on Science, 26(2):157–184, March/April 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00271.

Bloor:2005:TSE

[Blo05] David Bloor. Toward a sociology of epistemic things. Perspectives on Science, 13(3):285–312, Fall 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v013/13.
3bloor.html; http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.3bloor.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361405774287982.  Special Issue:

Technoscientific Productivity, Part 2. See reply [Rhe05b].

Blustein:1993:MBN

[Blu93] Bonnie Ellen Blustein. Medicine as biology: Neuropsychiatry at the University of Chicago, 1928–1939. Perspectives on Science, 1 (3):416–444, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Baltas:2004:APS

[BM04] Aristeides Baltas and Peter K. Machamer. Athens–Pittsburgh symposium in the history and philosophy of science and technology. Perspectives on Science, 12(3):243, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v012/12.
3baltas.html; http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.3baltas.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/1063614042795444.

Baird:1999:EIF

[BN99] Davis Baird and Alfred Nordmann. Editors’ introduction: Forays into the trading zone of Image and Logic. Perspectives on Science, 7(2):147–150, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v007/7.2baird01.html;
http://www.mitpressjournals.org/doi/abs/10.1162/posc.1999.
7.2.147.

Bohlin:2012:FSM

[Boh12] Ingemar Bohlin. Formalizing syntheses of medical knowledge: The rise of meta-analysis and systematic reviews. Perspectives on Science, 20(3):273–309, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v020/20.3.bohlin.
html;   http://muse.jhu.edu/journals/perspectives_on_science/

v020/20.3.bohlin.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_a_00075.

Bokulich:2001:RTE

[Bok01] Alisa Bokulich. Rethinking thought experiments. Perspectives on Science, 9(3):285–307, Fall 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v009/9.3bokulich.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v009/9.3bokulich.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/10636140160176152.

Bokulich:2008:PDE

[Bok08] Alisa Bokulich. Paul Dirac and the Einstein–Bohr debate. Perspectives on Science, 16(1):103–114, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v016/16.
1bokulich.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.1bokulich.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2008.16.1.103.

Bokulich:2014:PPP

[Bok14] Alisa Bokulich. Pluto and the ‘planet problem’: Folk concepts and natural kinds in astronomy. Perspectives on Science, 22(4): 464–490, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.4.bokulich.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v022/22.
4.bokulich.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00146.

Borg:2019:DIH

[Bor19] George Borg. Discovery and instrumentation: How surplus knowledge contributes to progress in science. Perspectives on Science, 27 (6):861–890, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00328.

Borrelli:2021:CSE

[Bor21] Arianna Borrelli. Constructing strangeness: Exploratory modeling and concept formation. Perspectives on Science, 29(4):388–408, July/August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/388/1959224/posc_a_00376.pdf; https:/
/muse.jhu.edu/article/804404.

Bouchard:2016:RIT

[Bou16] Frédéric Bouchard. The roles of institutional trust and distrust in grounding rational deference to scientific expertise. Perspectives on Science, 24(5):582–608, September/October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://muse.jhu.edu/article/628782;   http://muse.jhu.
edu/article/628782/pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00224.

Bowen:2002:SEH

[Bow02] Alan C. Bowen. Simplicius and the early history of Greek planetary theory. Perspectives on Science, 10(2):155–167, Summer 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.2bowen.html; http://muse.jhu.edu/journals/
perspectives_on_science/v010/10.2bowen.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361402321147504.

New foundations in the history of astronomy (Pittsburgh, PA, 1999).

Bowen:2007:DPT

[Bow07] Alan C. Bowen. The demarcation of physical theory and astronomy by Geminus and Ptolemy. Perspectives on Science, 15(3): 327–358, Fall 2007. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v015/15.3bowen.pdf; http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2007.15.3.327.

Barahona:2020:SIH

[BR20a] Ana Barahona and Marsha L. Richmond. Special issue: Heredity and evolution in an Ibero–American context. Perspectives on Science, 28(2):119–126, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_e_00335.

Bedessem:2020:CSS

[BR20b] Baptiste Bedessem and Stéphanie Ruphy. Citizen science and scientific objectivity: Mapping out epistemic risks and benefits. Perspectives on Science, 28(5):630–654, September/October 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00353.

Brockliss:1995:DGR

[Bro95] Laurence W. B. Brockliss. Descartes, Gassendi, and the reception of the mechanical philosophy at the French Collèges de Plein Exercice, 1640–1730. Perspectives on Science, 3(4):450–479, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi.

Brown:2014:ESU

[Bro14] James Robert Brown. Explaining, seeing, and understanding in thought experiments. Perspectives on Science, 22(3): 357–376, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v022/22.3.brown.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v022/22.3.brown.pdf.

Bromberg:2016:CNE

[Bro16a] Joan Lisa Bromberg. Coherence and noise in the era of the maser. Perspectives on Science, 24(1):93–111, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v024/24.1.bromberg.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v024/24.1.bromberg.
pdf.

Brown:2016:PP

[Bro16b] James Robert Brown. Patents and progress. Perspectives on Science, 24(5):505–528, September/October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://muse.jhu.edu/article/628779;   http://muse.jhu.
edu/article/628779/pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00221.

Brown:2018:HDF

[Bro18a] James Robert Brown. How do Feynman diagrams work? Perspectives on Science, 26(4):423–442, July/August 2018. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00281.

Brown:2018:I

[Bro18b] James Robert Brown. Introduction. Perspectives on Science, 26(4):419–422, July/August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00280.

Brooks:2023:TCE

[Bro23] Daniel S. Brooks. Themes of consolidation in Eugene P. Odum’s publicization of the levels concept in ecology textbooks, 1953– 1975. Perspectives on Science, 31(4):437–464, ???? 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

https://muse.jhu.edu/pub/6/article/902862.

Brush:1993:PTE

[Bru93] Stephen G. Brush. Prediction and theory evaluation: Cosmic microwaves and the revival of the Big Bang. Perspectives on Science, 1(4):565–602, Winter 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Buchwald:2001:IDE

[BS01] Jed Z. Buchwald and George E. (George Edwin) Smith. Incommensurability and the discontinuity of evidence. Perspectives on Science, 9(4):463–498, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v009/9.4buchwald.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v009/9.4buchwald.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361401760375839.

Buchwald:2010:RTK

[Buc10] Jed Z. Buchwald. A reminiscence of Thomas Kuhn. Perspectives on Science, 18(3):279–283, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
3.buchwald.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.3.buchwald.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00009.

Bueter:2019:MDP

[Bue19] Anke Bueter. A multi-dimensional pluralist response to the DSMcontroversies. Perspectives on Science, 27(2):316–343, March/April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00309.

Bullock:2016:RMA

[Bul16] Shawn M. Bullock. Radar, modems, and air defense systems: Noise as a data communication problem in the 1950s. Perspectives on Science, 24(1):73–92, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v024/24.
1.bullock.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v024/24.1.bullock.pdf.

Burkhardt:1999:SVM

[Bur99] Jeffrey Burkhardt. Scientific values and moral education in the teaching of science. Perspectives on Science, 7(1):87–110, Spring 1999. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.1burkhardt.html; http:/
/www.mitpressjournals.org/doi/abs/10.1162/posc.1999.7.
1.87.

Burian:2001:DCS

[Bur01] Richard M. Burian. The dilemma of case studies resolved: The virtues of using case studies in the history and philosophy of science. Perspectives on Science, 9(4):383–404, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.4burian.html; http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.4burian.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361401760375794.

Burian:2002:CPR

[Bur02] Richard M. Burian. Comments on the precarious relationship between history and philosophy of science. Perspectives on Science, 10(4):398–407, Winter 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v010/10.4burian.html;
http://muse.jhu.edu/journals/perspectives_on_science/v010/
10.4burian.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361402322288039.

Burman:2008:ERP

[Bur08] Jeremy Trevelyan Burman. Experimenting in relation to Piaget: Education is a chaperoned process of adaptation. Perspectives on Science, 16(2):160–195, Summer 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v016/16.2.burman.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v016/16.2.burman.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2008.16.2.160.

Burman:2012:MMB

[Bur12] Jeremy Trevelyan Burman. The misunderstanding of memes: Biography of an unscientific object, 1976-1999. Perspectives on Science, 20(1):75–104, Spring 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v020/20.1.burman.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v020/20.1.burman.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_a_00057.

Campbell:1993:PCB

[Cam93] Donald T. Campbell. Plausible coselection of belief by referent: All the objectivity that is possible. Perspectives on Science, 1(1): 88–108, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Camilleri:2007:ILC

[Cam07] Kristian Camilleri. Indeterminacy and the limits of classical concepts: The transformation of Heisenberg’s thought. Perspectives on Science, 15(2):178–201, Summer 2007. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.2camilleri.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v015/15.2camilleri.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2007.15.2.178.

Camilleri:2009:CMC

[Cam09a] Kristian Camilleri. Constructing the myth of the Copenhagen interpretation. Perspectives on Science, 17(1):26–57, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_

on_science/v017/17.1.camilleri.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v017/17.1.camilleri.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2009.17.1.26.

Campos:2009:FFP

[Cam09b] Daniel G. Campos. The framing of the fundamental probability set: a historical case study on the context of mathematical discovery. Perspectives on Science, 17(4):385–416, Winter 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.4.campos.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v017/17.
4.campos.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2009.17.4.385.

Camposampiero:2016:BIC

[Cam16] Matteo Favaretti Camposampiero. Bodies of inference: Christian Wolff’s epistemology of the life sciences and medicine. Perspectives on Science, 24(3):361–379, May/June 2016. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/618110;   http://muse.
jhu.edu/article/618110/pdf.

Cartwright:1999:LES

[Car99] Nancy Cartwright. The limits of exact science, from economics to physics. Perspectives on Science, 7(3):318–336, Fall 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v007/7.3cartwright.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v007/7.3cartwright.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.1999.7.3.318.

Carraud:2007:PAA

[Car07] Vincent Carraud. Pascal’s anti-Augustinianism. Perspectives on Science, 15(4):450–492, Winter 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v015/15.4carraud.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v015/15.4carraud.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2007.15.4.450.

Carusi:2011:CBL

[Car11]  Annamaria Carusi.  Computational biology and the lim-

its of shared vision. Perspectives on Science, 19(3):300– 336, Fall 2011. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.3.carusi.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v019/19.
3.carusi.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00044.

Carrier:2017:FCC

[Car17] Martin Carrier. Facing the credibility crisis of science: On the ambivalent role of pluralism in establishing relevance and reliability. Perspectives on Science, 25(4):439–464, July/August 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00249.

Cassan:2021:ILM

[Cas21a] Elodie Cassan. Introduction: Logic and methodology in the Early Modern Period. Perspectives on Science, 29(3):237–254, May/June 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/3/237/1922999/posc_e_00367.pdf; https:/
/muse.jhu.edu/article/796019.

Cassan:2021:NLB

[Cas21b] Elodie Cassan. A new logic: Bacon’s Novum Organum. Perspectives on Science, 29(3):255–274, May/June 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/29/3/255/
1922995/posc_a_00368.pdf; https://muse.jhu.edu/article/
796020.

Cat:2007:SGG

[Cat07] Jordi Cat. Switching gestalts on gestalt psychology: On the relation between science and philosophy. Perspectives on Science, 15(2):131–177, Summer 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/journals/perspectives_on_science/v015/15.2cat.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v015/15.2cat.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2007.15.2.131.

Cavicchi:2006:FPE

[Cav06] Elizabeth Cavicchi. Faraday and Piaget: Experimenting in relation with the world. Perspectives on Science, 14(1):66–96, Spring 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.1cavicchi.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v014/14.1cavicchi.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.

2006.14.1.66. Special Issue: Cognitive Psychology and Faraday’s Experiments.

Collins:2017:DFS

[CBRG17] Harry Collins, Andrew Bartlett, and Luis Reyes-Galindo. Demarcating fringe science for policy. Perspectives on Science, 25(4):411– 438, July/August 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00248.

Chalozin-Dovrat:2019:HSB

[CD19] Lin Chalozin-Dovrat. The history of SPACE between science and ordinary language: What can words tell us about conceptual change? Perspectives on Science, 27(2):244–277, March/April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00307.

Canizares-Esguerra:2004:ISR

[CE04] Jorge Cañizares-Esguerra. Iberian science in the Renaissance: Ignored how much longer? Perspectives on Science, 12(1):86– 124, Spring 2004. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.1esguerra.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v012/12.
1esguerra.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/106361404773843355.

Claveau:2015:ECM

[CF15] François Claveau and Melissa Vergara Fernández. Epistemic contributions of models: Conditions for propositional learning. Perspectives on Science, 23(4):405–423, Winter 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v023/23.

4.claveau.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.4.claveau.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00181.

Corneanu:2012:I

[CGJ12] Sorana Corneanu, Guido Giglioni, and Dana Jalobeanu. Introduction. Perspectives on Science, 20(2):135–138, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.2.corneanu.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.2.corneanu.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_e_00059.

Chang:1995:CRM

[Cha95] Hasok Chang. Circularity and reliability in measurement. Perspectives on Science, 3(2):153–172, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Chabas:2002:DAT

[Cha02] José Chabás. The diffusion of the Alfonsine Tables: The case of the Tabulae resolutae. Perspectives on Science, 10(2): 168–178, Summer 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v010/10.2chabas.html;
http://muse.jhu.edu/journals/perspectives_on_science/v010/
10.2chabas.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361402321147513. New foundations in the his-

tory of astronomy (Pittsburgh, PA, 1999).

Chaplin:2006:BFS

[Cha06] Joyce E. Chaplin. Benjamin Franklin and science, continuing opportunities. Perspectives on Science, 14(2):232–251, Summer 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.2chaplin.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v014/14.
2chaplin.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2006.14.2.232.

Charalampous:2021:CAJ

[Cha21] Charis Charalampous. The confined atom: James Clerk Maxwell on the fundamental particles and the limits of scientific knowledge.

Perspectives on Science, 29(2):189–214, March/April 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://muse.jhu.edu/article/787908.

Chandler:2023:ANT

[Cha23] David Chandler. Actor network theory and sensing governance: From causation to correlation. Perspectives on Science, 31(1):139– 158, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/31/1/139/2073129/posc_a_00584.pdf.

Clever:2022:PMI

[CHB22] Iris Clever, Jaehwan Hyun, and Elise K. Burton. People in motion: Introduction to transnational movements and transwar connections in the anthropological and genetic study of human populations. Perspectives on Science, 30(1):1–12, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
1/1/1987178/posc_e_00400.pdf.

Chen:1994:HDS

[Che94] Xiang Chen. How do scientists have disagreements about experiments? Incommensurability in the use of goal derived categories. Perspectives on Science, 2(3):275–301, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Chen:2000:SSU

[Che00] Xiang Chen. To see or not to see: The uses of photometers and measurements of reflective power. Perspectives on Science, 8(1):1–28, Spring 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v008/8.1chen.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361400567993.

Chen:2005:TTK

[Che05] Xiang Chen. Transforming temporal knowledge: Conceptual change between event concepts. Perspectives on Science, 13(1): 49–73, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v013/13.1chen.html; http://muse.
jhu.edu/journals/perspectives_on_science/v013/13.1chen.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
1063614053714108.

Chen:2014:WWR

[Che14] Xiang Chen. Why are we reluctant to act immediately on climate change?: From ontological assumptions to core cognition. Perspectives on Science, 22(4):574–592, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
4.chen.html; http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.4.chen.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00150.

Chirimuuta:2020:RMC

[Chi20] M. Chirimuuta. The reflex machine and the cybernetic brain: The critique of abstraction and its application to computationalism. Perspectives on Science, 28(3):421–457, May/June 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00346.

Chvaja:2020:WDM

[Chv20] Radim Chvaja. Why did memetics fail? Comparative case study. Perspectives on Science, 28(4):542–570, July/August 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00350.

Cittadino:1993:MCP

[Cit93] Eugene Cittadino. A “marvelous cosmopolitan preserve”: The Dunes, Chicago, and the dynamic ecology of Henry Cowles. Perspectives on Science, 1(3):520–559, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Campbell-Kelly:2013:RMM

[CK13] Martin Campbell-Kelly. Remembering Michael S. Mahoney. Perspectives on Science, 21(3):379–383, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v021/21.
3.campbell-kelly.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.3.campbell-kelly.pdf; http:
//www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00104.

Clark:1993:REM

[Cla93a] William Clark. Review essay: The misogyny of scholars. Perspectives on Science, 1(2):342–357, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Clarke:1993:MSL

[Cla93b] Adele E. Clarke. Money, sex and legitimacy at Chicago, circa 1892–1940: Lillie’s Center of Reproductive Biology. Perspectives on Science, 1(3):367–415, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Clever:2022:MTA

[Cle22] Iris Clever. Miriam Tildesley and the anthropological politics of standardizing racial measurements. Perspectives on Science, 30(1): 13–47, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/1/13/1987218/posc_a_00401.pdf.

Clough:2008:SEN

[Clo08] Sharyn Clough. Solomon’s empirical/non-empirical distinction and the proper place of values in science. Perspectives on Science, 16 (3):265–279, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.3.clough.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v016/16.
3.clough.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2008.16.3.265.

Carrillo:2023:SIM

[CM23] Natalia Carrillo and Sergio Martínez. Scientific inquiry: From metaphors to abstraction. Perspectives on Science, 31(2):233–261, March/April 2023. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/31/2/233/2077439/posc_a_00571.pdf.

Cobb:2012:JFW

[Cob12] Aaron D. Cobb. Is John F. W. Herschel an inductivist about hypothetical inquiry? Perspectives on Science, 20(4):409–439, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.4.cobb.html; http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.4.cobb.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/POSC_a_00080.

Coko:2020:MDM

[Cok20] Klodian Coko. The multiple dimensions of multiple determination. Perspectives on Science, 28(4):505–541, July/August 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00349.

Collacciani:2018:CHC

[Col18] Domenico Collacciani. A contextualist history of Cartesian philosophy: Roger Ariew’s Descartes and the first Cartesians. Perspectives on Science, 26(5):521–532, September/October 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00286.

Cooper:2004:CSU

[Coo04] Rachel Cooper. Can sociologists understand other forms of life? Perspectives on Science, 12(1):29–54, Spring 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.1cooper.html; http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.1cooper.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/106361404773843337.

Copeland:2018:FLU

[Cop18] Samantha Marie Copeland. ”Fleming leapt on the unusual like a weasel on a vole”: Challenging the paradigms of discovery in science. Perspectives on Science, 26(6):694–721, November/December 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00294.

Corneanu:2021:LMR

[Cor21] Sorana Corneanu. Logic and the movement of reasoning: Pierre Gassendi on the three acts of the mind. Perspectives on Science, 29(3):292–326, May/June 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/3/292/1923003/posc_
a_00370.pdf; https://muse.jhu.edu/article/796017.

Cozzoli:2010:DMO

[Coz10] Daniele Cozzoli. The development of Mersenne’s optics. Perspectives on Science, 18(1):9–25, Spring 2010. CODEN PRSIEU.

ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
1.cozzoli.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.1.cozzoli.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2010.18.1.9.

Coutinho:1999:RFC

[CP99] Marilia Coutinho and João Carlos Pinto Dias. The rise and fall of Chagas disease. Perspectives on Science, 7(4):447–485, Winter 1999. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.4coutinho.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v007/7.
4coutinho.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.1999.7.4.447.

Craig:2015:NDE

[Cra15] Lindsay R. Craig. Neo-Darwinism and evo-devo: an argument for theoretical pluralism in evolutionary biology. Perspectives on Science, 23(3):243–279, Fall 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v023/23.3.craig.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.3.craig.pdf.

Collins:2023:HSS

[CSB^{+}23] Harry Collins, Jeff Shrager, Andrew Bartlett, Shannon Conley, Rachel Hale, and Robert Evans. Hypernormal science and its significance. Perspectives on Science, 31(2):262–292, March/April 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
2/262/2077442/posc_a_00572.pdf.

Currie:2021:SFL

[Cur21] Adrian Currie. Stepping forwards by looking back: Underdetermination, epistemic scarcity and legacy data. Perspectives on Science, 29(1):104–132, January/February 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00362.

Corneanu:2012:IIF

[CV12] Sorana Corneanu and Koen Vermeir. Idols of the imagination: Francis Bacon on the imagination and the medicine of the mind. Perspectives on Science, 20(2):183–206, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.2.corneanu01.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.2.corneanu01.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00062.

Currie:2018:NIO

[CW18] Adrian Currie and Kirsten Walsh. Newton on Islandworld: Onticdriven explanations of scientific method. Perspectives on Science, 26(1):119–156, January/February 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00270.

Dambock:2022:PCN

[Dam22] Christian Damböck. The politics of Carnap’s non-cognitivism and the scientific world-conception of left-wing logical empiricism. Perspectives on Science, 30(4):493–524, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
4/493/2038157/posc_a_00372.pdf.

Dasgupta:2008:SCL

[Das08] Subrata Dasgupta. Shedding computational light on human creativity. Perspectives on Science, 16(2):121–136, Summer 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.2.dasgupta.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v016/16.2.dasgupta.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2008.16.2.121.

Dasgupta:2016:SPC

[Das16] Deepanwita Dasgupta. Scientific practice in the contexts of peripheral science: C. V. Raman and his construction of a mechanical violin-player. Perspectives on Science, 24(4):381–395, July/ August 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/625736;
 http://muse.jhu.edu/article/625736/pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/POSC_a_00212.

Dawes:2016:ESG

[Daw16] Gregory Dawes. Experiment, speculation, and Galileo’s scientific reasoning. Perspectives on Science, 24(3):343–360, May/June 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/618109;   http://muse.
jhu.edu/article/618109/pdf.

deBruyckere:2013:BGR

[dBV13] Anna de Bruyckere and Maarten Van Dyck. Being in or getting at the real: Kochan on Rouse, Heidegger and minimal realism. Perspectives on Science, 21(4):453–462, Winter 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.4.de-bruyckere.html;   http://muse.
jhu.edu/journals/perspectives_on_science/v021/21.4.de-
bruyckere.pdf; http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00112.

deChadarevian:2017:MHP

[dC17] Soraya de Chadarevian. Making human populations. Perspectives on Science, 25(5):698–703, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00260.

deCarvalho:2020:MEB

[dC20] Tito Brige de Carvalho. Modern evolutionary biology and Brazilian population genetics: Theodosius Dobzhansky at the University of São paulo. Perspectives on Science, 28(2):223–243, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00339.

deChadarevian:2022:CNT

[dC22a] Soraya de Chadarevian. Commentary: Nationalism and transnationalism in anthropological research. Perspectives on Science, 30(1):194–198, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/1/194/1987182/posc_
e_00407.pdf.

deCourtenay:2022:PSR

[dC22b] Nadine de Courtenay. On the philosophical significance of the reform of the International System of Units (SI): a double-adjustment account of scientific enquiry. Perspectives on Science, 30(4):549– 620, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/4/549/2038163/posc_a_00397.pdf.

deCourtenay:2021:IMC

[dCGLP21] Nadine de Courtenay, Fabien Grégis, Jan Lacki, and Christine Proust. Introduction: Measurement at the crossroads. Perspectives on Science, 29(6):681–700, November/December 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/29/6/681/
1972154/posc_e_00390.pdf; https://muse.jhu.edu/article/
839328.

DeCeglia:2004:BWM

[De 04] Francesco Paolo De Ceglia. The blood, the worm, the moon, the witch: Epilepsy in Georg Ernst Stahl’s pathological architecture. Perspectives on Science, 12(1):1–28, Spring 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.1ceglia.html; http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.1ceglia.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/106361404773843328.

DeCeglia:2006:RCD

[De 06] Francesco Paolo De Ceglia. Rotten corpses, a disembowelled woman, a flayed man. Images of the body from the end of the 17th to the beginning of the 19th Century. Florentine wax models in the first-hand accounts of visitors. Perspectives on Science, 14(4):417–456, Winter 2006. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.4de_ceglia.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v014/14.4de_ceglia.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2006.14.4.417.

Dede:2020:IHD

[Ded20] O. Çaglar Dede. Integrating Heather Douglas’ inductive risk framework with an account of scientific evidence: Why and how? Perspectives on Science, 28(6):737–763, November/December 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00358.

DeHart:1999:HMG

[DeH99] Scott M. DeHart. Hippocratic medicine and the Greek body image. Perspectives on Science, 7(3):349–382, Fall 1999. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v007/7.3dehart.html; http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.3dehart.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.1999.7.3.349.

Deichmann:1999:EJC

[Dei99] Ute Deichmann. The expulsion of Jewish chemists and biochemists from academia in Nazi Germany. Perspectives on Science, 7(1): 1–86, Spring 1999. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.1deichmann.html; http:/
/www.mitpressjournals.org/doi/abs/10.1162/posc.1999.7.
1.1.

Deltete:1993:WDA

[Del93] Robert Deltete. What does the anthropic principle explain? Perspectives on Science, 1(2):285–305, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

DesChene:1995:ERC

[Des95] Dennis Des Chene. Essay review: Cartesiomania: Early receptions of Descartes. Perspectives on Science, 3(4):534–581, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi.

DesChene:2000:LER

[Des00] Dennis Des Chene. On laws and ends: a response to Hattab and Menn. Perspectives on Science, 8(2):144–163, Summer 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.2des_chene.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v008/8.2des_chene.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361400568055.

DesChene:2003:LAD

[Des03] Dennis Des Chene. Life after Descartes: Régis on generation. Perspectives on Science, 11(4):410–420, Winter 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.4deschene.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v011/11.4deschene.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361403773082243.

Special Issue: Early Modern Philosophy and Biological Thought.

Desmond:2019:SGG

[Des19] Hugh Desmond. Shades of grey: Granularity, pragmatics, and non-causal explanation. Perspectives on Science, 27(1):68–87, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00300.

Detlefsen:2003:SOP

[Det03] Karen Detlefsen. Supernaturalism, occasionalism, and preformation in Malebranche. Perspectives on Science, 11(4):443– 483, Winter 2003. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v011/11.4detlefsen.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v011/11.
4detlefsen.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361403773082261. Special Issue: Early Modern

Philosophy and Biological Thought.

Deltete:1995:MAA

[DG95] Robert Deltete and Reed Guy. Much ado about nothing: a critique of quantum cosmogenesis. Perspectives on Science, 3(3):??, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

DiFate:2011:AUR

[Di 11] Victor Joseph Di Fate. Arguing for uniformity: Rethinking Lyell’s Principles of Geology. Perspectives on Science, 19 (2):136–153, Summer 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/journals/perspectives_on_science/v019/19.2.di-
fate.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.2.di-fate.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00034.

Dietrich:1993:ULI

[Die93] Michael R. Dietrich. Underdetermination and the limits of interpretative flexibility. Perspectives on Science, 1(1):109–126, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Dietrich:1996:MGG

[Die96] Michael R. Dietrich. On the mutability of genes and geneticists: The ‘Americanization’ of Richard Goldschmidt and Victor Jollos. Perspectives on Science, 4(3):321–345, Fall 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://search.ebscohost.com/login.aspx?direct=true&db=
aph&AN=6369692&site=ehost-live.

Dijksterhuis:2008:ROR

[Dij08] Fokko Jan Dijksterhuis. Reading up on the Opticks. Refashioning Newton’s theories of light and colors in Eighteenth-Century textbooks. Perspectives on Science, 16(4):309–327, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.4.dijksterhuis.html;   http://muse.
jhu.edu/journals/perspectives_on_science/v016/16.4.dijksterhuis.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2008.16.4.309.

deMelo-Martin:2014:WAD

[dMMI14] Inmaculada de Melo-Martín and Kristen Intemann. Who’s afraid of dissent?: Addressing concerns about undermining scientific consensus in public policy developments. Perspectives on Science, 22(4): 593–615, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.4.de-melo-martin.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v022/22.4.de-melo-martin.pdf; http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00151.

deOliveira:2017:TKI

[dO17] J. C. Pinto de Oliveira. Thomas Kuhn, the image of science and the image of art: The first manuscript of structure. Perspectives on Science, 25(6):746–??, November/December 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00264.

Domski:2010:KIG

[Dom10] Mary Domski. Kant on the imagination and geometrical certainty. Perspectives on Science, 18(4):409–431, Winter 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.4.domski.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v018/18.4.domski.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00016.

Donnelly:2006:IAI

[Don06] Sue Donnelly. Introduction to the archives of Imre Lakatos, 1922–1974. Perspectives on Science, 14(3):347–353, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.3donnelly.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v014/14.3donnelly.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2006.14.3.347.

Dorn:2000:SMH

[Dor00] Harold Dorn. Science, Marx, and history: Are there still research frontiers? Perspectives on Science, 8(3):223–254, Fall 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.3dorn.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v008/8.3dorn.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361400750340488.

Dorries:2001:PON

[Dör01] Matthias Dörries. Purity and objectivity in Nineteenth-Century metrology and literature. Perspectives on Science, 9(2):233– 250, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.2doerries.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v009/9.
2doerries.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/106361401317447309.

Douglas:2016:SPV

[Dou16] Heather E. Douglas. Science, policy, values: Exploring the nexus. Perspectives on Science, 24(5):475–480, September/ October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/628777;
 http://muse.jhu.edu/article/628777/pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/POSC_e_00219.

Downing:1997:LNL

[Dow97] Lisa J. Downing. Locke’s Newtonian and Lockean Newtonianism. Perspectives on Science, 5(3):285–310, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://search.ebscohost.com/login.aspx?direct=true&db=
aph&AN=6370269&site=ehost-live. Special issue: Selections

from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Downes:2009:MPU

[Dow09] Stephen M. Downes. Models, pictures, and unified accounts of representation: Lessons from aesthetics for philosophy of science. Perspectives on Science, 17(4):417–428, Winter 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v017/17.4.downes.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v017/17.4.downes.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2009.17.4.417.

deRegt:2014:VTU

[dR14] Henk W. de Regt. Visualization as a tool for understanding. Perspectives on Science, 22(3):377–396, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
3.de-regt.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.3.de-regt.pdf.

Dorato:2018:NRF

[DR18] Mauro Dorato and Emanuele Rossanese. The nature of representation in Feynman diagrams. Perspectives on Science, 26(4):443–458, July/August 2018. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00282.

deRobertis:2024:TST

[dR24] Sofia Sandreschi de Robertis. Tics, slips of the tongue and habit between Maine de Biran and Victor Egger. Perspectives on Science, 32(1):91–101, January-February 2024. CODEN PRSIEU.

ISSN 1063-6145 (print), 1530-9274 (electronic). URL https:/

/direct.mit.edu/posc/article-pdf/32/1/91/2210166/posc_
a_00600.pdf.

Dresow:2023:URE

[Dre23] Max Dresow. Uniformitarianism re-examined, or the present is the key to the past, except when it isn’t (and even then it kind of is). Perspectives on Science, 31(4):405–436, ???? 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

https://muse.jhu.edu/pub/6/article/902861.

deRegt:2014:ISV

[dRP14] Henk W. de Regt and Wendy S. Parker. Introduction: Simulation, visualization, and scientific understanding. Perspectives on Science, 22(3):311–317, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/journals/perspectives_on_science/v022/22.3.de-
regt01.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.3.de-regt01.pdf.

Dietrich:2007:MUS

[DS07] Michael R. Dietrich and Robert A. Skipper. Manipulating underdetermination in scientific controversy: The case of the molecular clock. Perspectives on Science, 15(3):295–326, Fall 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.3dietrich.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v015/15.3dietrich.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2007.15.3.295.

Dent:2017:UFD

[DS17] Rosanna Dent and Ricardo Ventura Santos. “An unusual and fast disappearing opportunity”: Infectious disease, indigenous populations, and new biomedical knowledge in Amazonia, 1960– 1970. Perspectives on Science, 25(5):585–605, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00255.

Duchesneau:1998:LTS

[Duc98] François Duchesneau. Leibniz’s theoretical shift in the Phoranomus and Dynamica de Potentia. Perspectives on Science, 6(1–2):77–109, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v006/6.1duchesneau.html. Special

Issue: Leibniz and the Sciences.

Duchesneau:2003:LMA

[Duc03] François Duchesneau. Leibniz’s model for analyzing organic phenomena. Perspectives on Science, 11(4):378–409, Winter 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.4duchesneau.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v011/11.4duchesneau.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

106361403773082234. Special Issue: Early Modern Philosophy and Biological Thought.

Ducheyne:2010:FQS

[Duc10] Steffen Ducheyne. Fundamental questions and some new answers on philosophical, contextual and scientific Whewell: Some reflections on recent Whewell scholarship and the progress made therein. Perspectives on Science, 18(2):242–272, Summer 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.2.ducheyne.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v018/18.2.ducheyne.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2010.18.2.242.

Dupre:2004:UCG

[Dup04] John Dupré. Understanding contemporary genomics. Perspectives on Science, 12(3):320–338, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v012/12.
3dupre.html; http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.3dupre.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/1063614042795435.

Dupre:2008:NTP

[Dup08] Sven Dupré. Newton’s telescope in print: The role of images in the reception of Newton’s instrument. Perspectives on Science, 16(4):328–359, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v016/16.4.dupre.html;

http://muse.jhu.edu/journals/perspectives_on_science/v016/
16.4.dupre.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2008.16.4.328.

Dupre:2012:CPS

[Dup12] John Dupré. Comments on Philosophy of Science after Feminism, by Janet Kourany. Perspectives on Science, 20(3): 310–319, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v020/20.3.dupre.html;
http://muse.jhu.edu/journals/perspectives_on_science/v020/
20.3.dupre.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00074.

Durant:2010:PPM

[Dur10] Darrin Durant. Public participation in the making of science policy. Perspectives on Science, 18(2):189–225, Summer 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.2.durant.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v018/18.
2.durant.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2010.18.2.189.

Earl:2013:MAC

[Ear13] Graeme Earl. Modeling in archaeology: Computer graphic and other digital pasts. Perspectives on Science, 21(2):226– 244, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.2.earl.html; http://muse.
jhu.edu/journals/perspectives_on_science/v021/21.2.earl.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

POSC_a_00096. Special Issue on Modeling in the Social Sciences.

Eck:2016:SCS

[Eck16] David Eck. Social coordination in scientific communities. Perspectives on Science, 24(6):770–800, November/December 2016. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/635026;   http://muse.
jhu.edu/article/635026/pdf.

Eigi:2020:ERN

[Eig20] Jaana Eigi. Are experts representative of non-experts? Elective modernism, aspects of representation, and the argument from inductive risk. Perspectives on Science, 28(4):459–481, July/August 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00347.

Elgin:2014:FTE

[Elg14] Catherine Z. Elgin. Fiction as thought experiment. Perspectives on Science, 22(2):221–241, Summer 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
2.elgin.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.2.elgin.pdf.

Elkins:1999:LIA

[Elk99] James Elkins. Logic and images in art history. Perspectives on Science, 7(2):151–180, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v007/7.
2elkins.html; http://www.mitpressjournals.org/doi/abs/
10.1162/posc.1999.7.2.151.

Elliott:2000:CCP

[Ell00] Kevin Elliott. Conceptual clarification and policy-related science: The case of chemical hormesis. Perspectives on Science, 8(4):346–366, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v008/8.4elliott.html;
http://muse.jhu.edu/journals/perspectives_on_science/v008/
8.4elliott.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361400753373731.

Elliott:2016:SSD

[Ell16] Kevin C. Elliott. Standardized study designs, value judgments, and financial conflicts of interest in research. Perspectives on Science, 24(5):529–551, September/October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/628780;   http://muse.jhu.edu/article/
628780/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00222.

Fagan:2011:SES

[Fag11] Melinda B. Fagan. Social experiments in stem cell biology. Perspectives on Science, 19(3):235–262, Fall 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v019/19.
3.fagan.html; http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.3.fagan.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00045.

Feest:2007:SES

[Fee07] Uljana Feest. Science and experience/science of experience: Gestalt psychology and the anti-metaphysical project of the Aufbau. Perspectives on Science, 15(1):1–25, Spring 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v015/15.
1feest.html; http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.1feest.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2007.15.1.1.

Fellows:2017:TSB

[Fel17] Jennifer Jill Fellows. Trust without shared belief: Pluralist realism and polar bear conservation. Perspectives on Science, 25 (1):36–66, January/February 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/647988;   http://muse.jhu.edu/article/
647988/pdf.

FernandezPinto:2023:MCB

[Fer23] Manuela Fernández Pinto. Methodological and cognitive biases in science: Issues for current research and ways to counteract them. Perspectives on Science, 31(5):535–554, September-October 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
5/535/2158903/posc_a_00589.pdf.

Fritzman:2012:SHE

[FG12] J. M. Fritzman and Molly Gibson. Schelling, Hegel, and evolutionary progress. Perspectives on Science, 20(1):105–128, Spring 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.1.fritzman.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.1.fritzman.

pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00058.

Fisher:2021:EME

[FGS21] Grant Fisher, Axel Gelfert, and Friedrich Steinle. Exploratory models and exploratory modeling in science: Introduction. Perspectives on Science, 29(4):355–358, July/August 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/29/4/355/
1959230/posc_e_00374.pdf; https://muse.jhu.edu/article/
804402.

Farrell:2007:ASDa

[FH07a] Robert P. Farrell and C. A. (Clifford Alan) Hooker. Applying self-directed anticipative learning to science I: Agency, error, and the interactive exploration of possibility space in early ape-language research. Perspectives on Science, 15(1):87–124, Spring 2007. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v015/15.1farrell.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v015/15.
1farrell.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2007.15.1.87.

Farrell:2007:ASDb

[FH07b] Robert P. Farrell and C. A. (Clifford Alan) Hooker. Applying selfdirected anticipative learning to science II: Learning how to learn across a revolution in early ape language research. Perspectives on Science, 15(2):222–255, Summer 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v015/15.2farrell.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v015/15.2farrell.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2007.15.2.222.

Fisher:2003:EMP

[Fis03a] Saul Fisher. Early modern philosophy and biological thought. Perspectives on Science, 11(4):373–377, Winter 2003. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/106361403773082225.

Fisher:2003:GAA

[Fis03b] Saul Fisher. Gassendi’s atomist account of generation and heredity in plants and animals. Perspectives on Science, 11(4):484– 512, Winter 2003. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v011/11.4fisher02.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v011/11.
4fisher02.pdf;   http://www.mitpressjournals.org/doi/abs/

10.1162/106361403773082270. Special Issue: Early Modern Philosophy and Biological Thought.

Fleming:2013:NLP

[Fle13] Leonore Fleming. The notion of limited perfect adaptedness in Darwin’s principle of divergence. Perspectives on Science, 21(1): 1–22, Spring 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.1.fleming.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v021/21.
1.fleming.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00084.

Formica:2010:NMS

[For10] Giambattista Formica. Von Neumann’s methodology of science: From incompleteness theorems to later foundational reflections. Perspectives on Science, 18(4):480–499, Winter 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.4.formica.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v018/18.4.formica.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00019.

Franklin:1993:DPJ

[Fra93a] Allan Franklin. Discovery, pursuit, and justification. Perspectives on Science, 1(2):252–284, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Franklin:1993:REE

[Fra93b] Allan Franklin. Review essay: Experimental questions. Perspectives on Science, 1(1):127–146, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Franklin:1995:RDR

[Fra95] Allan Franklin. The resolution of discordant results. Perspectives on Science, 3(3):346–420, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Franklin:1997:C

[Fra97] Allan Franklin. Calibration. Perspectives on Science, 5(1):31–80, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6664440&site=ehost-live.

Franklin:2010:GWN

[Fra10] Allan Franklin. Gravity waves and neutrinos: The later work of Joseph Weber. Perspectives on Science, 18(2):119–151, Summer 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.2.franklin.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v018/18.2.franklin.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2010.18.2.119.

Fransen:2019:AVL

[Fra19] Sietske Fransen. Antoni van Leeuwenhoek, his images and draughtsmen. Perspectives on Science, 27(3):485–544, May/June 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00314.

Freudenthal:2005:HGT

[Fre05] Gideon Freudenthal. The Hessen–Grossman thesis: an attempt at rehabilitation. Perspectives on Science, 13(2):166– 193, Summer 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v013/13.2freudenthal.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v013/13.2freudenthal.pdf; http://www.mitpressjournals.
org/doi/abs/10.1162/106361405774270575.  Special Issue:

Technoscientific Productivity, Part 1.

Friedman:2021:PPS

[Fri21] Michael Friedman. Paper, plaster, strings: Exploratory material mathematical models between the 1860s and 1930s. Perspectives on Science, 29(4):436–467, July/August 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/4/436/1959248/posc_
a_00378.pdf; https://muse.jhu.edu/article/804406.

Fehige:2014:I

[FS14a] Yiftach Fehige and Michael T. Stuart. Introduction. Perspectives on Science, 22(2):167–178, Summer 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v022/22.2.fehige.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v022/22.2.fehige.pdf.

Fehige:2014:OPT

[FS14b] Yiftach Fehige and Michael T. Stuart. On the origins of the philosophy of thought experiments: The forerun. Perspectives on Science, 22(2):179–220, Summer 2014. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.2.fehige01.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v022/22.2.fehige01.
pdf.

Fuchs:1993:POM

[Fuc93] Stephan Fuchs. Positivism is the organizational myth of science. Perspectives on Science, 1(1):1–23, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Garber:1998:LS

[GA98] Daniel Garber and Roger Ariew, editors. Leibniz and the sciences. MIT Press, Cambridge, MA, USA, 1998. ISSN 1063-6145 (print), 1530-9274 (electronic). i–iv and 1–208 pp. Perspect. Sci. 6 (1998), no. 1–2.

Gage:2008:SDN

[Gag08] John Gage. Signs of disharmony: Newton’s Opticks and the artists. Perspectives on Science, 16(4):360–377, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.4.gage.html; http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.4.gage.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2008.16.4.360.

Galison:1999:RIL

[Gal99] Peter Louis Galison. Reflections on Image and Logic: a material culture of microphysics. Perspectives on Science, 7(2): 255–284, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v007/7.2galison.html;
http://www.mitpressjournals.org/doi/abs/10.1162/posc.1999.
7.2.255.

Gal:2002:CPI

[Gal02] Ofer Gal. Constructivism for philosophers (be it a remark on realism). Perspectives on Science, 10(4):523–549, Winter 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.4gal.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v010/10.4gal.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361402322288075.

Gallo:2009:DOF

[Gal09] Jason Gallo. The discursive and operational foundations of the national Nanotechnology initiative in the history of the national science foundation. Perspectives on Science, 17(2): 174–211, Summer 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v017/17.2.gallo.html;
http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.2.gallo.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2009.17.2.174.

Garber:1995:ECC

[Gar95a] Daniel Garber. Experiment, community and the constitution of nature in the 17th Century. Perspectives on Science, 3(2):173– 205, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Garber:1995:IAO

[Gar95b] Daniel Garber. Introduction: Apples, oranges, and the role of Gassendi’s atomism in Seventeenth-Century science. Perspectives on Science, 3(4):425–428, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi.

Garber:2001:DSR

[Gar01] Daniel Garber. Descartes and the scientific revolution: Some Kuhnian reflections. Perspectives on Science, 9(4):405–422, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.4garber.html; http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.4garber.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361401760375802.

Garber:2004:FSR

[Gar04] Daniel Garber. On the frontlines of the scientific revolution: How Mersenne learned to love Galileo. Perspectives on Science, 12(2):135–163, Summer 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v012/12.2garber.html;
http://muse.jhu.edu/journals/perspectives_on_science/v012/
12.2garber.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361404323119853. Special Issue: Galileo in

Paris.

Garrod:2012:FNH

[Gar12] Raphaële Garrod. On fish: Natural history as spiritual materia medica: Calvinist pastoralism in Pierre Viret’s Instruction Chrestienne (1564). Perspectives on Science, 20(2):227– 245, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.2.garrod.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v020/20.
2.garrod.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00064.

Garau:2021:WWF

[Gar21] Rodolfo Garau. Who was the founder of empiricism after all? Gassendi and the ‘logic’ of Bacon. Perspectives on Science, 29(3):327–354, May/June 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/3/327/1923001/posc_
a_00371.pdf; https://muse.jhu.edu/article/796018.

Gaukroger:2000:RMT

[Gau00] Stephen Gaukroger. The role of matter theory in Baconian and Cartesian cosmologies. Perspectives on Science, 8(3):201–222, Fall 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.3gaukroger.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v008/8.3gaukroger.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361400750340479.

Georgescu:2017:DPG

[Geo17] Laura Georgescu. The disponent power in Gilbert’s De Magnete: From attraction to alignment. Perspectives on Science, 25(2):149– 176, March/April 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00239.

Gervais:2013:NCV

[Ger13] Raoul Gervais. Non-cognitive values and objectivity in scientific explanation: Egalitarianism and the case of the Movius line. Perspectives on Science, 21(4):429–452, Winter 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v021/21.
4.gervais.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.4.gervais.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00111.

Gingras:2011:UAS

[GG11] Yves Gingras and Alexandre Guay. The uses of analogies in Seventeenth and Eighteenth Century science. Perspectives on Science, 19(2):154–191, Summer 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v019/19.
2.gingras.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.2.gingras.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00035.

Goldstein:2005:KMO

[GH05] Bernard R. Goldstein and Giora Hon. Kepler’s move from orbs to orbits: documenting a revolutionary scientific concept. Perspectives on Science, 13(1):74–111, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.1goldstein.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v013/13.1goldstein.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
1063614053714126.

Gieryn:1994:REO

[Gie94] Thomas F. Gieryn. Review essay: Objectivity for these times. Perspectives on Science, 2(3):324–349, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Giere:2012:NPP

[Gie12] Ronald N. Giere. A new program for the philosophy of science? Perspectives on Science, 20(3):339–343, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.3.giere.html; http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.3.giere.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00073.

Giglioni:2012:PAT

[Gig12] Guido Giglioni. Philosophy according to Tacitus: Francis Bacon and the inquiry into the limits of human self-delusion. Perspectives on Science, 20(2):159–182, Summer 2012. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.2.giglioni.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.2.giglioni.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00061.

Gingerich:2002:KTN

[Gin02] Owen Gingerich. Kepler then and now. Perspectives on Science, 10(2):228–240, Summer 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v010/10.
2gingerich.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.2gingerich.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361402321147540.

Ginev:2016:HPS

[Gin16] Dimitri Ginev. Hermeneutic perspectives on science in Fleck’s work and hermeneutic critique of constructivist epistemology. Perspectives on Science, 24(2):228–253, March/April 2016. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00202.

Giovanelli:2017:SSP

[Gio17] Marco Giovanelli. The sensation and the stimulus: Psychophysics and the prehistory of the Marburg School. Perspectives on Science, 25(3):287–323, May/June 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00244.

Gissis:2002:LNC

[Gis02] Snait Gissis. Late Nineteenth-Century Lamarckism and French sociology. Perspectives on Science, 10(1):69–122, Spring 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.1gissis.html; http://muse.jhu.edu/journals/
perspectives_on_science/v010/10.1gissis.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/106361402762674807.

Gissis:2009:IBS

[Gis09] Snait Gissis. Interactions between social and biological thinking: The case of Lamarck. Perspectives on Science, 17(3):237– 306, Fall 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.3.gissis.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v017/17.
3.gissis.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2009.17.3.237.

Glouberman:1995:GGF

[Glo95] M. Glouberman. Gods, giants, fractals, and the geometry of Early Modernity: Descartes, Gassendi, and the rise of science. Perspectives on Science, 3(4):480–519, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi. See reply [Len95].

Giordani:2021:TDT

[GM21] Alessandro Giordani and Luca Mari. On theory dependence of truth in measurement. Perspectives on Science, 29(6):757– 781, November/December 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/6/757/1972162/posc_
a_00393.pdf; https://muse.jhu.edu/article/839324.

Greco:2021:CMC

[GMHP21] Gabriele Greco, Virginia Mastellari, Chris Holland, and Nicola M. Pugno. Comparing modern and classical perspectives on spider silks and webs. Perspectives on Science, 29(2):133–156, March/ April 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL https://muse.jhu.edu/article/787906.

Goeminne:2011:OUT

[Goe11] Gert Goeminne. Once upon a time I was a nuclear physicist: What the politics of sustainability can learn from the nuclear laboratory. Perspectives on Science, 19(1):1–31, Spring 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.1.goeminne.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v019/19.1.goeminne.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00024.

Goldstein:1997:PAL

[Gol97] Bernard R. Goldstein. The physical astronomy of Levi ben Gerson. Perspectives on Science, 5(1):1–30, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://search.ebscohost.com/login.aspx?direct=true&db=
aph&AN=6664438&site=ehost-live.

Goldberg:2013:TPX

[Gol13] Daniel S. Goldberg. The transformative power of X-rays in U.S. scientific & medical litigation: Mechanical objectivity in Smith v. Grant (1896). Perspectives on Science, 21(1):23–57, Spring 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.1.goldberg.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.1.goldberg.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00085.

Goldberg:2016:WHA

[Gol16a] Benjamin Goldberg. William Harvey on anatomy and experience. Perspectives on Science, 24(3):305–323, May/June 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/618107;   http://muse.
jhu.edu/article/618107/pdf.

Goldenberg:2016:PMS

[Gol16b] Maya J. Goldenberg. Public misunderstanding of science?: Reframing the problem of vaccine hesitancy. Perspectives on Science, 24(5):552–581, September/October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/628781;   http://muse.jhu.edu/article/
628781/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00223.

Gooding:2006:PFT

[Goo06] David Gooding. From phenomenology to field theory: Faraday’s visual reasoning. Perspectives on Science, 14(1):40–65, Spring 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.1gooding.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v014/14.
1gooding.pdf;   http://www.mitpressjournals.org/doi/abs/

10.1162/posc.2006.14.1.40. Special Issue: Cognitive Psychology and Faraday’s Experiments.

Goodwin:2018:CCC

[Goo18] William Goodwin. Conflicting conceptions of construction in Kant’s philosophy of geometry. Perspectives on Science, 26 (1):97–118, January/February 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00269.

Gorman:1996:MFC

[Gor96] Michael John Gorman. A matter of faith? Christoph Scheiner, Jesuit censorship, and the trial of Galileo. Perspectives on Science, 4(3):283–320, Fall 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6369660&site=ehost-live.

Gorham:2005:MRC

[Gor05] Geoffrey Gorham. The metaphysical roots of Cartesian physics: The law of rectilinear motion. Perspectives on Science, 13(4): 431–451, Winter 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v013/13.4gorham.html;
http://muse.jhu.edu/journals/perspectives_on_science/v013/

13.4gorham.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361405775466126.

Goulding:2022:GGT

[Gou22] Robert Goulding. Geometry and the gods: Theurgy in Proclus’s commentary on the first book of Euclid’s Elements. Perspectives on Science, 30(3):358–406, May/June 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/3/358/2023837/posc_
a_00420.pdf.

Grafton:2000:SMR

[Gra00] Anthony Grafton. Starry messengers: Recent work in the history of Western astrology. Perspectives on Science, 8(1): 70–83, Spring 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v008/8.1grafton.html;
http://muse.jhu.edu/journals/perspectives_on_science/v008/
8.1grafton.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361400568028.

Gradmann:2001:ICP

[Gra01] Christoph Gradmann. Isolation, contamination, and pure culture: Monomorphism and polymorphism of pathogenic micro-organisms as research problem 1860–1880. Perspectives on Science, 9(2):147– 172, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.2gradmann.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v009/9.
2gradmann.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/106361401317447264.

Grant:2004:SIM

[Gra04] Edward Grant. Scientific imagination in the middle ages. Perspectives on Science, 12(4):394–423, Winter 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v012/12.
4grant.html; http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.4grant.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/1063614042776021.

Granada:2010:MCG

[Gra10] Miguel A. Granada. Mersenne’s critique of Giordano Bruno’s conception of the relation between God and the Universe: a reappraisal. Perspectives on Science, 18(1):26–49, Spring 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.1.granada.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v018/18.1.granada.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2010.18.1.26.

Grene:1993:ACT

[Gre93a] Marjorie Grene. Aristotelico–Cartesian themes in natural philosophy: Some Seventeenth-Century cases. Perspectives on Science, 1 (1):66–87, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Grene:1993:RER

[Gre93b] Marjorie Grene. Review essay: Recent biographies of Darwin: The complexity of context. Perspectives on Science, 1(4):659–675, Winter 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Grene:1997:REC

[Gre97] Marjorie Grene. Review essay: Current issues in the philosophy of biology. Perspectives on Science, 5(2):255–281, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=10780097&site=ehost-live.

Grene:2000:RWA

[Gre00] Marjorie Glicksman Grene. Recent work on Aristotelian biology. Perspectives on Science, 8(4):444–459, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.4grene.html; http://muse.jhu.edu/journals/
perspectives_on_science/v008/8.4grene.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361400753373777.

Greif:2021:EMM

[Gre21] Hajo Greif. Exploring minds: Modes of modeling and simulation in artificial intelligence. Perspectives on Science, 29(4):409–435, July/August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/409/1962220/posc_a_00377.pdf; https:/
/muse.jhu.edu/article/804405.

GuerreroMcManus:2017:PM

[Gue17] Siobhan F. Guerrero McManus. Populations of misrecognition. Perspectives on Science, 25(5):712–??, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00262.

Guicciardini:2013:HDN

[Gui13]  Niccolò Guicciardini.  Harper and Ducheyne on Newton.

Perspectives on Science, 21(4):463–481, Winter 2013. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.4.guicciardini.html;   http://muse.
jhu.edu/journals/perspectives_on_science/v021/21.4.guicciardini.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00113.

Gundersen:2020:VFY

[Gun20] Torbjørn Gundersen. Value-free yet policy-relevant? The normative views of climate scientists and their bearing on philosophy. Perspectives on Science, 28(1):89–118, January/February 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://www.mitpressjournals.org/doi/abs/10.1162/
posc_a_00334;   https://muse.jhu.edu/article/749744.

Gurka:2006:MLI

[Gur06] Dezső Gurka. A missing link: The infuence of László Kalmár’s empirical view on Lakatos’ philosophy of mathematics. Perspectives on Science, 14(3):263–281, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
3gurka.html; http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.3gurka.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.3.263.

Gutmann:2021:DFS

[Gut21] Mathias Gutmann. The disappearance of form? Some methodological considerations on a lost conceptual dimension in biology. Perspectives on Science, 29(5):666–680, September/ October 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/5/666/1964247/posc_a_00389.pdf; https:/
/muse.jhu.edu/article/810916.

Grune-Yanoff:2013:RBT

[GY13] Till Grüne-Yanoff. Relations between theory and model in psychology and economics. Perspectives on Science, 21(2):196–201, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.2.grune-yanoff.html;
http://muse.jhu.edu/journals/perspectives_on_science/v021/
21.2.grune-yanoff.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_a_00093. Special Issue on Modeling in

the Social Sciences.

Hacques:2024:BSA

[Hac24] Romain Hacques. The Biranian spiritualism of Alexis Bertrand: a philosophy of one’s own body? Perspectives on Science, 32(1): 70–90, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/32/1/70/2210170/posc_a_00599.pdf.

Hamilton:2001:SPC

[Ham01] Kelly Ann Hamilton. Some philosophical consequences of Wittgenstein’s aeronautical research. Perspectives on Science, 9(1):1– 37, Spring 2001. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.1hamilton.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v009/9.
1hamilton.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/10636140152947777.

Hamilton:2002:DPM

[Ham02] Kelly Ann Hamilton. Darstellungen in The Principles of Mechanics and the Tractatus: The representation of objects in relation in Hertz and Wittgenstein. Perspectives on Science, 10(1):28–68, Spring 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v010/10.
1hamilton.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.1hamilton.pdf.

Hamm:2016:MHS

[Ham16] Elizabeth Hamm. Modeling the heavens: Sphairopoiia and Ptolemy’s planetary hypotheses. Perspectives on Science, 24 (4):416–424, July/August 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/625738;   http://muse.jhu.edu/article/
625738/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00214.

Hands:2006:PFE

[Han06] D. Wade Hands. Priority fights in economic science: Paradox and resolution. Perspectives on Science, 14(2):215–231, Summer 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.2hands.html; http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.2hands.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2006.14.2.215.

Hansson:2022:CUQ

[Han22] Sven Ove Hansson. Can uncertainty be quantified? Perspectives on Science, 30(2):210–236, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/2/210/2005008/posc_
a_00412.pdf.

Harris:1994:CRS

[Har94a] R. Allen Harris. The Chomskyian revolution I: Syntax, semantics, and science. Perspectives on Science, 2(1):38–75, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Harris:1994:CRI

[Har94b] R. Allen Harris. The Chomskyian revolution II: Sturm und Drang. Perspectives on Science, 2(2):176–230, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Harley:1996:REA

[Har96] David Harley. Review essay: Anglo–American perspectives on early modern medicine: Society, religion, and science. Perspectives on Science, 4(3):346–386, Fall 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6369699&site=ehost-live.

Harwood:1997:REG

[Har97] Jonathan Harwood. Review essay: German science and technology under National Socialism. Perspectives on Science, 5(1):128–151, Spring 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6664449&site=ehost-live.

Harwood:2005:CAP

[Har05a] Jonathan Harwood. Comments on Andrew Pickering’s paper. Perspectives on Science, 13(3):411–415, Fall 2005. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.3harwood02.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v013/13.3harwood02.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

106361405774287991. Special Issue: Technoscientific Productivity, Part 2.

Harwood:2005:GTC

[Har05b] Jonathan Harwood. On the genesis of technoscience: a case study of German agricultural education. Perspectives on Science, 13(3):329–351, Fall 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v013/13.
3harwood01.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.3harwood01.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361405774288017.  Special Issue:

Technoscientific Productivity, Part 2.

Harris:2009:AGR

[Har09] Randy Harris. Alan Gross and the rhetoric of science. Perspectives on Science, 17(3):346–380, Fall 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v017/17.3.harris.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v017/17.3.harris.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2009.17.3.346.

Harrison:2012:FBN

[Har12] Peter Harrison. Francis Bacon, natural philosophy, and the cultivation of the mind. Perspectives on Science, 20(2):139–158, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.2.harrison.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.2.harrison.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00060.

Hatfield:1997:WPS

[Hat97] Gary Hatfield. Wundt and psychology as science: Disciplinary transformations. Perspectives on Science, 5(3):349–382, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6370277&site=ehost-live. Special is-

sue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Hattab:2000:PSC

[Hat00] Helen Hattab. The problem of secondary causation in Descartes: a response to Des Chene. Perspectives on Science, 8(2):93– 118, Summer 2000. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v008/8.2hattab.html; http://muse.
jhu.edu/journals/perspectives_on_science/v008/8.2hattab.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
106361400568037.

Hattab:2016:AAC

[Hat16] Helen Hattab. Aristotelianism and atomism combined: Gorlaeus on knowledge of universals. Perspectives on Science, 24(3):285–304, May/June 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/618106;
 http://muse.jhu.edu/article/618106/pdf.

Halpern:2022:SED

[HE22] Megan K. Halpern and Kevin C. Elliott. Science as experience: a Deweyan model of science communication. Perspectives on Science, 30(4):621–656, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/30/4/621/2038161/posc_a_00398.pdf.

Hedfors:2007:FC

[Hed07] Eva Hedfors. Fleck in context. Perspectives on Science, 15(1): 49–86, Spring 2007. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v015/15.1hedfors.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v015/15.
1hedfors.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2007.15.1.49.

Heidelberger:2007:NKC

[Hei07] Michael Heidelberger. From neo-Kantianism to critical realism: Space and the mind-body problem in Riehl and Schlick. Perspectives on Science, 15(1):26–48, Spring 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v015/15.
1heidelberger.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.1heidelberger.pdf; http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2007.15.1.26.

Heis:2010:CPB

[Hei10] Jeremy Heis. “Critical philosophy begins at the very point where logistic leaves off”: Cassirer’s response to Frege and Russell. Perspectives on Science, 18(4):383–408, Winter 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
4.heis.html; http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.4.heis.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00015.

Heilbron:2012:RWG

[Hei12] J. L. Heilbron. Robert Westman on Galileo and related matters. Perspectives on Science, 20(3):379–388, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.3.heilbron.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.3.heilbron.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00072.

Heilbron:2013:RW

[Hei13] J. L. Heilbron. Reply to Westman. Perspectives on Science, 21 (3):386, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.3.heilbron.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v021/21.
3.heilbron.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_c_00106.

Heilbron:2015:RTP

[Hei15] J. L. Heilbron. Roman thought police and early-modern astrology. Perspectives on Science, 23(2):228–240, Summer 2015.

CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.2.heilbron.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v023/23.2.heilbron.
pdf.

Henderson:2019:RHV

[Hen19] Felicity Henderson. Robert Hooke and the visual world of the early Royal Society. Perspectives on Science, 27(3):395–434, May/ June 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://www.mitpressjournals.org/doi/abs/
10.1162/posc_a_00312.

Hicks:2017:GMC

[Hic17] Daniel J. Hicks. Genetically modified crops, inclusion, and democracy. Perspectives on Science, 25(4):488–520, July/August 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00251.

Hlabangane:2018:CMS

[Hla18] Nokuthula Hlabangane. Can a methodology subvert the logics of its principal? Decolonial meditations. Perspectives on Science, 26 (6):658–693, November/December 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00293.

Hadden:1996:OPR

[HO96] Richard W. Hadden and Michael A. Overington. Ontological porcupine: The road to hegemony and back in science studies. Perspectives on Science, 4(1):1–23, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6663242&site=ehost-live.

Hoffmann:2001:DDP

[Hof01] Christoph Hoffmann. The design of disturbance: Physics institutes and physics research in Germany, 1870–1910. Perspectives on Science, 9(2):173–195, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v009/9.
2hoffmann02.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.2hoffmann02.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361401317447273.

Hof:2018:CGM

[Hof18] Barbara E. Hof. The cybernetic “general model theory”: Unifying science or epistemic change? Perspectives on Science, 26(1):76–96, January/February 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00268.

Hoffmann:2020:MD

[Hof20] Christoph Hoffmann. Multiple data. Perspectives on Science, 28 (6):684–699, November/December 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00356.

Hofler:2021:MOM

[Höf21] Carolin Höfler. Model operations: Morphogenesis and the design process. Perspectives on Science, 29(5):602–626, September/October 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/5/602/1964258/posc_a_00386.pdf; https:/
/muse.jhu.edu/article/810913.

Holowchak:2012:WFA

[Hol12] M. Andrew Holowchak. When Freud (almost) met Chaplin: The science behind Freud’s “Especially Simple, Transparent Case”. Perspectives on Science, 20(1):44–74, Spring 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.1.holowchak.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.1.holowchak.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00056.

Hong:1998:UOT

[Hon98] Sungook Hong. Unfaithful offspring? Technologies and their trajectories. Perspectives on Science, 6(3):259–287, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v006/6.3hong.html.

Hooker:2018:NPS

[Hoo18] Cliff Hooker. A new problem-solving paradigm for philosophy of science. Perspectives on Science, 26(2):266–291, March/April 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00275.

Hopp:2014:ET

[Hop14] Walter Hopp. Experiments in thought. Perspectives on Science, 22(2):242–263, Summer 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
2.hopp.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.2.hopp.pdf.

Horn:2005:ESS

[Hor05] Jeff Horn. Enlightenment science and the state in Revolutionary France: The legacy of Charles Coulston Gillispie. Perspectives on Science, 13(1):112–132, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v013/13.
1horn.html; http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.1horn.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/1063614053714090.

Howard:1993:WER

[How93] Don Howard. Was Einstein really a realist? Perspectives on Science, 1(2):204–251, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Howard:1997:I

[How97] Don Howard. Introduction. Perspectives on Science, 5(3):iii–iv, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special issue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Hoffmann:2001:SMD

[HS01] Christoph Hoffmann and Jutta Schickore. Secondary matters: On disturbances, contamination, and waste as objects of research. Perspectives on Science, 9(2):123–125, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/106361401317447246.

Hangel:2017:SCG

[HS17] Nora Hangel and Jutta Schickore. Scientists’ conceptions of good research practice. Perspectives on Science, 25(6):766–??, November/December 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00265.

Hudson:2016:WWS

[Hud16] Robert Hudson. Why we should not reject the value-free ideal of science. Perspectives on Science, 24(2):167–191, March/April 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00199.

Hughes:2006:TPB

[Hug06] R. I. G. Hughes. Theoretical practice: the Bohm-Pines quartet. Perspectives on Science, 14(4):457–524, Winter 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.4hughes.html; http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.4hughes.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/posc.2006.14.
4.457.

Hull:1998:SSS

[Hul98] David L. Hull. Studying the study of science scientifically. Perspectives on Science, 6(3):209–231, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v006/6.3hull.html.

Huneman:2019:MLH

[Hun19] Philippe Huneman. The multifaceted legacy of the Human Genome Program for evolutionary biology: an epistemological perspective. Perspectives on Science, 27(1):117–152, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00302.

Hutchison:1996:WDP

[Hut96] Keith Hutchison. Why does Plato urge rulers to study astronomy? Perspectives on Science, 4(1):24–58, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://search.ebscohost.com/login.aspx?direct=true&db=
aph&AN=6663244&site=ehost-live.

Hutchins:2016:DDP

[Hut16] Barnaby R. Hutchins. Does Descartes have a principle of life?: Hierarchy and interdependence in Descartes’s physiology. Perspectives on Science, 24(6):744–769, November/December 2016. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/635025;   http://muse.
jhu.edu/article/635025/pdf.

Hyder:2003:FCH

[Hyd03] David Jalal Hyder. Foucault, Cavaillès, and Husserl on the historical epistemology of the sciences. Perspectives on Science, 11(1): 107–129, Spring 2003. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v011/11.1hyder.html; http://muse.
jhu.edu/journals/perspectives_on_science/v011/11.1hyder.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

106361403322286724. Special issue: History of science and philosophy of science (Montreal, QC, 2002).

Hyun:2022:NHA

[Hyu22] Jaehwan Hyun. In the name of human adaptation: Japanese American “hybrid children” and racial anthropology in postwar Japan. Perspectives on Science, 30(1):167–193, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
1/167/1987188/posc_a_00406.pdf.

Hon:2009:KOP

[HZ09] Giora Hon and Yaakov Zik. Kepler’s Optical Part of Astronomy (1604): Introducing the ecliptic instrument. Perspectives on Science, 17(3):307–345, Fall 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v017/17.
3.hon.html; http://muse.jhu.edu/journals/perspectives_
on_science/v017/17.3.hon.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2009.17.3.307.

Israel-Jost:2016:CIP

[IJ16] Vincent Israel-Jost. Computer image processing: an epistemological aid in scientific investigation. Perspectives on Science, 24(6):

669–695, November/December 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/635022;   http://muse.jhu.edu/article/
635022/pdf.

Inkpen:2017:DND

[Ink17] S. Andrew Inkpen. Demarcating nature, defining ecology: Creating a rationale for the study of nature’s “primitive conditions”. Perspectives on Science, 25(3):355–392, May/June 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00246.

Jackson:2001:MSD

[Jac01] Myles W. Jackson. Music and science during the scientific revolution. Perspectives on Science, 9(1):106–115, Spring 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.1jackson.html; http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.1jackson.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/10636140152947812.

Jackson:2008:PSB

[Jac08] Myles W. Jackson. Putting the subject back into color: Accessibility in Goethe’s Zur Farbenlehre. Perspectives on Science, 16(4):378–391, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v016/16.
4.jackson.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.4.jackson.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2008.16.4.378.

Jacobs:2010:JBC

[Jac10] Struan Jacobs. J. B. Conant’s other assistant: Science as depicted by Leonard K. Nash, including reference to Thomas Kuhn. Perspectives on Science, 18(3):328–351, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
3.jacobs.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.3.jacobs.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00010.

Jackson:2015:HGP

[Jac15a] Myles W. Jackson. How gene patents are challenging intellectual property law: The history of the CCR5 gene patent. Perspectives on Science, 23(1):80–105, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v023/23.
1.jackson01.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.1.jackson01.pdf.

Jackson:2015:I

[Jac15b] Myles W. Jackson. Introduction. Perspectives on Science, 23(1): 1–12, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.1.jackson.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v023/23.
1.jackson.pdf.

Jackson:2021:UMD

[Jac21a] Rebecca L. Jackson. “The uncertain method of drops”: How a non-uniform unit survived the century of standardization. Perspectives on Science, 29(6):802–841, November/ December 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/6/802/1972176/posc_a_00395.pdf; https:/
/muse.jhu.edu/article/839326.

Jacoby:2021:ARN

[Jac21b] Franklin Jacoby. Acids and rust: a new perspective on the chemical revolution. Perspectives on Science, 29(2):215–236, March/ April 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL https://muse.jhu.edu/article/787909.

Jalobeanu:2012:INH

[Jal12] Dana Jalobeanu. Idolatry, natural history, and spiritual medicine: Francis Bacon and the neo-stoic Protestantism of the late Sixteenth Century. Perspectives on Science, 20(2):207–226, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.2.jalobeanu.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.2.jalobeanu.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00063.

Jalobeanu:2016:DEF

[Jal16] Dana Jalobeanu. Disciplining experience: Francis Bacon’s experimental series and the art of experimenting. Perspectives on Science, 24(3):324–342, May/June 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/618108;   http://muse.jhu.edu/article/
618108/pdf.

Janssen:2002:CSE

[Jan02] Michel Janssen. COI stories: Explanation and evidence in the history of science. Perspectives on Science, 10(4):457– 522, Winter 2002. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v010/10.4janssen.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v010/10.
4janssen.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/106361402322288066.

Janiak:2004:KPS

[Jan04] Andrew Janiak. Kant as philosopher of science. Perspectives on Science, 12(3):339–363, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v012/12.3janiak.html;
http://muse.jhu.edu/journals/perspectives_on_science/v012/
12.3janiak.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/1063614042795453.

Jebeile:2018:ESW

[Jeb18] Julie Jebeile. Explaining with simulations: Why visual representations matter. Perspectives on Science, 26(2):213–238, March/April 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00273.

Jesseph:1993:HMM

[Jes93] Douglas M. Jesseph. Hobbes and mathematical method. Perspectives on Science, 1(2):306–341, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Jesseph:1998:LFC

[Jes98] Douglas Michael Jesseph. Leibniz on the foundations of the calculus: The question of the reality of infinitesimal magnitudes. Perspectives on Science, 6(1–2):6–40, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v006/6.1jesseph.html. Special Issue: Leibniz and

the Sciences.

Jesseph:2004:GHB

[Jes04] Douglas Michael Jesseph. Galileo, Hobbes, and the Book of Nature. Perspectives on Science, 12(2):191–211, Summer 2004. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.2jesseph.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v012/12.
2jesseph.pdf;   http://www.mitpressjournals.org/doi/abs/

10.1162/106361404323119871. Special Issue: Galileo in Paris.

Jesseph:2007:DPE

[Jes07] Douglas Michael Jesseph. Descartes, Pascal, and the epistemology of mathematics: The case of the cycloid. Perspectives on Science, 15(4):410–433, Winter 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v015/15.4jesseph.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v015/15.4jesseph.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/posc.2007.15.4.410.

Jesseph:2010:MMD

[Jes10] Douglas Jesseph. Machines, mechanism, and the development of mechanics: Contemporary understandings. Perspectives on Science, 18(1):98–112, Spring 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
1.jesseph.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.1.jesseph.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2010.18.1.98.

Jha:2006:BTP

[Jha06a] Stefania Ruzsits Jha. The bid to transcend Popper, and the Lakatos-Polanyi connection. Perspectives on Science, 14(3):318– 346, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.3jha02.html; http://muse.
jhu.edu/journals/perspectives_on_science/v014/14.3jha02.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2006.14.3.318.

Jha:2006:EIH

[Jha06b] Stefania Ruzsits Jha. Editor’s introduction: Hungarian studies in Lakatos’ philosophies of mathematics and science. Perspectives on Science, 14(3):257–262, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
3jha01.html; http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.3jha01.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.3.257.

Jacobs:2017:AMM

[JM17] Struan Jacobs and Phil Mullins. Anthropological materials in the making of Michael Polanyi’s metascience. Perspectives on Science, 25(2):261–285, March/April 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00243.

Johnson:2005:RTK

[Joh05] Ann Johnson. Revisiting technology as knowledge. Perspectives on Science, 13(4):554–573, Winter 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v013/13.4johnson.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v013/13.4johnson.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361405775466090.

Johnson:2009:MMC

[Joh09] Ann Johnson. Modeling molecules: Computational Nanotechnology as a knowledge community. Perspectives on Science, 17(2):144–173, Summer 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v017/17.
2.johnson.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v017/17.2.johnson.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2009.17.2.144.

Josefowicz:2003:HD

[Jos03] Diane Greco Josefowicz. Histories of discovery. Perspectives on Science, 11(3):346–364, Fall 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v011/11.
3josefowicz.html;   http://muse.jhu.edu/journals/perspectives_

on_science/v011/11.3josefowicz.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361403773062687.

Josefowicz:2005:EPS

[Jos05] Diane Greco Josefowicz. Experience, pedagogy, and the study of terrestrial magnetism. Perspectives on Science, 13(4):452– 494, Winter 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v013/13.4josefowicz.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v013/13.
4josefowicz.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361405775466108.

Joy:1995:RAF

[Joy95] Lynn S. Joy. Rationality among the friends of truth. The Gassendi– Descartes controversy. Perspectives on Science, 3(4):429–449, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi.

Jost:2023:SRM

[JR23] Jürgen Jost and Guillermo Restrepo. Self-reinforcing mechanisms driving the evolution of the chemical space. Perspectives on Science, 31(5):555–593, September-October 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
5/555/2158912/posc_a_00588.pdf.

Jukola:2017:SEI

[Juk17] Saana Jukola. A social epistemological inquiry into biases in journal peer review. Perspectives on Science, 25(1):124–148, January/ February 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/647991;
 http://muse.jhu.edu/article/647991/pdf.

Jaton:2023:PAO

[JV23] Florian Jaton and Dominique Vinck. Politicizing algorithms by other means: Toward inquiries for affective dissensions. Perspectives on Science, 31(1):84–118, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
1/84/2073144/posc_a_00582.pdf.

Jacobson:1998:IOP

[JZ98] David Jacobson and Charles A. Ziegler. Insider and outsider perspectives in the anthropology of science: a cautionary tale. Perspectives on Science, 6(4):361–380, Winter 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v006/6.4jacobson.html.

Kaiser:1996:TMS

[Kai96] Matthias Kaiser. Toward more secrecy in science? Comments on some structural changes in science — and on their implications for an ethics of science. Perspectives on Science, 4(2):207–230, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://search.ebscohost.com/login.
aspx?direct=true&db=aph&AN=6369837&site=ehost-live.

Karaca:2017:REP

[Kar17] Koray Karaca. Representing experimental procedures through diagrams at CERN’s Large Hadron Collider: The communicatory value of diagrammatic representations in collaborative research. Perspectives on Science, 25(2):177–203, March/April 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00240.

Kaiser:2022:MNR

[KBG22] Matthias Kaiser, Tatjana Buklijas, and Peter Gluckman. Models and numbers: Representing the world or imposing order? Perspectives on Science, 30(4):525–548, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
4/525/2038166/posc_a_00373.pdf.

Kitcher:1996:SER

[KC96] Philip Kitcher and Nancy D. Cartwright. Science and ethics: Reclaiming some neglected questions. Perspectives on Science, 4(2): 145–153, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://search.ebscohost.com/
login.aspx?direct=true&db=aph&AN=6369821&site=ehost-live.

Keller:1994:BNM

[Kel94] Evelyn Fox Keller. The body of a new machine: Situating the organism between telegraphs and computers. Perspectives on Science, 2(3):302–323, Fall 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Kevles:2015:IYN

[Kev15] Daniel J. Kevles. Inventions, yes; nature, no: The productsof-nature doctrine from the American Colonies to the U.S. courts. Perspectives on Science, 23(1):13–34, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.1.kevles.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v023/23.1.kevles.pdf.

Kuhn:2012:SSR

[KH12] Thomas S. Kuhn and Ian Hacking. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA and London, UK, fourth edition, 2012. ISBN 0-226-45811-3 (hardcover), 0-226-45812-1 (paperback), 0-226-45814-8 (e-book). xlvi + 217 pp. LCCN Q175 .K95 2012. With an introductory essay by Ian Hacking.

Kedar:2017:RBC

[KH17] Yael Kedar and Giora Hon. Roger Bacon (c. 1220–1292) and his system of laws of nature: Classification, hierarchy and significance. Perspectives on Science, 25(6):719–??, November/December 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00263.

Khalili:2024:RPR

[Kha24] Mahdi Khalili. Reality as persistence and resistance. Perspectives on Science, 32(2):184–206, March-April 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/32/2/184/2345640/posc_
a_00604.pdf.

Kingsland:1993:HSC

[Kin93] Sharon E. Kingsland. A humanistic science: Charles Judson Herrick and the struggle for psychobiology at the University of Chicago. Perspectives on Science, 1(3):445–477, Fall 1993. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Kindi:2005:RHS

[Kin05] Vasso Kindi. The relation of history of science to philosophy of science in The Structure of Scientific Revolutions and Kuhn’s later philosophical work. Perspectives on Science, 13(4):495–530, Winter 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.4kindi.html; http://muse.jhu.edu/journals/
perspectives_on_science/v013/13.4kindi.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361405775466117.

Kindi:2010:NRA

[Kin10] Vasso Kindi. Novelty and revolution in art and science: The connection between Kuhn and Cavell. Perspectives on Science, 18(3):284–310, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v018/18.3.kindi.html;
http://muse.jhu.edu/journals/perspectives_on_science/v018/
18.3.kindi.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00011.

Kiss:2006:HML

[Kis06] Olga Kiss. Heuristic, methodology or logic of discovery? Lakatos on patterns of thinking. Perspectives on Science, 14(3):302–317, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.3kiss.html; http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.3kiss.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2006.14.3.302.

Kitcher:1999:URI

[Kit99] Philip Kitcher. Unification as a regulative ideal. Perspectives on Science, 7(3):337–348, Fall 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v007/7.3kitcher.html;
http://muse.jhu.edu/journals/perspectives_on_science/v007/
7.3kitcher.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.1999.7.3.337.

Katz:2011:CC

[KK11] Karin U. Katz and Mikhail G. Katz. Cauchy’s continuum. Perspectives on Science, 19(4):426–452, Winter 2011. CODEN PRSIEU.

ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v019/19.
4.katz.html; http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.4.katz.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00047.

Knuuttila:2021:BCV

[KL21] Tarja Knuuttila and Andrea Loettgers. Biological control variously materialized: Modeling, experimentation and exploration in multiple media. Perspectives on Science, 29(4):468–492, July/ August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/468/1959260/posc_a_00379.pdf; https:/
/muse.jhu.edu/article/804407.

Klein:2005:EIE

[Kle05a] Ursula Klein. Experiments at the intersection of experimental history, technological inquiry, and conceptually driven analysis: a case study from early Nineteenth-Century France. Perspectives on Science, 13(1):1–48, Spring 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v013/13.
1klein.html; http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.1klein.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/1063614053714063.

Klein:2005:ITP

[Kle05b] Ursula Klein. Introduction: Technoscientific productivity. Perspectives on Science, 13(2):139–141, Summer 2005. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/106361405774270548.

Klein:2005:TAL

[Kle05c] Ursula Klein. Technoscience avant la lettre. Perspectives on Science, 13(2):226–266, Summer 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v013/13.2klein02.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v013/13.2klein02.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361405774270557. Special Issue: Techno-

scientific Productivity, Part 1.

Kuorikoski:2015:BPE

[KM15]  Jaakko Kuorikoski and Caterina Marchionni.  Broadening

the perspective: Epistemic, social, and historical aspects of scientific modelling. Perspectives on Science, 23(4):381– 385, Winter 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v023/23.4.kuorikoski.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.4.kuorikoski.pdf; http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_e_00179.

Kochan:2011:GRR

[Koc11] Jeff Kochan. Getting real with Rouse and Heidegger. Perspectives on Science, 19(1):81–115, Spring 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v019/19.1.kochan.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v019/19.1.kochan.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_a_00026.

Kochan:2021:AAL

[Koc21] Jeff Kochan. Animism, Aristotelianism, and the legacy of William Gilbert’s De Magnete. Perspectives on Science, 29(2):157–188, March/April 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://muse.jhu.edu/article/
787907.

Kochan:2022:IAE

[Koc22] Jeff Kochan. Ingold’s animism and European science. Perspectives on Science, 30(4):783–817, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/4/783/2038155/posc_
a_00425.pdf.

Kokkov:2019:WMR

[Kok19] Kristin Kokkov. Warrants, middle-range theories, and inferential scaffolding in archaeological interpretation. Perspectives on Science, 27(2):171–186, March/April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00304.

Konoval:2014:WDD

[Kon14] Brandon Konoval. What has Dayton to do with Sils-Maria?: Nietzsche and the Scopes Trial. Perspectives on Science, 22(4):545– 573, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.4.konoval.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v022/22.
4.konoval.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00149.

Konoval:2018:PPD

[Kon18] Brandon Konoval. Pythagorean pipe dreams? Vincenzo Galilei, Marin Mersenne, and the pneumatic mysteries of the pipe organ. Perspectives on Science, 26(1):1–51, January/February 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00266.

KormosBarkan:1994:SMC

[Kor94] Diana Kormos Barkan. Simply a matter of chemistry? The Nobel Prize for 1920. Perspectives on Science, 2(4):357–395, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Kosman:2004:MAP

[Kos04] Louis Aryeh Kosman. Mechanisms in ancient philosophy of science. Perspectives on Science, 12(3):244–261, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.3kosman.html; http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.3kosman.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/1063614042795417.

Kostic:2019:MSE

[Kos19a] Daniel Kostić. Minimal structure explanations, scientific understanding and explanatory depth. Perspectives on Science, 27 (1):48–67, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00299.

Kostic:2019:UDM

[Kos19b] Daniel Kostić. Unifying the debates: Mathematical and noncausal explanations. Perspectives on Science, 27(1):1–6, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_e_00296.

Koskinen:2023:SIR

[Kos23] Inkeri Koskinen. Societal impact in research collaborations beyond the boundaries of science. Perspectives on Science, 31(6):744–770, November-December 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/31/6/744/2172269/posc_a_00593.pdf.

Kourany:2012:ISR

[Kou12] Janet A. Kourany. The ideal of socially responsible science: Reply to Dupré, Rolin, Solomon, and Giere. Perspectives on Science, 20(3):344–352, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v020/20.
3.kourany.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.3.kourany.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00070.

Kouw:2016:SSG

[Kou16] Matthijs Kouw. Standing on the shoulders of giants — and then looking the other way? Epistemic opacity, immersion, and modeling in hydraulic engineering. Perspectives on Science, 24(2):206– 227, March/April 2016. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00201.

Kragh:1997:EUG

[Kra97] Helge Kragh. The electrical universe: Grand cosmological theory versus mundane experiments. Perspectives on Science, 5(2):199–231, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
10780091&site=ehost-live.

Kragh:2013:MPI

[Kra13] Helge Kragh. “The Most Philosophically Important of All the Sciences”: Karl Popper and physical cosmology. Perspectives on Science, 21(3):325–357, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v021/21.3.kragh.html;

http://muse.jhu.edu/journals/perspectives_on_science/v021/
21.3.kragh.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00102.

Krohs:2021:EBR

[Kro21] Ulrich Krohs. The epistemology of biomimetics: The role of models and of morphogenetic principles. Perspectives on Science, 29(5):583–601, September/October 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/5/583/1964249/posc_
a_00385.pdf; https://muse.jhu.edu/article/810912.

Katz:2013:AEM

[KSS13] Mikhail G. Katz, David M. Schaps, and Steven Shnider. Almost equal: The method of adequality from Diophantus to Fermat and beyond. Perspectives on Science, 21(3):283–324, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.3.katz.html; http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.3.katz.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/POSC_a_00101.

Kuhn:1962:SSR

[Kuh62] Thomas S. Kuhn. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA and London, UK, 1962. xv + 172 pp. LCCN Q175 .K95.

Kuhn:1970:SSR

[Kuh70] Thomas S. Kuhn. The Structure of Scientific Revolutions. International encyclopedia of unified science. Foundations of the unity of science, v. 2, no. 2. University of Chicago Press, Chicago, IL, USA and London, UK, second enlarged edition, 1970. ISBN 0-226-45803-2 (hardcover), 0-226-45804-0 (paperback), 0-226-45807-5. xii + 210 pp. LCCN Q175 .K95 1970.

Kuhn:1996:SSR

[Kuh96] Thomas S. Kuhn. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA and London, UK, third edition, 1996. ISBN 0-226-45807-5 (hardcover), 0-226-45808-3 (paperback). xiv + 212 pp. LCCN Q175 .K95 1996. URL http:/

/www.loc.gov/catdir/description/uchi051/96013195.html;
 http://www.loc.gov/catdir/toc/uchi051/96013195.html.

Kuhlmann:2021:EFA

[Kuh21] Meinard Kuhlmann. On the exploratory function of agentbased modeling. Perspectives on Science, 29(4):510–536, July/ August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/510/1959236/posc_a_00381.pdf; https:/
/muse.jhu.edu/article/804409.

Kukla:2008:NO

[Kuk08] Rebecca Kukla. Naturalizing objectivity. Perspectives on Science, 16(3):285–302, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v016/16.3.kukla.html;
http://muse.jhu.edu/journals/perspectives_on_science/v016/
16.3.kukla.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2008.16.3.285.

Kusukawa:2019:ERS

[Kus19a] Sachiko Kusukawa. The early Royal Society and visual culture. Perspectives on Science, 27(3):350–394, May/June 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00311.

Kusukawa:2019:IMV

[Kus19b] Sachiko Kusukawa. Introduction to making visible: The visual and graphic practices of the early Royal Society. Perspectives on Science, 27(3):345–349, May/June 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_e_00310.

Kyllingstad:2022:NAH

[Kyl22] Jon Røyne Kyllingstad. The Norwegian Association for Heredity Research and the organized international eugenics movement. Expertise, authority, transnational networks and international organization in Norwegian genetics and eugenics (1919– 1934). Perspectives on Science, 30(1):77–107, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
1/77/1987184/posc_a_00403.pdf.

LaBerge:1999:HSH

[La 99] Ann Elizabeth Fowler La Berge. The history of science and the history of microscopy. Perspectives on Science, 7(1): 111–142, Spring 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v007/7.1laberge.html;
http://www.mitpressjournals.org/doi/abs/10.1162/posc.1999.
7.1.111.

LaBerge:2004:DSP

[La 04] Ann Elizabeth Fowler La Berge. Debate as scientific practice in Nineteenth-Century Paris: The controversy over the microscope. Perspectives on Science, 12(4):424–453, Winter 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.4la_berge.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v012/12.4la_berge.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/1063614042775996.

Lampert:2003:PTA

[Lam03] Timm Lampert. Psychophysical and tractarian analysis. Perspectives on Science, 11(3):285–317, Fall 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v011/11.3lampert.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v011/11.3lampert.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361403773062650.

Lancaster:2012:NHR

[Lan12] James A. T. Lancaster. Natural histories of religion: a (Baconian) ‘Science’? Perspectives on Science, 20(2):246–267, Summer 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.2.lancaster.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v020/20.2.lancaster.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00065.

Lange:2014:DER

[Lan14] Marc Lange. Did Einstein really believe that principle theories are explanatorily powerless? Perspectives on Science, 22 (4):449–463, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v022/22.4.lange.html;
http://muse.jhu.edu/journals/perspectives_on_science/v022/
22.4.lange.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00145.

Lange:2019:TBC

[Lan19] Marc Lange. Are there both causal and non-causal explanations of a rocket’s acceleration? Perspectives on Science, 27(1):7–25, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00297.

Laubichler:2000:ODL

[Lau00] Manfred D. Laubichler. The organism is dead. Long live the organism! Perspectives on Science, 8(3):286–315, Fall 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.3laubichler.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v008/8.3laubichler.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
106361400750340505.

LeGrand:1993:CHR

[Le 93] Homer Le Grand. Choke-holds, radiolarian cherts, and Davy Jones’s locker. Perspectives on Science, 1(1):24–65, Spring 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Lefevre:2005:SL

[Lef05] Wolfgang Lefèvre. Science as labor. Perspectives on Science, 13(2): 194–225, Summer 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v013/13.2lefevre.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v013/13.
2lefevre.pdf;   http://www.mitpressjournals.org/doi/abs/

10.1162/106361405774270539. Special Issue: Technoscientific Productivity, Part 1.

Lefkaditou:2022:BAR

[Lef22] Ageliki Lefkaditou. Blood affairs: Racial blood group research and nation building in Greece, 1920s–1940s. Perspectives on Science, 30 (1):48–76, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/30/1/48/1987174/posc_a_00402.pdf.

Lehoux:2003:TFE

[Leh03] Daryn Lehoux. Tropes, facts, and empiricism. Perspectives on Science, 11(3):326–345, Fall 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v011/11.
3lehoux.pdf; http://www.mitpressjournals.org/doi/abs/10.
1162/106361403773062678.

Lennon:1995:DGR

[Len95] Thomas M. Lennon. Descartes and Gassendi: a reply to Glouberman. Perspectives on Science, 3(4):520–533, Winter 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special Issue: Descartes versus Gassendi. See [Glo95].

Lennox:2010:DGC

[Len10] James G. Lennox. The Darwin/Gray correspondence 1857–1869: an intelligent discussion about chance and design. Perspectives on Science, 18(4):456–479, Winter 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v018/18.4.lennox.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v018/18.4.lennox.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_a_00018.

Lenhard:2015:SRC

[Len15] Johannes Lenhard. Simulation, representation, and cartography: Compiling a virtual atlas. Perspectives on Science, 23(4):386– 404, Winter 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.4.lenhard.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v023/23.
4.lenhard.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00180.

Leonelli:2014:DID

[Leo14] Sabina Leonelli. Data interpretation in the digital age. Perspectives on Science, 22(3):397–417, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
3.leonelli.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.3.leonelli.pdf.

Levere:1996:RER

[Lev96] Trevor H. Levere. Review essay: Romanticism, natural philosophy, and the sciences: a review and bibliographic essay. Perspectives on Science, 4(4):463–488, Winter 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6370249&site=ehost-live.

Levine:2010:TKC

[Lev10] Alex Levine. Thomas Kuhn’s cottage. Perspectives on Science, 18 (3):369–377, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.3.levine.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v018/18.
3.levine.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_r_00012.

Lewis:2001:LMC

[Lew01] Eric Lewis. The legacy of Margaret Cavendish. Perspectives on Science, 9(3):341–365, Fall 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v009/9.
3lewis.html; http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.3lewis.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/10636140160176189.

Lewis:2006:RH

[Lew06]  Rhodri Lewis.  Robert Hooke at 371.  Perspectives on

Science, 14(4):558–573, Winter 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
4lewis.html; http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.4lewis.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.4.558.

Lewis:2007:CR

[Lew07]  Eric P. Lewis.  Cartesianism revisited.  Perspectives on

Science, 15(4):493–522, Winter 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v015/15.
4lewis.html; http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.4lewis.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2007.15.4.493.

LiVigni:2021:REC

[Li 21] Fabrizio Li Vigni. Regimes of evidence in complexity sciences. Perspectives on Science, 29(1):62–103, January/February 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00361.

LiVigni:2023:PCS

[Li 23] Fabrizio Li Vigni. The promises of complexity sciences: a critique. Perspectives on Science, 31(4):465–502, ???? 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

https://muse.jhu.edu/pub/6/article/902863.

Lindemann:2019:SOS

[Lin19] Anna Lindemann. Scientific objectivity and subjectivity in Eighteenth Century pharmacology. Perspectives on Science, 27(6): 787–809, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00325.

Liu:2016:ESE

[Liu16] Jennifer A. Liu. Emerging science, emerging democracy: Stem cell research and policy in Taiwan. Perspectives on Science, 24 (5):609–636, September/October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/628783;   http://muse.jhu.edu/article/
628783/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00225.

Lenoir:1995:IMD

[LL95] Timothy Lenoir and Christophe Lecuyer. Instrument makers and discipline builders: The case of nuclear magnetic resonance. Perspectives on Science, 3(3):276–345, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Llored:2021:CMS

[Llo21] Jean-Pierre Llored. Chemistry and measurement: Some philosophical lessons. Perspectives on Science, 29(6):782–801, November/December 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/6/782/1972158/posc_a_00394.pdf; https:/
/muse.jhu.edu/article/839325.

Longino:2008:NNC

[Lon08] Helen E. Longino. Norms and naturalism: Comments on Miriam Solomon’s social empiricism. Perspectives on Science, 16(3):241– 245, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.3.longino.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v016/16.
3.longino.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2008.16.3.241.

Look:2002:MTL

[Loo02] Brandon Look. Marks and traces: Leibnizian scholarship past, present, and future. Perspectives on Science, 10(1):123–146, Spring 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.1look.html; http://muse.jhu.edu/journals/
perspectives_on_science/v010/10.1look.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361402762674816.

Lopes:2009:DSS

[Lop09] Dominic McIver Lopes. Drawing in a social science: Lithic illustration. Perspectives on Science, 17(1):5–25, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v017/17.1.lopes.html; http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.1.lopes.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/posc.2009.17.
1.5.

Lopez:2017:FSP

[Lóp17] Raúl Necochea López. Fertility surveyors and population-making technologies in Latin America. Perspectives on Science, 25(5): 631–654, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00257.

Lowood:1994:LTT

[LR94] Henry F. Lowood and Robin E. Rider. Literary technology and typographic culture: The instrument of print in early modern science. Perspectives on Science, 2(1):1–37, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Luggin:2022:TNO

[Lug22] Johanna Luggin. ”Trust no one but yourself”: William Gilbert’s use of experiment and rejection of authority, reconsidered. Perspectives on Science, 30(6):925–949, November/December 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
6/925/2059532/posc_a_00564.pdf.

Lundgren:2018:NRE

[Lun18] Jakob Lundgren. No “real” experts: Unexpected agreement over disagreement in STS and philosophy of science. Perspectives on Science, 26(6):722–735, November/December 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00295.

Luomanen:2013:SSM

[Luo13] Petri Luomanen. Social-scientific modeling in biblical and related studies. Perspectives on Science, 21(2):202–220, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.luomanen.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.2.luomanen.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

POSC_a_00094. Special Issue on Modeling in the Social Sciences.

Lupacchini:2014:IHA

[Lup14] Rossella Lupacchini. Introduction: Hilbert’s axiomatics as ‘symbolic form’? Perspectives on Science, 22(1):1–34, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.1.lupacchini.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v022/22.1.lupacchini.
pdf.

Luthy:2000:WDS

[Lüt00] Christoph Lüthy. What to do with Seventeenth-Century natural philosophy? A taxonomic problem. Perspectives on Science, 8(2):164–195, Summer 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v008/8.
2luthy.html; http://muse.jhu.edu/journals/perspectives_

on_science/v008/8.2luthy.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400568064.

Lohse:2020:IPS

[LWR20] Simon Lohse, Martin S. Wasmer, and Thomas A. C. Reydon. Integrating philosophy of science into research on ethical, legal and social issues in the life sciences. Perspectives on Science, 28(6): 700–736, November/December 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00357.

Lynch:2019:BKS

[Lyn19] William T. Lynch. Between kin selection and cultural relativism: Cultural evolution and the origin of inequality. Perspectives on Science, 27(2):278–315, March/April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00308.

Lyngs:2024:SCE

[Lyn24] Allan Lyngs. Science as a collective effort: Collaboration at the Zoophysiological Laboratory 1911–1945. Perspectives on Science, 32(2):141–183, March-April 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/32/2/141/2345642/posc_a_00620.pdf.

Maglo:2003:RNG

[Mag03] Koffi Maglo. The reception of Newton’s gravitational theory by Huygens, Varignon, and Maupertuis: How normal science may be revolutionary. Perspectives on Science, 11(2):135–169, Summer 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.2maglo.html; http://muse.jhu.edu/journals/
perspectives_on_science/v011/11.2maglo.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361403322495876.

Maglo:2011:CAB

[Mag11] Koffi N. Maglo. The case against biological realism about race: From Darwin to the post-genomic era. Perspectives on Science, 19(4):361–390, Winter 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v019/19.4.maglo.html;
http://muse.jhu.edu/journals/perspectives_on_science/v019/

19.4.maglo.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00048.

Majer:2014:AMC

[Maj14] Ulrich Majer. The “Axiomatic Method” and its constitutive role in physics. Perspectives on Science, 22(1):56–79, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.1.majer.html;   http:/
/muse.jhu.edu/journals/perspectives_on_science/v022/22.
1.majer.pdf.

Maki:1999:SFM

[Mäk99] Uskali Mäki. Science as a free market: a reflexivity test in an economics of economics. Perspectives on Science, 7 (4):486–509, Winter 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/journals/perspectives_on_science/v007/7.4maki.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v007/7.4maki.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.1999.7.4.486.

Malet:2022:MDH

[Mal22] Antoni Malet. Milliet Dechales as historian of mathematics. Perspectives on Science, 30(3):463–492, May/June 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
3/463/2023843/posc_a_00423.pdf.

Mansnerus:2013:MSS

[Man13] Erika Mansnerus. Modeling in the social sciences: Interdisciplinary comparison. Perspectives on Science, 21(2):267–272, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.mansnerus.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.2.mansnerus.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

POSC_a_00099. Special Issue on Modeling in the Social Sciences.

Manzo:2019:HAE

[Man19] Silvia Manzo. Historiographical approaches on experience and empiricism in the Early Nineteenth-Century: Degérando and Tennemann. Perspectives on Science, 27(5):655–679, September/October 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00320.

Marcum:1996:ETC

[Mar96] James A. Marcum. Experimentation and theory choice: Is thrombin an enzyme? Perspectives on Science, 4(4):434–462, Winter 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6370218&site=ehost-live.

Marcum:2009:NLC

[Mar09] James A. Marcum. The nature of light and color: Goethe’s “Der Versuch als Vermittler” versus Newton’s Experimentum Crucis. Perspectives on Science, 17(4):457–481, Winter 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v017/17.4.marcum.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v017/17.4.marcum.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2009.17.4.457.

Marchionni:2013:MBE

[Mar13] Caterina Marchionni. Model-based explanation in the social sciences: Modeling kinship terminologies and romantic networks. Perspectives on Science, 21(2):175–180, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.marchionni.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.2.marchionni.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/POSC_

a_00091. Special Issue on Modeling in the Social Sciences.

Martin:2016:AEM

[Mar16] Craig Martin. The aeolipile as experimental model in early modern natural philosophy. Perspectives on Science, 24(3):264–284, May/June 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/618105;
 http://muse.jhu.edu/article/618105/pdf.

Martin:2020:WPG

[Mar20] Marta Velasco Martín. Women and partnership genealogies in Drosophila population genetics. Perspectives on Science, 28(2):277–317, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00341.

Martin:2022:GCM

[Mar22] Craig Martin. Girolamo Cardano’s meteorological predictions: Hippocratism, weather signs, winds, and the limits of astrology. Perspectives on Science, 30(5):851–873, September/October 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
5/851/2053599/posc_a_00562.pdf.

Massimi:2010:GMN

[Mas10] Michela Massimi. Galileo’s mathematization of nature at the crossroad between the empiricist and the Kantian tradition. Perspectives on Science, 18(2):152–188, Summer 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
2.massimi.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.2.massimi.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2010.18.2.152.

Mattila:2005:IMM

[Mat05] Erika Mattila. Interdisciplinarity “In the Making”: Modeling infectious diseases. Perspectives on Science, 13(4):531– 553, Winter 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v013/13.4mattila.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v013/13.
4mattila.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/106361405775466081.

Mate:2006:ASI

[Mát06] András Máté. Árpád Szabó and Imre Lakatos, or the relation between history and philosophy of mathematics. Perspectives on Science, 14(3):282–301, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
3mate.html; http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.3mate.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.3.282.

Matei:2020:EMW

[Mat20] Oana Matei. Experimenting with matter in the works of Gabriel Plattes. Perspectives on Science, 28(3):398–420, May/June 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00345.

Montuschi:2024:UWP

[MB24] Eleonora Montuschi and Baptiste Bedessem. Understanding what in public understanding of science. Perspectives on Science, 32(2): 207–229, March-April 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/32/2/207/2345638/posc_a_00603.pdf.

Malet:2010:MMM

[MC10] Antoni Malet and Daniele Cozzoli. Mersenne and mixed mathematics. Perspectives on Science, 18(1):1–8, Spring 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.1.malet.html; http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.1.malet.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/posc.2010.18.
1.1.

Moroni:2022:UPC

[MC22] Stefano Moroni and Daniele Chiffi. Uncertainty and planning: Cities, technologies and public decision-making. Perspectives on Science, 30(2):237–259, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/2/237/2005006/posc_
a_00413.pdf.

McAllister:2016:RES

[McA16] James W. McAllister. Rhetoric of effortlessness in science. Perspectives on Science, 24(2):145–166, March/April 2016. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00198.

McDonald:2003:DSP

[McD03] Patrick Joseph McDonald. Demonstration by simulation: The philosophical significance of experiment in Helmholtz’s theory of perception. Perspectives on Science, 11(2):170–207, Summer 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.2mcdonald.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v011/11.2mcdonald.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361403322495885.

McDonough:2010:LOC

[McD10] Jeffrey K. McDonough. Leibniz’s optics and contingency in nature. Perspectives on Science, 18(4):432–455, Winter 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.4.mcdonough.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v018/18.4.mcdonough.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00017.

McIntyre:2001:APC

[McI01] Lee C. McIntyre. Accommodation, prediction, and confirmation. Perspectives on Science, 9(3):308–323, Fall 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.3mcintyre.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v009/9.3mcintyre.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/10636140160176161.

Menn:2000:DCP

[Men00] Stephen Philip Menn. On Dennis Des Chene’s Physiologia. Perspectives on Science, 8(2):119–143, Summer 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.2menn.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400568046.

Merino:2023:LPP

[Mer23] Noemí Sanz Merino. Latour on politics: Political turn in epistemology or ontological turn in politics? Perspectives on Science, 31(1): 119–138, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/31/1/119/2073135/posc_a_00583.pdf.

Methuen:1997:CNS

[Met97] Charlotte Methuen. ‘the comet or new star’: Theology and the interpretation of the Nova 1572. Perspectives on Science, 5(4):499– 515, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://search.ebscohost.com/login.
aspx?direct=true&db=aph&AN=6370933&site=ehost-live.

Methot:2015:SSP

[Mét15] Pierre-Olivier Méthot. Science and science policy: Regulating “select agents” in the age of synthetic biology. Perspectives on Science, 23(3):280–309, Fall 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v023/23.3.methot.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.3.methot.pdf.

Meynell:2018:PFD

[Mey18] Letitia Meynell. Picturing Feynman diagrams and the epistemology of understanding. Perspectives on Science, 26(4):459–481, July/August 2018. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00283.

Melogno:2023:TGT

[MG23] Pablo Melogno and Leandro Giri. Towards a genealogy of Thomas Kuhn’s semantics. Perspectives on Science, 31(4):385–404, ???? 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://muse.jhu.edu/pub/6/article/902860.

Morgan:2013:MPS

[MGY13] Mary S. Morgan and Till Grüne-Yanoff. Modeling practices in the social and human sciences. An interdisciplinary exchange. Perspectives on Science, 21(2):143–156, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.morgan01.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.2.morgan01.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

POSC_a_00089. Special Issue on Modeling in the Social Sciences.

Mahoney:2011:HC

[MH11] Michael S. (Michael Sean) Mahoney and Thomas Haigh, editors. Histories of computing. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-05568-3. vi + 250 pp. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh. See the Mahoney obituary [CK13] for comments on the author and this book.

Mikkelson:1997:MMC

[Mik97] Gregory M. Mikkelson. Methods and metaphors in community ecology: The problem of defining stability. Perspectives on Science, 5(4):481–498, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6370927&site=ehost-live.

Miller:2018:REC

[Mil18] David Marshall Miller. Regressus and empiricism in the controversy about Galileo’s lunar observations. Perspectives on Science, 26(3):293–324, May/June 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00276.

Mirza:2022:CHK

[Mir22] Ali Mirza. Of chimeras, harmony, and kintsugi: Towards a historicist epistemology of paleontological reconstruction, theory-change, and exploring heuristics. Perspectives on Science, 30(4):657–695, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/30/4/657/2038179/posc_a_00399.pdf.

Mizrahi:2020:CSM

[Miz20] Moti Mizrahi. The case study method in philosophy of science: An empirical study. Perspectives on Science, 28(1):63–88, January/ February 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://www.mitpressjournals.org/
doi/abs/10.1162/posc_a_00333;   https://muse.jhu.edu/article/
749743.

Malaterre:2023:MS

[MJLG23] Christophe Malaterre, Emmanuelle J. Javaux, and Purificación López-García. Misconceptions in science. Perspectives on Science, 31(6):717–743, November-December 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/31/6/717/2172274/posc_
a_00590.pdf.

Mladenovic:2007:MHR

[Mla07] Bojana Mladenović. ”Muckraking in History”: The role of the history of science in Kuhn’s philosophy. Perspectives on Science, 15(3):261–294, Fall 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v015/15.
3mladenovic.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.3mladenovic.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2007.15.3.261.

McCabe:2015:GPP

[MM15] Linda L. McCabe and Edward R. B. McCabe. Gene patents: Perspective from the clinic and the laboratory. Perspectives on Science, 23(1):66–79, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v023/23.1.mccabe.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.1.mccabe.pdf.

Mitman:1993:CBB

[MMC93] Gregg Mitman, Jane Maienschein, and Adele E. Clarke. Crossing the borderlands: Biology at Chicago. Perspectives on Science, 1 (3):359–366, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special issue on biology at the University of Chicago in the 20th century.

MacLeod:2019:IIP

[MMMN19] Miles MacLeod, Martina Merz, Uskali Mäki, and Michiru Nagatsu. Investigating interdisciplinary practice: Methodological challenges (introduction). Perspectives on Science, 27(4):545–552, July/August 2019. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_e_00315.

Mody:2009:I

[Mod09] Cyrus C. M. Mody. Introduction. Perspectives on Science, 17(2): 111–122, Summer 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.2.mody.html; http://muse.
jhu.edu/journals/perspectives_on_science/v017/17.2.mody.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2009.17.2.111.

Molina:2017:AEM

[Mol17] Santiago José Molina. Amerindians, Europeans, Makiritare, Mestizos, Puerto Rican, and Quechua: Categorical heterogeneity in Latin American human biology. Perspectives on Science, 25(5): 655–679, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00258.

Montuschi:2017:TNN

[Mon17] Eleonora Montuschi. There is “noise,” and noise. Perspectives on Science, 25(2):204–225, March/April 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00241.

Morgan:2013:ELT

[Mor13] Mary S. Morgan. Experiencing life through modeling. Perspectives on Science, 21(2):245–249, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v021/21.2.morgan.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v021/21.2.morgan.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/POSC_a_00097. Special Issue on Modeling in

the Social Sciences.

Morawski:2021:MA

[Mor21] Roman Z. Morawski. Measurement as abduction. Perspectives on Science, 29(6):742–756, November/December 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/29/6/742/
1972167/posc_a_00392.pdf; https://muse.jhu.edu/article/
839323.

Moynahan:2003:HCP

[Moy03] Gregory B. Moynahan. Hermann Cohen’s Das Prinzip der Infinitesimalmethode, Ernst Cassirer, and the politics of science in Wilhelmine Germany. Perspectives on Science, 11(1):35–75, Spring 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.1moynahan.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v011/11.1moynahan.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361403322286706.

Special issue: History of science and philosophy of science (Montreal, QC, 2002).

Marletto:2014:QPT

[MR14] Chiara Marletto and Mario Rasetti. Quantum physics, topology, formal languages, computation: a categorical view as homage to David Hilbert. Perspectives on Science, 22(1):98–114, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.1.marletto.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v022/22.1.marletto.
pdf.

McCoy:1994:CCC

[MSF94] E. D. McCoy and K. S. Shrader-Frechette. The community concept in community ecology. Perspectives on Science, 2(4):455–475, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Nanay:2011:PDA

[Nan11] Bence Nanay. Popper’s Darwinian analogy. Perspectives on Science, 19(3):337–354, Fall 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v019/19.3.nanay.html;
http://muse.jhu.edu/journals/perspectives_on_science/v019/
19.3.nanay.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00043.

Nash:2018:DTS

[Nas18] Erin J. Nash. In defense of “targeting”: Some dissent about science. Perspectives on Science, 26(3):325–359, May/June 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00277.

Needham:1996:SDE

[Nee96] Paul Needham. Substitution: Duhem’s explication of a chemical paradigm. Perspectives on Science, 4(4):408–433, Winter 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6370214&site=ehost-live.

Nejeschleba:2019:LVP

[Nej19] Tomás Nejeschleba. Light and void. the philosophical background of Valerian Magni’s vacuum experiments. Perspectives on Science, 27(6):767–786, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00324.

Nelson:2023:ALM

[Nel23] Lee Claiborne Nelson. Attending to Latour’s militaristic rhetoric and politics “With other means”. Perspectives on Science, 31(1): 57–83, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/31/1/57/2073131/posc_a_00581.pdf.

Nersessian:2019:IAC

[Ner19] Nancy J. Nersessian. Interdisciplinarities in action: Cognitive ethnography of bioengineering sciences research laboratories. Perspectives on Science, 27(4):553–581, July/August 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00316.

Newman:2009:AAA

[New09a] William R. Newman. Alchemical atoms or artisanal “Building Blocks”?: a response to Klein. Perspectives on Science, 17(2):212– 231, Summer 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.2.newman.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v017/17.
2.newman.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2009.17.2.212.

Newman:2009:BVA

[New09b] William R. Newman. Brian Vickers on alchemy and the occult: a response. Perspectives on Science, 17(4):482–506, Winter 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.4.newman.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v017/17.
4.newman.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2009.17.4.482.

Niss:2016:BML

[Nis16] Martin Niss. Brownian motion as a limit to physical measuring processes: a chapter in the history of noise from the physicists’ point of view. Perspectives on Science, 24(1):29–44, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v024/24.1.niss.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v024/24.1.niss.pdf.

Niss:2018:MDP

[Nis18] Martin Niss. A mathematician doing physics: Mark Kac’s work on the modeling of phase transitions. Perspectives on Science, 26(2):185–212, March/April 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00272.

Nickel:2022:MUT

[NKvdP22] Philip J. Nickel, Olya Kudina, and Ibo van de Poel. Moral uncertainty in technomoral change: Bridging the explanatory gap. Perspectives on Science, 30(2):260–283, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
2/260/2005012/posc_a_00414.pdf.

Nordmann:1994:EAA

[Nor94] Alfred Nordmann. The evolutionary analysis: Apparent error, certified belief, and the defects of asymmetry. Perspectives on Science, 2(2):131–175, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Nordmann:1999:ECI

[Nor99] Alfred Nordmann. Establishing commensurability: Intercalation, global meaning and the unity of science. Perspectives on Science, 7(2):181–195, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v007/7.2nordmann.
html; http://www.mitpressjournals.org/doi/abs/10.1162/
posc.1999.7.2.181.

Nordmann:2002:ANW

[Nor02] Alfred Nordmann. Another new Wittgenstein: The scientific and engineering background of the Tractatus. Perspectives on Science, 10(3):356–384, Fall 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v010/10.
3nordmann.html;   http://muse.jhu.edu/journals/perspectives_

on_science/v010/10.3nordmann.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361402321899087.

Nordmann:2009:ION

[Nor09] Alfred Nordmann. Invisible origins of nanotechnology: Herbert gleiter, materials science, and questions of prestige. Perspectives on Science, 17(2):123–143, Summer 2009. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v017/17.2.nordmann.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v017/17.2.nordmann.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2009.17.2.123.

Nordmann:2021:BBS

[Nor21] Alfred Nordmann. BioTechnology as BioParody — strategies for salience. Perspectives on Science, 29(5):568–582, September/October 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/5/568/1964245/posc_a_00384.pdf; https:/
/muse.jhu.edu/article/810911.

Norton:2023:CCT

[Nor23] John D. Norton. Chance combinatorics: The theory that history forgot. Perspectives on Science, 31(6):771–810, November-December 2023. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/31/6/771/2172271/posc_a_00621.pdf.

Novoa:2020:SSG

[Nov20] Adriana Novoa. Science, sensibility and gender in Argentina, 1820–1852. Perspectives on Science, 28(2):318–340, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00342.

Nukaga:2019:DHR

[Nuk19] Yoshio Nukaga. Development of the hybrid rule and the concept of justice: The selection of subjects in biomedical research. Perspectives on Science, 27(6):891–924, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00329.

Nurmi:2013:MDP

[Nur13] Jari-Erik Nurmi. Modeling developmental processes in psychology. Perspectives on Science, 21(2):181–195, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.nurmi.html; http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.2.nurmi.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00092.

Special Issue on Modeling in the Social Sciences.

Ongaro:2022:NEU

[OA22] Malvina Ongaro and Mattia Andreoletti. Non-empirical uncertainties in evidence-based decision making. Perspectives on Science, 30(2):305–320, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/30/2/305/2005014/posc_a_00416.pdf.

Oberdan:2014:RPM

[Obe14] Thomas Oberdan. Russell’s principles of mathematics and the revolution in Marburg neo-Kantianism. Perspectives on Science, 22(4):523–544, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
4.oberdan.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.4.oberdan.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00148.

Omodeo:2017:PCS

[Omo17] Pietro Daniel Omodeo. Post-Copernican science in Galileo’s Italy. Perspectives on Science, 25(4):393–410, July/August 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00247.

Omodeo:2022:RIL

[Omo22] Pietro Daniel Omodeo. Resources of intellectual legitimacy in Italian cosmological affairs: Cremonini and Bellarmine’s authority conflict (c. 1616). Perspectives on Science, 30(5):874–902, September/October 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/5/874/2053608/posc_a_00563.pdf.

Osbeck:2017:EII

[ON17] Lisa M. Osbeck and Nancy J. Nersessian. Epistemic identities in interdisciplinary science. Perspectives on Science, 25(2):226–260, March/April 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00242.

Oreskes:2008:DHD

[Ore08] Naomi Oreskes. The Devil is in the (historical) details: Continental drift as a case of normatively appropriate consensus? Perspectives on Science, 16(3):253–264, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v016/16.
3.oreskes.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.3.oreskes.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2008.16.3.253.

Orel:2010:SSM

[Ore10] Vítezslav Orel. Science studies and Mendel’s paradigm. Perspectives on Science, 18(2):226–241, Summer 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
2.orel.html; http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.2.orel.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2010.18.2.226.

OMalley:2018:MSM

[OS18] Maureen A. O’Malley and Derek J. Skillings. Methodological strategies in microbiome research and their explanatory implications. Perspectives on Science, 26(2):239–265, March/April 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00274.

Osler:2011:SHG

[Osl11] Margaret J. Osler. The search for the historical Gassendi. Perspectives on Science, 19(2):212–229, Summer 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v019/19.
2.osler.html; http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.2.osler.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00037.

Palmer:1993:LIH

[Pal93] Eric Palmer. Lakatos’s “internal history” as historiography. Perspectives on Science, 1(4):603–626, Winter 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Palmerino:2004:GRG

[Pal04] Carla Rita Palmerino. Gassendi’s reintrepretation of the Galilean theory of tides. Perspectives on Science, 12(2):212–237, Summer 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.2palmerino.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v012/12.2palmerino.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

106361404323119880. Special Issue: Galileo in Paris.

Palmerino:2010:EMP

[Pal10] Carla Rita Palmerino. Experiments, mathematics, physical causes: How Mersenne came to doubt the validity of Galileo’s Law of Free Fall. Perspectives on Science, 18(1):50–76, Spring 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.1.palmerino.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v018/18.1.palmerino.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2010.18.1.50.

Palombi:2015:PBM

[Pal15] Luigi Palombi. The patenting of biological materials in the United States: a state of policy confusion. Perspectives on Science, 23(1):35–65, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v023/23.
1.palombi.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.1.palombi.pdf.

Pannese:2015:ASN

[Pan15] Alessia Pannese. “Anything that is strang”: Normality, deviance, and the tradescants’ collecting legacy. Perspectives on Science, 23(3):335–360, Fall 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v023/23.
3.pannese.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.3.pannese.pdf.

Parker:2014:SUS

[Par14] Wendy S. Parker. Simulation and understanding in the study of weather and climate. Perspectives on Science, 22(3):336– 356, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.3.parker.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v022/22.
3.parker.pdf.

Pashby:2012:DPP

[Pas12] Thomas Pashby. Dirac’s prediction of the positron: a case study for the current realism debate. Perspectives on Science, 20(4): 440–475, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.4.pashby.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v020/20.
4.pashby.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00081.

Pietarinen:2023:RNC

[PB23] Ahti-Veikko Pietarinen and Majid D. Beni. Reinvigorating the Nineteenth Century scientific method: a Peirce-pective on science. Perspectives on Science, 31(5):684–715, September-October 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
5/684/2158905/posc_a_00605.pdf.

Porter:2019:MIR

[PCC^{+}19] Alan L. Porter, Stephen F. Carley, Caitlin Cassidy, Jan Youtie, David J. Schoeneck, Seokbeom Kwon, and Gregg E. A. Solomon. Measuring interdisciplinary research categories and knowledge transfer: a case study of connections between cognitive science and education. Perspectives on Science, 27(4):582–618, July/August 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00317.

Pearce:2017:WLM

[Pea17] Jacob Pearce. ”Why these laws?” — multiverse discourse as a scene of response. Perspectives on Science, 25(3):324–354, May/ June 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00245.

Pecharman:2018:RAF

[Péc18] Martine Pécharman. Roger Ariew and “the first Cartesians”. Perspectives on Science, 26(5):548–562, September/October 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00288.

Pedersen:2008:LEW

[Ped08] Kurt Møller Pedersen. Leonhard Euler’s wave theory of light. Perspectives on Science, 16(4):392–416, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.4.pedersen.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v016/16.4.pedersen.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2008.16.4.392.

Petrescu:2013:DHL

[Pet13] Lucian Petrescu. Descartes on the heartbeat: The Leuven affair. Perspectives on Science, 21(4):397–428, Winter 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.4.petrescu.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.4.petrescu.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00110.

Petrescu:2018:SLC

[Pet18] Lucian Petrescu. Scholastic logic and Cartesian logic. Perspectives on Science, 26(5):533–547, September/October 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00287.

Poznic:2021:STS

[PH21] Michael Poznic and Rafaela Hillerbrand. Scenarios as tools of the scientific imagination: The case of climate projections. Perspectives on Science, 29(1):36–61, January/February 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00360.

Piazza:2024:PSP

[Pia24] Marco Piazza. Proust on the subconscious: Psychic splitting, halfsleep, and Metempsychosis. Perspectives on Science, 32(1):127– 139, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/32/1/127/2210160/posc_a_00602.pdf.

Pickering:1995:CHW

[Pic95] Andy Pickering. Cyborg history and the World War II regime. Perspectives on Science, 3(1):1–48, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). See errata [Ano95c] with replacement for Figure 4, pages 28–29.

Pickering:2001:RS

[Pic01] Andy Pickering. Reading the Structure. Perspectives on Science, 9(4):499–510, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v009/9.4pickering.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v009/9.4pickering.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361401760375848.

Pickering:2005:DSS

[Pic05a] Andrew Pickering. Decentering sociology: Synthetic dyes and social theory. Perspectives on Science, 13(3):352–405, Fall 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.3pickering01.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v013/13.3pickering01.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/106361405774287955.

Special Issue: Technoscientific Productivity, Part 2.

Pickering:2005:DIR

[Pic05b] Andrew Pickering. From dyes to Iraq: a reply to Jonathan Harwood. Perspectives on Science, 13(3):416–425, Fall 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.3pickering02.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v013/13.3pickering02.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/106361405774288008.

Special Issue: Technoscientific Productivity, Part 2.

Pickstone:2005:KAL

[Pic05c] John V. Pickstone. On knowing, acting, and the location of technoscience: a response to Barry Barnes. Perspectives on Science, 13(2):267–278, Summer 2005. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.2pickstone.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v013/13.2pickstone.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

106361405774270566. Special Issue: Technoscientific Productivity, Part 1.

Pierson:1993:CCL

[Pie93] Stuart Pierson. Corpore cadente …: historians discuss Newton’s Second Law. Perspectives on Science, 1(4):627–658, Winter 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Pierson:1994:TMT

[Pie94] Stuart Pierson. Two mathematics, two Gods: Newton and the Second Law. Perspectives on Science, 2(2):231–253, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Pietarinen:2006:IPC

[Pie06] Ahti-Veikko Pietarinen. Interdisciplinarity and Peirce’s classification of the sciences: a Centennial reassessment. Perspectives on Science, 14(2):127–152, Summer 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
2pietarinen.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.2pietarinen.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.2.127.

Pihlstrom:2008:HWH

[Pih08] Sami Pihlström. How (not) to write the history of pragmatist philosophy of science? Perspectives on Science, 16(1):26–69, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.1pihlstrom.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v016/16.1pihlstrom.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2008.16.1.26.

Pinto:2015:CLW

[Pin15] Manuela Fernández Pinto. Commercialization and the limits of well-ordered science. Perspectives on Science, 23(2): 173–191, Summer 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v023/23.2.pinto.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.2.pinto.pdf.

PintodeOliveira:2020:KCC

[Pin20] J. C. Pinto de Oliveira. Kuhn, Condorcet, and Comte: On the justification of the “old” historiography of science. Perspectives on Science, 28(3):375–397, May/June 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00344.

Pitt:2000:NE

[Pit00] Joseph C. Pitt. Note from the Editor. Perspectives on Science, 8 (4):327, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/106361400753373704.

Pitt:2001:DCS

[Pit01a] Joseph C. Pitt. The dilemma of case studies: Toward a Heraclitian philosophy of science. Perspectives on Science, 9(4):373– 382, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.4pitt02.html; http://muse.
jhu.edu/journals/perspectives_on_science/v009/9.4pitt02.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
106361401760375785.

Pitt:2001:RLT

[Pit01b] Joseph C. Pitt. Reconsidering the legacy of Thomas Kuhn; Editor’s introduction. Perspectives on Science, 9(4):371–372, Winter 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.4pitt01.html; http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.4pitt01.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361401760375776.

Pojman:2011:IBP

[Poj11] Paul Pojman. The influence of biology and psychology upon physics: Ernst Mach revisited. Perspectives on Science, 19(2):121– 135, Summer 2011. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v019/19.2.pojman.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v019/19.
2.pojman.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00033.

Psillos:2004:GSC

[Psi04] Stathis Psillos. A glimpse of the Secret Connexion: Harmonising mechanisms with counterfactuals. Perspectives on Science, 12(3): 288–319, Fall 2004. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.3psillos.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v012/12.
3psillos.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/1063614042795426.

Porcelli:2019:AEF

[PT19] Apollonya Maria Porcelli and Amy S. Teller. Asymmetric epistemology: Field notes from training in two disciplines. Perspectives on Science, 27(2):187–213, March/April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00305.

Purves:2018:FSO

[Pur18] Gordon Michael Purves. Fictionalism, semantics, and ontology. Perspectives on Science, 26(1):52–75, January/February 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00267.

Rainger:1993:BGN

[Rai93] Ronald Rainger. Biology, geology, neither, or both: Vertebrate paleontology at the University of Chicago, 1892–1950. Perspectives on Science, 1(3):478–519, Fall 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Ramsey:1998:RER

[Ram98] Jeffry L. Ramsey. Review essay: Recent work in the history and philosophy of chemistry. Perspectives on Science, 6(4):409– 427, Winter 1998. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v006/6.4ramsey.html.

Richmond:2022:EMS

[RB22] Marsha Richmond and Ana Barahona. Erratum: Maria Santesmases, “Women in Early Human Cytogenetics: an Essay on a Gendered History of Chromosome Imaging,” in the special issue, “Heredity and Evolution in an Ibero-American Context,” Perspectives on Science (2020) 28(2): 170–200. https://doi.org/10.1162/posc_{a}_{0}0337.PerspectivesonScience, 30(1) : 199, January/F ebruary2022. CODENP RSIEU. ISSN1063 −

6145(print), 1530−9274(electronic).URLhttpsdirectmiteduposcarticlepdf3011991987186posce00418pdf.See[San20].

Ritson:2015:CBS

[RC15] Sophie Ritson and Kristian Camilleri. Contested boundaries: The string theory debates and ideologies of science. Perspectives on Science, 23(2):192–227, Summer 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v023/23.2.ritson.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.2.ritson.pdf.

Read:2013:MCI

[Rea13] Dwight Read. Modeling cultural idea systems: The relationship between theory models and data models. Perspectives on Science, 21(2):157–174, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v021/21.
2.read.html; http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.read.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00090. Special Issue on Modeling

in the Social Sciences.

Redei:2014:HPA

[Réd14] Miklós Rédei. Hilbert’s 6th Problem and axiomatic quantum field theory. Perspectives on Science, 22(1):80–97, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.1.redei.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v022/22.1.redei.pdf.

Reisch:1997:EEL

[Rei97] George A. Reisch. Economist, epistemologist … and censor? On Otto Neurath’s ‘Index Verborum Prohibitorum’. Perspectives on Science, 5(3):452–480, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=

6370284&site=ehost-live. Special issue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Reinhart:2019:RWF

[Rei19] Katherine M. Reinhart. Richard Waller and the fusion of visual and scientific practice in the early Royal Society. Perspectives on Science, 27(3):435–484, May/June 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00313.

Remmert:2004:WNA

[Rem04] Volker R. Remmert. What’s Nazi about Nazi Science? Recent trends in the history of science in Nazi Germany. Perspectives on Science, 12(4):454–475, Winter 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v012/12.4remmert.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v012/12.4remmert.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/1063614042776003.

Resnik:1998:CIS

[Res98] David B. Resnik. Conflicts of interest in science. Perspectives on Science, 6(4):381–408, Winter 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v006/6.4resnik.html.

Resnik:2000:FIR

[Res00] David B. Resnik. Financial interests and research bias. Perspectives on Science, 8(3):255–285, Fall 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v008/8.
3resnik.html; http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.3resnik.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400750340497.

Richardson:2003:CPS

[RH03a] Alan Richardson and Don Howard. The contexts of philosophy of science. Perspectives on Science, 11(1):1–2, Spring 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/106361403322286689.

Richardson:2003:SIH

[RH03b] Alan Richardson and Don Howard, editors. Special issue: History of science and philosophy of science. MIT Press, Cambridge, MA, USA, 2003. ISSN 1063-6145 (print), 1530-9274 (electronic). i–ii and 1–129 pp. Papers from the 4th Biennial Conference of the International Society for the History of Philosophy of Science (HOPOS 2002) held in Montreal, QC, June 2002, Perspect. Sci. 11 (2003), no. 1.

Rheinberger:2005:GBN

[Rhe05a] Hans-Jörg Rheinberger. Gaston Bachelard and the notion of “Phenomenotechnique”. Perspectives on Science, 13(3):313–328, Fall 2005. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v013/13.3rheinberger01.html;   http://muse.
jhu.edu/journals/perspectives_on_science/v013/13.3rheinberger01.
pdf; http://www.mitpressjournals.org/doi/abs/10.1162/106361405774288026.

Special Issue: Technoscientific Productivity, Part 2.

Rheinberger:2005:RDB

[Rhe05b] Hans-Jörg Rheinberger. A reply to David Bloor: “Toward a Sociology of Epistemic Things”. Perspectives on Science, 13(3): 406–410, Fall 2005. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v013/13.3rheinberger02.html;
http://muse.jhu.edu/journals/perspectives_on_science/v013/
13.3rheinberger02.pdf; http://www.mitpressjournals.org/
doi/abs/10.1162/106361405774287973. Special Issue: Techno-

scientific Productivity, Part 2. See [Blo05].

Richards:1996:REO

[Ric96] Joan L. Richards. Review essay: Observing science in early Victorian England: Recent scholarship on William Whewell. Perspectives on Science, 4(2):231–247, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6369841&site=ehost-live.

Richardson:1997:THS

[Ric97] Alan Richardson. Toward a history of scientific philosophy. Perspectives on Science, 5(3):418–451, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=

6370280&site=ehost-live. Special issue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Richards:2002:KVC

[Ric02a] Richard C. Richards. Kuhnian values and Cladistic parsimony. Perspectives on Science, 10(1):1–27, Spring 2002. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.1richards.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v010/10.1richards.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361402762674780.

Richardson:2002:NHR

[Ric02b] Alan W. Richardson. Narrating the history of reason itself: Friedman, Kuhn, and a constitutive A Priori for the Twenty-First Century. Perspectives on Science, 10(3):253–274, Fall 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.3richardson.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v010/10.3richardson.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
106361402321899050.

Richardson:2008:SSC

[Ric08] Alan Richardson. Solomon’s science without conscience, or, on the coherence of epistemic Newtonianism. Perspectives on Science, 16(3):246–252, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v016/16.3.richardson.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v016/16.3.richardson.pdf; http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2008.16.3.246.

Richmond:2020:SAF

[Ric20] Marsha L. Richmond. South American fieldwork/cytogenetic knowledge: The cytogenetic research program of Sally Hughes-Schrader and Franz Schrader. Perspectives on Science, 28(2):127– 169, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00336.

Rios:2013:MMS

[Rio13] Diego Rios. Models and modeling in the social sciences. Perspectives on Science, 21(2):221–225, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v021/21.
2.rios.html; http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.rios.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00095. Special Issue on Modeling

in the Social Sciences.

Risjord:2000:PEO

[Ris00] Mark W. Risjord. The politics of explanation and the origins of ethnography. Perspectives on Science, 8(1):29–52, Spring 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.1risjord.html; http://muse.jhu.edu/journals/
perspectives_on_science/v008/8.1risjord.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/106361400568000.

Rivelli:2019:MEE

[Riv19] Luca Rivelli. Multilevel ensemble explanations: a case from theoretical biology. Perspectives on Science, 27(1):88–116, January/February 2019. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00301.

Roberts:2017:PPL

[Rob17] Elizabeth F. S. Roberts. Practicing population in Latin America. Perspectives on Science, 25(5):704–711, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00261.

Rojas:2020:DGF

[Roj20] Erica Torrens Rojas. From darkness to gloom: The feminine presence in the teaching of human evolution in Mexico. Perspectives on Science, 28(2):341–373, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00343.

Rolin:1999:CGI

[Rol99] Kristina Rolin. Can gender ideologies influence the practice of the physical sciences? Perspectives on Science, 7(4):510–533, Winter 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v007/7.4rolin.html; http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.4rolin.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.1999.7.4.510.

Rolin:2012:FAV

[Rol12] Kristina Rolin. A feminist approach to values in science. Perspectives on Science, 20(3):320–330, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v020/20.
3.rolin.html; http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.3.rolin.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00068.

Roque:2022:TIP

[Roq22] Ricardo Roque. Transnational isolates: Portuguese colonial race science and the foreign world. Perspectives on Science, 30(1):108– 136, January/February 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/1/108/1987180/posc_a_00404.pdf.

Rossini:2022:IIN

[Ros22] Paolo Rossini. Introduction: The idiosyncratic nature of Renaissance mathematics. Perspectives on Science, 30(3):353–357, May/ June 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL https://direct.mit.edu/posc/article-pdf/
30/3/353/2023835/posc_e_00419.pdf.

Roux:2018:EYA

[Rou18] Sophie Roux. Everything you always wanted to know about the summa quadripartita that Descartes never wrote. Perspectives on Science, 26(5):563–578, September/October 2018.

CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00289.

Rowe:1997:REP

[Row97] David E. Rowe. Review essay: Perspective on Hilbert. Perspectives on Science, 5(4):533–570, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6370944&site=ehost-live.

Rehg:2008:CCA

[RS08] William Rehg and Kent W. (Kent Wade) Staley. The CDF collaboration and argumentation theory: The role of process in objective knowledge. Perspectives on Science, 16(1):1–25, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.1rehg.html; http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.1rehg.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2008.16.1.1.

Rudge:1998:BAS

[Rud98] David Wÿss Rudge. A Bayesian analysis of strategies in evolutionary biology. Perspectives on Science, 6(4):341–360, Winter 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v006/6.4rudge.html.

Rudge:2001:KES

[Rud01] David Wÿss Rudge. Kettlewell from an error Statisticians’s point of view. Perspectives on Science, 9(1):59–77, Spring 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.1rudge.html; http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.1rudge.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/10636140152947795.

Ruphy:2006:EAW

[Rup06] Stéphanie Ruphy. ”Empiricism all the way down”: a defense of the value-neutrality of science in response to Helen Longino’s contextual empiricism. Perspectives on Science, 14(2):189–214, Summer 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_

on_science/v014/14.2ruphy.html; http://muse.jhu.edu/journals/
perspectives_on_science/v014/14.2ruphy.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2006.14.2.189.

Rusu:2017:RSS

[Rus17] Doina-Cristina Rusu. Rethinking Sylva sylvarum: Francis Bacon’s use of Giambattista Della Porta’s Magia naturalis. Perspectives on Science, 25(1):1–35, January/February 2017. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://muse.jhu.edu/article/647987;   http://muse.
jhu.edu/article/647987/pdf.

Ryckman:2003:SSS

[Ryc03] Thomas A. Ryckman. Surplus structure from the standpoint of transcendental idealism: The “World Geometries” of Weyl and Eddington. Perspectives on Science, 11(1): 76–106, Spring 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v011/11.1ryckman.pdf;
http://www.mitpressjournals.org/doi/abs/10.1162/106361403322286715.

Special issue: History of science and philosophy of science (Montreal, QC, 2002).

Sabra:1998:CUA

[Sab98] A. I. Sabra. Configuring the universe: Aporetic, problem solving, and kinematic modeling as themes of Arabic astronomy. Perspectives on Science, 6(3):288–330, Fall 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v006/6.

3sabra.html. See comments [Sal00, Sab00].

Sabra:2000:RAV

[Sab00] A. I. Sabra. Reply to: “Arabic versus Greek astronomy: a debate over the foundations of science. Comments on: ‘Configuring the universe: aporetic, problem solving, and kinematic modeling as themes of Arabic astronomy’ ” [Perspect. Sci. 8 (2000), no. 4, 328–341; MR1873205 (2003b:01014)] by G. Saliba. Perspectives on Science, 8(4):342–345, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v008/8.
4sabra.html; http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.4sabra.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400753373722. See [Sab98, Sal00].

Saliba:2000:AVG

[Sal00] George Saliba. Arabic versus Greek astronomy: a debate over the foundations of science. Comments on: “Configuring the universe: aporetic, problem solving, and kinematic modeling as themes of Arabic astronomy” [Perspect. Sci. 6 (1998), no. 3, 288–330; MR1696321 (2001h:01009)] by A. I. Sabra. Perspectives on Science, 8(4):328–341, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v008/8.
4saliba.html; http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.4saliba.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400753373713. See [Sab98, Sab00].

Santesmases:2020:WEH

[San20] María Jesús Santesmases. Women in early human cytogenetics: an essay on a gendered history of chromosome imaging. Perspectives on Science, 28(2):170–200, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00337. See

erratum [RB22].

Sanchez:2024:SMS

[San24] Grégoire Sanchez. Sympathy and moral sentiments in Maine de Biran’s philosophy. Perspectives on Science, 32(1):28–46, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/32/1/28/2210158/posc_a_00597.pdf.

Steinle:2002:IHS

[SB02] Friedrich Steinle and Richard M. Burian. Introduction: History of science and philosophy of science. Perspectives on Science, 10(4): 391–397, Winter 2002. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/106361402322288020.

Sullivan-Clarke:2019:MMP

[SC19] Andrea Sullivan-Clarke. Misled by metaphor: The problem of ingrained analogy. Perspectives on Science, 27(2):153–170, March/ April 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://www.mitpressjournals.org/doi/abs/
10.1162/posc_a_00303.

Schuster:1995:RED

[Sch95] John A. Schuster. Review essay: Descartes Agonistes: new tales of Cartesian natural philosophy. Perspectives on Science, 3(1): 99–145, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Schickore:2001:EPI

[Sch01] Jutta Schickore. Ever-present impediments: Exploring instruments and methods of microscopy. Perspectives on Science, 9(2):126– 146, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v009/9.2schickore.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v009/9.
2schickore.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361401317447255.

Schickore:2002:UPP

[Sch02] Jutta Schickore. (ab)using the past for present purposes: Exposing contextual and trans-contextual features of error. Perspectives on Science, 10(4):433–456, Winter 2002. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.4schickore.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v010/10.4schickore.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
106361402322288057.

Schmaus:2003:KRF

[Sch03] Warren Schmaus. Kant’s reception in France: Theories of the categories in academic philosophy, psychology, and social science. Perspectives on Science, 11(1):3–34, Spring 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v011/11.
1schmaus.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.1schmaus.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361403322286698. Special issue: His-

tory of science and philosophy of science (Montreal, QC, 2002).

Schonfeld:2006:ACP

[Sch06] Martin Schönfeld. Animal consciousness: Paradigm change in the life sciences. Perspectives on Science, 14(3):354–381, Fall 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_

on_science/v014/14.3schonfeld.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v014/14.3schonfeld.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2006.14.3.354.

Schickore:2011:MTH

[Sch11] Jutta Schickore. More thoughts on HPS: Another 20 years later. Perspectives on Science, 19(4):453–481, Winter 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.4.schickore.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v019/19.4.schickore.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00049.

Schindler:2014:MKT

[Sch14] Samuel Schindler. A matter of Kuhnian theory-choice?: The GWS model and the neutral current. Perspectives on Science, 22(4): 491–522, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v022/22.4.schindler.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v022/22.
4.schindler.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00147.

Schulz:2015:HDS

[Sch15] Armin W. Schulz. The heuristic defense of scientific models: an incentive-based assessment. Perspectives on Science, 23(4):424– 442, Winter 2015. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v023/23.4.schulz.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v023/23.
4.schulz.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00182.

Schmaltz:2018:FCS

[Sch18] Tad M. Schmaltz. French Cartesian scholasticism: Remarks on Descartes and the first Cartesians. Perspectives on Science, 26 (5):579–598, September/October 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00290.

Schickore:2020:MSW

[Sch20] Jutta Schickore. Mess in science and wicked problems. Perspectives on Science, 28(4):482–504, July/August 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00348.

Schlaudt:2021:WTD

[Sch21] Oliver Schlaudt. “Who is there that doesn’t calculate?”: Homo Economicus as a measuring instrument in non-market accounting. Perspectives on Science, 29(6):842–868, November/ December 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/6/842/1972156/posc_a_00396.pdf; https:/
/muse.jhu.edu/article/839327.

Schiaffonati:2022:EEP

[Sch22] Viola Schiaffonati. Explorative experiments: a paradigm shift to deal with severe uncertainty in autonomous robotics. Perspectives on Science, 30(2):284–304, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/30/2/284/2005004/posc_
a_00415.pdf.

Suarez-Diaz:2017:BDB

[SD17] Edna Suárez-Díaz. Blood diseases in the backyard: Mexican “indígenas” as a population of cognition in the mid-1960s. Perspectives on Science, 25(5):606–630, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00253.

Suarez-Diaz:2017:PCP

[SDGDV17] Edna Suárez-Díaz, Vivette García-Deister, and Emily E. Vasquez. Populations of cognition: Practices of inquiry into human populations in Latin America. Perspectives on Science, 25(5): 551–563, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/POSC_a_00256.

Schweer:2023:DMC

[SE23] Julie Schweer and Marcus Elstner. Dealing with molecular complexity. atomistic computer simulations and scientific explanation. Perspectives on Science, 31(5):594–626, September-October 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
5/594/2158907/posc_a_00594.pdf.

Segev:2022:SMJ

[Seg22] Ran Segev. ‘For the sciences migrate, just like people’: The case of botanical knowledge in the Early Modern Iberian empires. Perspectives on Science, 30(4):732–756, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
4/732/2038170/posc_a_00409.pdf.

Sent:2000:HSC

[Sen00] Esther-Mirjam Sent. Herbert A. Simon as a cyborg scientist. Perspectives on Science, 8(4):380–406, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.4sent.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v008/8.4sent.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361400753373759.

Seselja:2021:ESI

[Šeš21] Dunja Šešelja. Exploring scientific inquiry via: Agent-based modelling. Perspectives on Science, 29(4):537–557, July/ August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/537/1959232/posc_a_00382.pdf; https:/
/muse.jhu.edu/article/804410.

Shrader-Frechette:2000:RHM

[SF00] K. S. (Kristin Sharon) Shrader-Frechette. Radiobiogical hormesis, methodological value judgments, and metascience. Perspectives on Science, 8(4):367–379, Winter 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v008/8.
4frechette.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v008/8.4frechette.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361400753373740.

Schmidt:2001:RAV

[SG01] Dietmar Schmidt and Andrew Gledhill. Refuse archaeology: Virchow–Schliemann–Freud. Perspectives on Science, 9(2):210– 232, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v009/9.2schmidt.html;
http://muse.jhu.edu/journals/perspectives_on_science/v009/
9.2schmidt.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361401317447291.

Sgarbi:2022:GEA

[Sga22] Marco Sgarbi. Galileo and the epistemology of anatomy. Perspectives on Science, 30(5):903–923, September/October 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
5/903/2053601/posc_a_00568.pdf.

Smith:2017:CMD

[SGD17] Lindsay A. Smith and Vivette García-Deister. Capturing Los Migrantes Desaparecidos: Crisis, unknowability, and the making of the missing. Perspectives on Science, 25(5):680–697, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00259.

Steel:2018:GSV

[SGMB18] Daniel Steel, Chad Gonnerman, Aaron M. McCright, and Itai Bavli. Gender and scientists’ views about the value-free ideal. Perspectives on Science, 26(6):619–657, November/December 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00292.

Shapin:1995:CLC

[Sha95] Steven Shapin. Cordelia’s love: Credibility and the social studies of science. Perspectives on Science, 3(3):255–275, Fall 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Shapiro:1996:GAN

[Sha96] Alan E. Shapiro. The gradual acceptance of Newton’s theory of light and color, 1672–1727. Perspectives on Science, 4(1):59–140, Spring 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6663277&site=ehost-live.

Shank:2002:RPP

[Sha02] Michael H. Shank. Regiomontanus on Ptolemy, physical orbs, and astronomical fictionalism: Goldsteinian themes in the “Defense of Theon against George of Trebizond”. Perspectives on Science, 10(2):179–207, Summer 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v010/10.
2shank.html; http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.2shank.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361402321147522. New foundations in

the history of astronomy (Pittsburgh, PA, 1999).

Shapiro:2003:HBM

[Sha03] Lisa Shapiro. The health of the body-machine? or Seventeenth Century mechanism and the concept of health. Perspectives on Science, 11(4):421–442, Winter 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v011/11.4shapiro.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v011/11.4shapiro.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361403773082252. Special Issue: Early

Modern Philosophy and Biological Thought.

Shapiro:2008:TNY

[Sha08] Alan E. Shapiro. Twenty-nine years in the making: Newton’s Opticks. Perspectives on Science, 16(4):417–438, Winter 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.4.shapiro.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v016/16.4.shapiro.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2008.16.4.417.

Shavit:2016:LIR

[Sha16] Ayelet Shavit. “Location” incommensurability and “replication” indeterminacy: Clarifying an entrenched conflation by using an involved approach. Perspectives on Science, 24(4):425–442, July/ August 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/625739;
 http://muse.jhu.edu/article/625739/pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/POSC_a_00215.

Shea:1994:REG

[She94] William R. Shea. Review essay: Galileo in the Nineties. Perspectives on Science, 2(4):476–487, Winter 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Shipworth:2013:VAH

[Shi13] David Shipworth. The vernacular architecture of household energy models. Perspectives on Science, 21(2):250–266, Summer 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.2.shipworth.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.2.shipworth.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/

POSC_a_00098. Special Issue on Modeling in the Social Sciences.

Sieg:2014:WHA

[Sie14] Wilfried Sieg. The ways of Hilbert’s axiomatics: Structural and formal. Perspectives on Science, 22(1):133–157, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.1.sieg.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v022/22.1.sieg.pdf.

Simons:2019:BLS

[Sim19] Massimiliano Simons. Bruno Latour and the secularization of science. Perspectives on Science, 27(6):925–954, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00330.

Simons:2022:HEH

[Sim22] Massimiliano Simons. History as engagement: The historical epistemology of Raymond Aron. Perspectives on Science, 30(4): 757–782, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/4/757/2038159/posc_a_00410.pdf.

Sinnema:2014:WAM

[Sin14] Peter W. Sinnema. “We have adventured to make the Earth hollow”: Edmond Halley’s extravagant hypothesis. Perspectives on Science, 22(4):423–448, Winter 2014. CODEN PRSIEU.

ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
4.sinnema.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.4.sinnema.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00144.

Sismondo:1995:SDF

[Sis95] Sergio Sismondo. The scientific domains of feminist standpoints. Perspectives on Science, 3(1):49–65, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Shew:2018:CAT

[SJ18] Ashley Shew and Keith Johnson. Companion animals as technologies in biomedical research. Perspectives on Science, 26(3):400– 417, May/June 2018. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00279.

Skipper:2004:CLM

[Ski04] Robert A. Skipper, Jr. Calibration of laboratory models in population genetics. Perspectives on Science, 12(4):369–393, Winter 2004. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v012/12.4skipper.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v012/12.
4skipper.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/1063614042776012.

Seguin:2023:BLS

[SL23] Eve Seguin and Laurent-Olivier Lord. Bruno Latour’s science is politics by other means: Between politics and ontology. Perspectives on Science, 31(1):9–39, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
1/9/2074866/posc_a_00579.pdf.

Slowik:2009:NMS

[Slo09] Edward Slowik. Newton’s metaphysics of space: a “Tertium Quid” betwixt substantivalism and relationism, or merely a “God of the Rational Mechanical Gaps”? Perspectives on Science, 17(4):429– 456, Winter 2009. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v017/17.4.slowik.html;   http:

//muse.jhu.edu/journals/perspectives_on_science/v017/17.
4.slowik.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2009.17.4.429.

Siao:2020:DDG

[SM20] José Franco Monte Sião and Lilian Al-Chueyr Pereira Martins. Dobzhansky and Dreyfus’s group: The introduction of natural population genetics studies in Brazil (1943–1960). Perspectives on Science, 28(2):244–276, March/April 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00340.

Smith:2009:UGP

[Smi09] Justin E. H. Smith. “The Unity of the Generative Power”: Modern taxonomy and the problem of animal generation. Perspectives on Science, 17(1):78–104, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v017/17.1.smith.html;
http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.1.smith.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2009.17.1.78.

Smith:2012:CKH

[Smi12] Justin E. H. Smith. “Curious Kinks of the Human Mind”: Cognition, natural history, and the concept of race. Perspectives on Science, 20(4):504–529, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v020/20.4.smith.html;
http://muse.jhu.edu/journals/perspectives_on_science/v020/
20.4.smith.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00083.

Smith:2018:ICT

[Smi18] Eden T. Smith. Interdependent concepts and their independent uses: Mental imagery and hallucinations. Perspectives on Science, 26(3):360–399, May/June 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00278.

Snyder:1997:MWD

[Sny97] Laura J. Snyder. The Mill–Whewell debate: much ado about induction. Perspectives on Science, 5(2):159–198, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=10780088&site=ehost-live.

Solomon:1995:PTN

[Sol95] Miriam Solomon. The pragmatic turn in naturalistic philosophy of science. Perspectives on Science, 3(2):206–230, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Solomon:1996:IEI

[Sol96] Miriam Solomon. Information and the ethics of information control in science. Perspectives on Science, 4(2):195–206, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6369832&site=ehost-live.

Solomon:2003:WLR

[Sol03] Miriam Solomon. The whiptail lizard reconsidered. Perspectives on Science, 11(3):318–325, Fall 2003. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v011/11.3solomon.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v011/11.3solomon.pdf;   http://www.mitpressjournals.org/
doi/abs/10.1162/106361403773062669.

Solomon:2008:RC

[Sol08] Miriam Solomon. Responses to critics. Perspectives on Science, 16 (3):280–284, Fall 2008. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.3.solomon.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v016/16.
3.solomon.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2008.16.3.280.

Solomon:2010:KAP

[Sol10] Stephanie Solomon. Kuhn’s alternative path: Science and the social resistance to criticism. Perspectives on Science, 18(3): 352–368, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.3.solomon.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v018/18.
3.solomon.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00013.

Solomon:2012:SRS

[Sol12] Miriam Solomon. Socially responsible science and the unity of values. Perspectives on Science, 20(3):331–338, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.3.solomon.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v020/20.3.solomon.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00069.

Simoncelli:2015:MCA

[SP15] Tania Simoncelli and Sandra S. Park. Making the case against gene patents. Perspectives on Science, 23(1):106–145, Spring 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.1.simoncelli.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v023/23.1.simoncelli.
pdf.

SilvadaCosta:2023:NCP

[SP23] Wigson Rafael Silva da Costa and Antonio Augusto Passos Videira. Nietzsche and cosmology: a possible way of enriching the practice of science. Perspectives on Science, 31(4):503–533, ???? 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://muse.jhu.edu/pub/6/article/902864.

Spear:2004:CCS

[Spe04] Joseph Howard Spear. Cumulative change in scientific production: Research technologies and the structuring of new knowledge. Perspectives on Science, 12(1):55–85, Spring 2004. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v012/12.
1spears.html; http://muse.jhu.edu/journals/perspectives_
on_science/v012/12.1spears.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/106361404773843346.

Spencer:2015:BBS

[Spe15] Matt Spencer. Brittleness and bureaucracy: Software as a material for science. Perspectives on Science, 23(4):466–484, Winter 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v023/23.4.spencer.html;   http://muse.jhu.edu/

journals/perspectives_on_science/v023/23.4.spencer.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00184.

Staley:1999:GES

[Sta99] Kent W. Staley. Golden events and statistics: What’s wrong with Galison’s image/logic distinction? Perspectives on Science, 7(2): 196–230, Summer 1999. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc.1999.7.2.196.

Staley:2012:DFT

[Sta12] Kent W. Staley. Dirac’s ‘Fine-Tuning Problem’: a constructive use of anachronism? Perspectives on Science, 20(4):476– 503, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.4.staley.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v020/20.
4.staley.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00082.

Steinle:2002:EHP

[Ste02] Friedrich Steinle. Experiments in history and philosophy of science. Perspectives on Science, 10(4):408–432, Winter 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v010/10.4steinle02.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v010/10.4steinle02.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
106361402322288048.

Stevens:2011:CSH

[Ste11] Hallam Stevens. Coding sequences: a history of sequence comparison algorithms as a scientific instrument. Perspectives on Science, 19(3):263–299, Fall 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v019/19.
3.stevens.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.3.stevens.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00042.

Steel:2016:CCS

[Ste16a] Daniel Steel. Climate change and second-order uncertainty: Defending a generalized, normative, and structural argument from inductive risk. Perspectives on Science, 24(6):696–721, November/ December 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://muse.jhu.edu/article/635023;
 http://muse.jhu.edu/article/635023/pdf.

Stegenga:2016:HHH

[Ste16b] Jacob Stegenga. Hollow hunt for harms. Perspectives on Science, 24(5):481–504, September/October 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/628778;   http://muse.jhu.edu/article/
628778/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00220.

Stillwell:2014:IEH

[Sti14] John Stillwell. Ideal elements in Hilbert’s geometry. Perspectives on Science, 22(1):35–55, Spring 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
1.stillwell.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.1.stillwell.pdf.

Stoltzner:2018:FDM

[Stö18] Michael Stöltzner. Feynman diagrams: Modeling between physics and mathematics. Perspectives on Science, 26(4):482–500, July/ August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://www.mitpressjournals.org/
doi/abs/10.1162/posc_a_00284.

Storni:2024:BDN

[Sto24] Marco Storni. Beyond Descartes: Noël Regnault and Eighteenth-Century French Cartesianism. Perspectives on Science, 32(2): 230–261, March-April 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/32/2/230/2345646/posc_a_00623.pdf.

Strazzoni:2023:UPD

[Str23] Andrea Strazzoni. The use and plagiarism of Descartes’s Traité de l’homme by Henricus Regius: a reassessment. Perspectives on Science, 31(5):627–683, September-October 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
5/627/2158918/posc_a_00587.pdf.

Stuewer:1994:OLD

[Stu94] Roger H. Stuewer. The origin of the liquid-drop model and the interpretation of nuclear fission. Perspectives on Science, 2(1):76– 129, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://alsos.wlu.edu/information.

aspx?id=1736. See errata [Ano94b].

Stump:1997:RUM

[Stu97] David J. Stump. Reconstructing the unity of mathematics circa 1900. Perspectives on Science, 5(3):383–417, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6370279&site=ehost-live. Special is-

sue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Stuart:2014:CST

[Stu14] Michael T. Stuart. Cognitive science and thought experiments: a refutation of Paul Thagard’s skepticism. Perspectives on Science, 22(2):264–287, Summer 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v022/22.2.stuart.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v022/22.2.stuart.pdf.

Suppe:1993:CSC

[Sup93] Frederick Suppe. Credentialling scientific claims. Perspectives on Science, 1(2):153–203, Summer 1993. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Seguin:2023:ISP

[SV23] Eve Seguin and Dominique Vinck. Introduction: Science is politics by other means revisited. Perspectives on Science, 31(1):1– 8, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/31/1/1/2073139/posc_e_00578.pdf.

Saunders:2002:TC

[SvJB02] Barbara Saunders and J. van (Jaap) Brakel. The trajectory of color. Perspectives on Science, 10(3):302–355, Fall 2002. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_

on_science/v010/10.3saunders.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v010/10.3saunders.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/106361402321899078.

Swerdlow:2012:CAA

[Swe12] N. M. Swerdlow. Copernicus and astrology, with an appendix of translations of primary sources. Perspectives on Science, 20(3): 353–378, Fall 2012. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.3.swerdlow.html;   http:
//muse.jhu.edu/journals/perspectives_on_science/v020/20.
3.swerdlow.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/POSC_a_00071.

Swerdlow:2013:PWR

[Swe13] N. M. Swerdlow. On Professor Westman’s reply to Copernicus and astrology, with an appendix of translations of additional primary sources. Perspectives on Science, 21(3):384–385, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.3.swerdlow.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v021/21.3.swerdlow.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_c_00105.

Sylla:1998:EMP

[Syl98] Edith Dudley Sylla. The emergence of mathematical probability from the perspective of the Leibniz–Jacob Bernoulli correspondence. Perspectives on Science, 6(1–2):41–76, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v006/6.1sylla.html. Special Issue: Leibniz and the

Sciences.

Tal:2021:TMR

[Tal21] Eran Tal. Two myths of representational measurement. Perspectives on Science, 29(6):701–741, November/December 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/29/6/701/
1972160/posc_a_00391.pdf; https://muse.jhu.edu/article/
839329.

Tamborini:2021:MTS

[Tam21a] Marco Tamborini. The material turn in the study of form: From bio-inspired robots to robotics-inspired morphology. Perspectives on Science, 29(5):643–665, September/October 2021. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/29/5/643/
1964253/posc_a_00388.pdf; https://muse.jhu.edu/article/
810915.

Tamborini:2021:MRF

[Tam21b] Marco Tamborini. Morphogenesis — “the riddles of form” in Twenty-First Century science. Perspectives on Science, 29(5):559–567, September/October 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://

direct.mit.edu/posc/article-pdf/29/5/559/1964251/posc_
e_00383.pdf; https://muse.jhu.edu/article/810917.

Tanghe:2019:FWW

[Tan19] Koen B. Tanghe. The fate of William Whewell’s four palætiological domains: a comparative study. Perspectives on Science, 27(6): 810–838, November/December 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00326.

Tautschnig:2022:CAP

[Tau22] Irina Tautschnig. Constructing authority in the paratext: The poems to Johannes Hevelius’ Selenographia. Perspectives on Science, 30(6):1005–1041, November/December 2022. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL https://direct.mit.edu/posc/article-pdf/30/6/1005/
2059526/posc_a_00569.pdf.

Taylor:1995:BCE

[Tay95] Peter J. Taylor. Building on construction: an exploration of heterogeneous constructionism, using an analogy from psychology and a sketch from socioeconomic modeling. Perspectives on Science, 3 (1):66–98, Spring 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Teira:2016:DMA

[Tei16] David Teira. Debiasing methods and the acceptability of experimental outcomes. Perspectives on Science, 24(6):722–743, November/December 2016. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/article/
635024;   http://muse.jhu.edu/article/635024/pdf.

Tessicini:2022:DCE

[Tes22] Dario Tessicini. Defining “Cosmology” in the early modern system of knowledge, 1530–1621. Perspectives on Science, 30(5):826– 850, September/October 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.edu/

posc/article-pdf/30/5/826/2054826/posc_a_00561.pdf.

Thagard:2014:TEC

[Tha14] Paul Thagard. Thought experiments considered harmful. Perspectives on Science, 22(2):288–305, Summer 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v022/22.
2.thagard.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.2.thagard.pdf.

Toepfer:2021:FFC

[Toe21] Georg Toepfer. Forms as forces: The causal regime of morphology in biology. Perspectives on Science, 29(5):627–642, September/October 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/5/627/1964243/posc_a_00387.pdf; https:/
/muse.jhu.edu/article/810914.

Tanghe:2021:IHE

[TPDB21] Koen B. Tanghe, Lieven Pauwels, Alexis De Tiège, and Johan Braeckman. Interpreting the history of evolutionary biology through a Kuhnian prism: Sense or nonsense? Perspectives on Science, 29(1):1–35, January/February 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00359.

Tsou:2003:RFA

[Tso03] Jonathan Y. Tsou. Reconsidering Feyerabend’s “Anarchism”. Perspectives on Science, 11(2):208–235, Summer 2003. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v011/11.2tsou.html; http://muse.jhu.edu/journals/
perspectives_on_science/v011/11.2tsou.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/106361403322495894.

Tuboly:2020:KMS

[Tub20] Adam Tamas Tuboly. Knowledge missemination: L. Susan Stebbing, C. E. M. Joad, and Philipp Frank on the philosophy of the physicists. Perspectives on Science, 28(1):1–34, January/February 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00331;   https://muse.jhu.edu/article/749741.

Tucker:2007:PTF

[Tuc07] Aviezer Tucker. The political theory of French science studies in context. Perspectives on Science, 15(2):202–221, Summer 2007. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.2tucker.html; http://muse.jhu.edu/journals/
perspectives_on_science/v015/15.2tucker.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/posc.2007.15.
2.202.

Turner:2023:LSP

[Tur23] Stephen Turner. Latour and Schmitt: Political theology and science. Perspectives on Science, 31(1):40–56, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
1/40/2073137/posc_a_00580.pdf.

Tuunainen:2001:COT

[Tuu01] Juha Tuunainen. Constructing objects and transforming experimental systems. Perspectives on Science, 9(1):78–105, Spring 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v009/9.1tuunainen.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v009/9.1tuunainen.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/10636140152947803.

Tavoni:2022:UIA

[TV22] Massimo Tavoni and Giovanni Valente. Uncertainty in integrated assessment modeling of climate change. Perspectives on Science, 30(2):321–351, March/April 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/30/2/321/2005016/posc_a_00417.pdf.

Tweney:2006:DDH

[Twe06a] Ryan D. Tweney. Discovering discovery: How Faraday found the first metallic colloid. Perspectives on Science, 14(1):97–121, Spring 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.1tweney02.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v014/14.1tweney02.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.

2006.14.1.97. Special Issue: Cognitive Psychology and Faraday’s Experiments.

Tweney:2006:TCH

[Twe06b] Ryan D. Tweney. Toward a cognitive-historical understanding of Michael Faraday’s research: Editor’s introduction. Perspectives on Science, 14(1):1–6, Spring 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v014/14.
1tweney01.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.1tweney01.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2006.14.1.1. Special Issue: Cogni-

tive Psychology and Faraday’s Experiments.

Uebel:2008:WRP

[Ueb08] Thomas E. (Thomas Ernst) Uebel. Writing a revolution: On the production and early reception of the Vienna Circle’s Manifesto. Perspectives on Science, 16(1):70–102, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.1uebel.html; http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.1uebel.pdf;   http://www.
mitpressjournals.org/doi/abs/10.1162/posc.2008.16.1.70.

Uebel:2013:LPL

[Ueb13] Thomas Uebel. “Logical Positivism”–“Logical Empiricism”: What’s in a name? Perspectives on Science, 21(1):58–99, Spring 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.1.uebel.html; http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.1.uebel.pdf;   http://
www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00086.

Uebel:2020:IAP

[Ueb20] Thomas Uebel. Intersubjective accountability: Politics and philosophy in the Left Vienna Circle. Perspectives on Science, 28(1):35–62, January/February 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

www.mitpressjournals.org/doi/abs/10.1162/posc_a_00332;
 https://muse.jhu.edu/article/749742.

Urreiztieta:2010:MAM

[Urr10] Carlos Calderón Urreiztieta. The monochord according to Marin Mersenne: Bits, atoms, and some surprises. Perspectives on Science, 18(1):77–97, Spring 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v018/18.
1.urreiztieta.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v018/18.1.urreiztieta.pdf; http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2010.18.1.77.

VanAcker:2011:IUV

[Van11] Wouter Van Acker. Internationalist utopias of visual education: The graphic and scenographic transformation of the Universal Encyclopaedia in the work of Paul Otlet, Patrick Geddes, and Otto Neurath. Perspectives on Science, 19(1):32–80, Spring 2011. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v019/19.1.van-acker.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v019/19.1.van-acker.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00025.

VanBouwel:2015:TDM

[Van15] Jeroen Van Bouwel. Towards democratic models of science: Exploring the case of scientific pluralism. Perspectives on Science, 23(2):149–172, Summer 2015. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v023/23.2.bouwel.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v023/23.2.bouwel.pdf.

Vanzo:2016:I

[Van16] Alberto Vanzo. Introduction. Perspectives on Science, 24(3):255– 263, May/June 2016. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/article/
618104;   http://muse.jhu.edu/article/618104/pdf.

Vargas-Dominguez:2017:POM

[VD17] Joel Vargas-Domínguez. The “problematic” Otomi: Metabolism, nutrition, and the classification of indigenous populations in Mexico in the 1930’s. Perspectives on Science, 25(5):564–584, September/October 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00254.

vanDijk:2023:CEP

[vD23] Niels van Dijk. Constitutional ecology of practices. Bringing law, robots and epigrams into Latourian cosmopolitics. Perspectives on Science, 31(1):159–185, January/February 2023. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/31/
1/159/2073133/posc_a_00585.pdf.

vanderSluijs:2012:CGT

[vdS12] Marinus A. van der Sluijs. Closing gaps in traditional sky lore. Perspectives on Science, 20(1):1–43, Spring 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v020/20.
1.van-der-sluijs.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v020/20.1.van-der-sluijs.pdf; http:
//www.mitpressjournals.org/doi/abs/10.1162/POSC_a_00055.

Valleriani:2022:HPH

[VFN22] Matteo Valleriani, Beate Federau, and Olya Nicolaeva. The hidden praeceptor: How Georg Rheticus taught geocentric cosmology to Europe. Perspectives on Science, 30(3):407–436, May/June 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
3/407/2023847/posc_a_00421.pdf.

Vicedo:1995:RES

[Vic95] Marga Vicedo. Review essay: Scientific styles: toward some common ground in the history, philosophy and sociology of science. Perspectives on Science, 3(2):231–254, Summer 1995. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Vicente:2000:SEP

[Vic00] Kim J. Vicente. Is science an evolutionay process? Evidence from miscitation of the scientific literature. Perspectives on Science, 8(1):53–69, Spring 2000. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v008/8.1vicente.html;
http://muse.jhu.edu/journals/perspectives_on_science/v008/
8.1vicente.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361400568019.

Vigni:2022:DMO

[Vig22] Fabrizio Li Vigni. Data and model operations in computational sciences: The examples of computational embryology and epidemiology. Perspectives on Science, 30(4):696–731, July/August 2022. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL https://direct.mit.edu/posc/article-pdf/30/
4/696/2038173/posc_a_00408.pdf.

Vincenti:2019:EEM

[Vin19] Denise Vincenti. Experience and experimentation: Medicine, psychiatry and experimental psychology in Paul Janet. Perspectives on Science, 27(5):704–738, September/October 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00322.

Vincenti:2024:PJP

[Vin24] Denise Vincenti. Pierre Janet: a psychological reading of Maine de Biran’s theory of the unconscious. Perspectives on Science, 32(1): 102–126, January-February 2024. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL https://direct.mit.

edu/posc/article-pdf/32/1/102/2210172/posc_a_00601.pdf.

Vismann:2001:LR

[Vis01] Cornelia Vismann. The Love of ruins. Perspectives on Science, 9(2):196–209, Summer 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v009/9.2vismann.html;
http://muse.jhu.edu/journals/perspectives_on_science/v009/
9.2vismann.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/106361401317447282.

Wagner:2008:PSR

[Wag08] Roy Wagner. Post-structural readings of a logico-mathematical text. Perspectives on Science, 16(2):196–230, Summer 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.2.wagner.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v016/16.2.wagner.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2008.16.2.196.

Walstad:2001:SFM

[Wal01] Allan Walstad. On science as a free market. Perspectives on Science, 9(3):324–340, Fall 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v009/9.3walstad.html;
http://muse.jhu.edu/journals/perspectives_on_science/v009/
9.3walstad.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/10636140160176170.

Watkins:1997:LMN

[Wat97] Eric Watkins. The laws of motion from Newton to Kant. Perspectives on Science, 5(3):311–348, Fall 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Special issue: Selections from the First Annual Conference for the History of Philosophy of Science (Roanoke, VA, 1996).

Watkins:1998:PEH

[Wat98] Eric Watkins. From pre-established harmony to physical influx: Leibniz’s reception in Eighteenth Century Germany. Perspectives on Science, 6(1–2):136–203, Spring/Summer 1998. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v006/6.1watkins.html. Special Issue: Leibniz and

the Sciences.

Waters:2014:LGM

[Wat14] Tony Waters. Of looking glasses, mirror neurons, culture, and meaning. Perspectives on Science, 22(4):616–649, Winter 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.4.waters.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v022/22.4.waters.pdf;

 http://www.mitpressjournals.org/doi/abs/10.1162/POSC_
a_00152.

Wayne:1996:TUC

[Way96] Andrew Wayne. Theoretical unity: The case of the Standard Model. Perspectives on Science, 4(4):391–407, Winter 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6370048&site=ehost-live.

Westman:2013:CQR

[Wes13] Robert S. Westman. The Copernican question revisited: a reply to Noel Swerdlow and John Heilbron. Perspectives on Science, 21(1):100–136, Spring 2013. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v021/21.
1.westman.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v021/21.1.westman.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00087.

Wieber:2020:MPS

[WH20] Frédéric Wieber and Alexandre Hocquet. Models, parameterization, and software: Epistemic opacity in computational chemistry. Perspectives on Science, 28(5):610–629, September/October 2020. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00352.

Whitt:1999:VBB

[Whi99] Laurie Anne Whitt. Value-bifurcation in bioscience: The rhetoric of research justification. Perspectives on Science, 7(4):413– 446, Winter 1999. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v007/7.4whitt.html; http://muse.
jhu.edu/journals/perspectives_on_science/v007/7.4whitt.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.1999.7.4.413.

Whistler:2019:TES

[Whi19] Daniel Whistler. ”True empiricism”: The stakes of the Cousin–Schelling controversy. Perspectives on Science, 27(5): 739–765, September/October 2019. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://www.

mitpressjournals.org/doi/abs/10.1162/posc_a_00323.

Wilson:2021:TEU

[Wil21] Joseph Wilson. Two exploratory uses for general circulation models in climate science. Perspectives on Science, 29(4):493–509, July/August 2021. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL https://direct.mit.edu/posc/
article-pdf/29/4/493/1959226/posc_a_00380.pdf; https:/
/muse.jhu.edu/article/804408.

Wittje:2016:CSN

[Wit16] Roland Wittje. Concepts and significance of noise in acoustics: Before and after the Great War. Perspectives on Science, 24(1): 7–28, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v024/24.1.wittje.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v024/24.1.wittje.pdf.

Wolfe:2009:CHN

[Wol09] Charles T. Wolfe. “Cabinet d’Histoire Naturelle,” or: The interplay of nature and artifice in Diderot’s naturalism. Perspectives on Science, 17(1):58–77, Spring 2009. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v017/17.1.wolfe.html;
http://muse.jhu.edu/journals/perspectives_on_science/v017/
17.1.wolfe.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/posc.2009.17.1.58.

Wray:1997:RC

[Wra97] K. Brad Wray. Rational communities. Perspectives on Science, 5(2):232–254, Summer 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
10780092&site=ehost-live.

Wray:2010:KC

[Wra10] K. Brad Wray. Kuhn’s constructionism. Perspectives on Science, 18 (3):311–327, Fall 2010. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v018/18.3.wray.html; http://muse.
jhu.edu/journals/perspectives_on_science/v018/18.3.wray.

pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
POSC_a_00014.

Wright:2016:PFT

[Wri16] Aaron Sidney Wright. The physics of forgetting: Thermodynamics of information at IBM 1959–1982. Perspectives on Science, 24(1):112–141, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v024/24.1.wright.
html;   http://muse.jhu.edu/journals/perspectives_on_science/
v024/24.1.wright.pdf.

Wuthrich:2018:EWS

[Wüt18] Adrian Wüthrich. The exigencies of war and the stink of a theoretical problem: Understanding the Genesis of Feynman’s quantum electrodynamics as mechanistic modelling at different levels. Perspectives on Science, 26(4):501–520, July/August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00285.

Wylie:1996:EDA

[Wyl96] Alison Wylie. Ethical dilemmas in archaeological practice: Looting, repatriation, stewardship, and the (trans)formation of disciplinary identity. Perspectives on Science, 4(2):154–194, Summer 1996. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://search.ebscohost.com/login.aspx?
direct=true&db=aph&AN=6369828&site=ehost-live.

Wylie:1999:RUW

[Wyl99] Alison Wylie. Rethinking unity as a “Working Hypothesis” for philosophy: How archaeologists exploit the disunities of science. Perspectives on Science, 7(3):293–317, Fall 1999. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v007/7.
3wylie.html; http://muse.jhu.edu/journals/perspectives_
on_science/v007/7.3wylie.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.1999.7.3.293.

Wylie:2008:MSE

[Wyl08] Alison Wylie. A more social epistemology: Decision vectors, epistemic fairness, and consensus in Solomon’s social empiricism. Perspectives on Science, 16(3):237–240, Fall 2008. CODEN PRSIEU.

ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v016/16.
3.wylie.html; http://muse.jhu.edu/journals/perspectives_
on_science/v016/16.3.wylie.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/posc.2008.16.3.237.

Yeang:2016:UNT

[YB16] Chen-Pang Yeang and Joan Lisa Bromberg. Understanding noise in Twentieth-Century physics and engineering. Perspectives on Science, 24(1):1–6, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

muse.jhu.edu/journals/perspectives_on_science/v024/24.
1.yeang01.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v024/24.1.yeang01.pdf.

Yeang:2016:TMA

[Yea16] Chen-Pang Yeang. Two mathematical approaches to random fluctuations. Perspectives on Science, 24(1):45–72, January/February 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v024/24.1.yeang.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v024/24.1.yeang.pdf.

Ylikoski:2014:ABS

[Yli14] Petri Ylikoski. Agent-based simulation and sociological understanding. Perspectives on Science, 22(3):318–335, Fall 2014. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v022/22.3.ylikoski.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v022/22.3.ylikoski.
pdf.

Young:2016:TTR

[You16] Mark Thomas Young. Technology and technique: The role of skill in the practice of scientific observation. Perspectives on Science, 24(4):396–415, July/August 2016. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/625737;   http://muse.jhu.edu/article/
625737/pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/POSC_a_00213.

Young:2017:MLM

[You17] Mark Thomas Young. Manual labor and ‘mean mechanicks’: Bacon’s mechanical history and the deprecation of craft skills in early modern science. Perspectives on Science, 25(4):521–550, July/August 2017. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00252.

Zahle:2017:PIC

[Zah17] Julie Zahle. Privacy, informed consent, and participant observation. Perspectives on Science, 25(4):465–487, July/August 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/POSC_a_00250.

ZamoraBonilla:2006:SST

[Zam06] Jesús P. Zamora Bonilla. Science studies and the theory of games. Perspectives on Science, 14(4):525–557, Winter 2006. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v014/14.4bonilla.html;   http://muse.jhu.edu/
journals/perspectives_on_science/v014/14.4bonilla.pdf;
 http://www.mitpressjournals.org/doi/abs/10.1162/posc.
2006.14.4.525.

Zik:2001:SIT

[Zik01] Yaakov Zik. Science and instruments: The telescope as a scientific instrument at the beginning of the Seventeenth Century. Perspectives on Science, 9(3):259–284, Fall 2001. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.

edu/journals/perspectives_on_science/v009/9.3zik.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v009/9.3zik.pdf;   http://www.mitpressjournals.org/doi/
abs/10.1162/10636140160176143.

Zollinger:1997:LPS

[Zol97] Heinrich Zollinger. Logic and psychology of scientific discoveries: a case study in contemporary chemistry. Perspectives on Science, 5(4):516–532, Winter 1997. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://

search.ebscohost.com/login.aspx?direct=true&db=aph&AN=
6370939&site=ehost-live.

Zuppone:2017:IAE

[Zup17] Romina Zuppone. An internal answer to the experimenters’ regress through the analysis of the semantics of experimental results and their representational content. Perspectives on Science, 25 (1):95–123, January/February 2017. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.

jhu.edu/article/647990;   http://muse.jhu.edu/article/
647990/pdf.